### v1

# Predicting Drug-Gene Relations via Analogy Tasks with Word Embeddings

Abstract
[…]
Introduction
Natural language processing (NLP) is a computer technology for processing human language. NLP is used in various applications, such as machine translation [1, 2], sentiment analysis [3, 4], and sentence similarity computation [5, 6]. Many of these applications use models such as skip-gram [7, 8] or BERT [4] to convert words in text to embeddings or distributed representations, which are feature vectors with hundreds of dimensions. In particular, skip-gram is a method that learns high-performance embeddings by predicting the surrounding words of a word in a sentence. These embeddings form the foundational building blocks of recent advances in large language models, enabling models to understand and generate human language with unprecedented accuracy and coherence.
Mikolov et al. [7] showed that skip-gram embeddings have properties such as $\mathbf{u}_{\mathrm{\emph{king}}}-\mathbf{u}_{\mathrm{\emph{man}}}+\mathbf{u}_{\mathrm{\emph{woman}}}\approx\mathbf{u}_{\mathrm{\emph{queen}}}$, where the vector $\mathbf{u}_{w}$ represents the embedding of word $w$. In the embedding space, the vector differences such as $\mathbf{u}_{\mathrm{\emph{king}}}-\mathbf{u}_{\mathrm{\emph{man}}}$ and $\mathbf{u}_{\mathrm{\emph{queen}}}-\mathbf{u}_{\mathrm{\emph{woman}}}$ can be seen as vectors representing the royalty relation. Since these relations are not explicitly taught during the training of skip-gram, the embeddings acquire these properties spontaneously. Solving questions such as “If man corresponds to king, what does woman correspond to?” requires understanding the relations between words. This type of problem is known as an analogy task and is used to evaluate a model’s language comprehension and reasoning skills. In this example, the model must understand the royalty relations from man to king and apply a similar relation to woman. Allen and Hospedales [9] explained why analogy computation with such simple vector arithmetic can effectively solve analogy tasks.
Recently, NLP methods have also gained attention in the field of biology [10, 11, 12]. However, since traditional skip-gram models are typically trained on web corpora, they do not properly handle the specialized terms that appear in biological texts. In particular, when multiple words represent the same concept, these concepts should be normalized beforehand and represented by the same embedding. With this in mind, Chen et al. [13] proposed a method to compute word embeddings of biological concepts using approximately 30 million PubMed abstracts in which the mentions of these concepts had been previously normalized using PubTator [14]. PubTator is an online tool that supports the automatic annotation of biomedical text and aims at extracting specific information efficiently. Particularly, it can identify the mentions of biological concepts and entities within text and classify them into appropriate categories (e.g., diseases, genes, drugs). They trained four embedding models, including skip-gram, and named their embeddings BioConceptVec, and assessed the usefulness of BioConceptVec in two ways: intrinsic and extrinsic evaluations. For intrinsic evaluations, they identified related genes based on drug-gene and gene-gene interactions using cosine similarity of the embeddings. For extrinsic evaluations, they performed protein-protein interaction prediction and drug-drug interaction extraction using neural network classifiers with the embeddings. However, the embeddings’ performance on solving the analogy tasks is yet to be explored.
In this study, we consider analogy tasks in biology using word embeddings. We trained skip-gram embeddings similarly to BioConceptVec, aiming to compare them with BioConceptVec embeddings and to develop embeddings from datasets divided by year in an extended experiment. To evaluate the performance of analogy computation, we focus on predicting target genes from a given drug. These target genes are associated with proteins that the drug interacts with, and these interactions are called drug-target interactions (DTIs) [15, 16]. In this paper, we will refer to these connections as drug-gene relations. Our research aims to show that embeddings learned from biological text data contain information about drug-gene relations. In the example of $\mathbf{u}_{\mathrm{\emph{king}}}-\mathbf{u}_{\mathrm{\emph{man}}}+\mathbf{u}_{\mathrm{\emph{woman}}}\approx\mathbf{u}_{\mathrm{\emph{queen}}}$, the vector differences $\mathbf{u}_{\mathrm{\emph{king}}}-\mathbf{u}_{\mathrm{\emph{man}}}$ and $\mathbf{u}_{\mathrm{\emph{queen}}}-\mathbf{u}_{\mathrm{\emph{woman}}}$ represent the royalty relation. However, there are multiple drug-gene pairs with drug-gene relations, and a single drug often has multiple target genes. Therefore, we calculate the vector differences between the embeddings of each drug and its target genes and average these vector differences to define a mean vector representing the relation.
To evaluate the performance in predicting drug-gene relations, we use data derived from KEGG [17] as the ground truth. We first consider analogy tasks in the global setting where all drugs and genes are included and demonstrate high performance in solving these tasks. Next, to consider more detailed analogy tasks, we use information common to both drugs and genes. Specifically, we focus on biological pathways to categorize drugs and genes. To do this, we use a list of human pathways from the KEGG API. We group all drugs and genes that are associated with the same pathway into a single category and define vectors representing drug-gene relations for each pathway. In this pathway-wise setting, we demonstrate that using these vectors in the analogy computation can improve its performance. Finally, as an application of analogy computation, we investigate the potential of our approach to predict unknown drug-gene relations. For this purpose, we divide the vocabulary by year to distinguish between known and unknown drug-gene relations. We then train skip-gram embeddings using PubMed abstracts published before the specified year and redefine vectors representing the known relations to predict unknown relations. The experimental results show that our approach can predict unknown relations to a certain extent.
The structure of this paper is as follows. First, we explain related work and detail our approach in this study. In the following sections, we perform experiments for each setting and evaluate the performance using metrics such as top-1 accuracy. Finally, we discuss specific prediction results and then conclude.
Related work
[…]
In the biomedical field, capturing the relations between biomedical concepts such as drug-drug and protein-protein interaction is an important issue. Word embeddings trained on biomedical domain corpora have been used to predict drug-drug interactions for new drugs [19] and to construct networks of gene-gene interactions [20]. Also, neural networks have been used in a number of studies to predict these relations [21, 22, 23, 24, 25].
Technologies developed in NLP, not limited to word embedding, have been applied to a wide variety of problems in the biomedical domain: Müller et al. [26] have created an online literature search and curation system using an “ontology dictionary” obtained by text mining, and Friedman et al. [27] have used syntactic parsing to extract and structure information about cellular pathways from biological literature. Yeganova et al. [28] normalize synonyms in the biomedical literature by word embedding similarity. Furthermore, Du et al. [29] applied the word embedding algorithm to gene co-expression across the transcriptome to compute vector representations of genes. In recent years, there has also been intensive research on fine-tuning BERT [4] and other pre-trained language models to the biomedical domain [10, 30, 31].
Methods
(a) Predicting queen from woman
(b) Predicting genes from a drug
Figure 1: Analogy computations by adding the relation vector. (a) Example of a basic analogy task. The question “If man corresponds to king, what does woman correspond to?” is solved by adding the relation vector $\mathbf{v}_{\emph{royalty}}$ to $\mathbf{u}_{\emph{woman}}$. (b) Example of an analogy task in setting G. The target genes of the drug $d$ are $g_{1}$, $g_{2}$, and $g_{3}$, predicted by adding the relation estimator $\hat{\mathbf{v}}$ to the drug embedding $\mathbf{u}_{d}$.
Analogy tasks
[…]
$\mathbf{u}_{\mathrm{\emph{king}}}-\mathbf{u}_{\mathrm{\emph{man}}}+\mathbf{u}_{\mathrm{\emph{woman}}}\approx\mathbf{u}_{\mathrm{\emph{queen}}}.$ (1)
This analogy computation is illustrated in Fig. 1. Here, the vector $\mathbf{v}_{\mathrm{\emph{royalty}}}$ represents the royalty relation between man and king, and is defined as the vector difference between $\mathbf{u}_{\mathrm{\emph{king}}}$ and $\mathbf{u}_{\mathrm{\emph{man}}}$:
$\mathbf{v}_{\mathrm{\emph{royalty}}}:=\mathbf{u}_{\mathrm{\emph{king}}}-\mathbf{u}_{\mathrm{\emph{man}}}.$ (2)
Adding the royalty relation vector $\mathbf{v}_{\mathrm{\emph{royalty}}}$ to $\mathbf{u}_{\mathrm{\emph{woman}}}$ then yields $\mathbf{u}_{\mathrm{\emph{queen}}}$:
$\mathbf{u}_{\mathrm{\emph{woman}}}+\mathbf{v}_{\mathrm{\emph{royalty}}}\approx\mathbf{u}_{\mathrm{\emph{queen}}}.$ (3)
Analogy tasks for drug-gene pairs
In this section, based on the basic analogy task from the previous section, we explain analogy tasks for predicting target genes from a drug. First, we consider the global setting, where all drugs and genes in the vocabulary set are used. Then, we consider the pathway-wise setting, where drugs and genes are categorized based on biological pathways.
Global setting
(a) Randomly sampled drug-gene pairs.
(b) Drugs and genes categorized in the ErbB signaling pathway.
Figure 2: 2-d visualization of embeddings using PCA. Drugs and genes are shown in blue and orange, respectively. Solid lines represent the relations. The symbols $\bigstar$ and $\blacksquare$ represent the mean embeddings of the drugs and genes, respectively. The direction of the dashed line connecting these two symbols can be considered as the drug-gene relation vector. (a) Randomly sampled $200$ drug-gene pairs from $\mathcal{R}$. The mean embeddings are computed by Eq. (9). (b) Drugs $d\in\mathcal{D}_{p}$ and genes $g\in\mathcal{G}_{p}$ for $p=$ErbB signaling pathway. Those that have drug-gene relations are shown in boxes. The mean embeddings are computed by Eq. (16).
As a first step, based on the basic analogy task, we consider analogy tasks in the global setting, where all drugs and genes are used. We define $\mathcal{D}\subset\mathcal{V}$ and $\mathcal{G}\subset\mathcal{V}$ as the sets of all drugs and genes, respectively. We also define $\mathcal{R}\subset\mathcal{D}\times\mathcal{G}$ as the set of drug-gene pairs with drug-gene relations. Thus, if a drug $d\in\mathcal{D}$ and a gene $g\in\mathcal{G}$ have a drug-gene relation, the pair $(d,g)$ is in $\mathcal{R}$. For illustrating the positions of drugs and genes in the $100$-dimensional BioConceptVec skip-gram embeddings, we sampled $200$ drug-gene pairs in $\mathcal{R}$ and plotted them in Fig. 2 using Principal Component Analysis (PCA). The vector difference between the mean embeddings of drugs and genes points in roughly the same direction as the vector differences between the embeddings of each sampled drug and its target gene.
To apply the analogy computation of Eq. (3) to drug-gene pairs $(d,g)\in\mathcal{R}$, we consider the analogy tasks for predicting the gene $g$ from the drug $d$. Using the vector $\mathbf{v}$ representing the drug-gene relation, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}$ to $\mathbf{u}_{d}$ as follows:
$\mathbf{u}_{d}+\mathbf{v}\approx\mathbf{u}_{g}.$ (4)
Thus, we need methods for estimating the relation vector $\mathbf{v}$. In Eq. (2), the royalty relation vector $\mathbf{v}_{\text{\emph{royalty}}}$ is calculated from the embeddings of the given pair (man, king). Therefore, we estimate the relation vector $\mathbf{v}$ from the embeddings of genes and drugs. The following equation gives the vector difference between the mean vectors of $\mathcal{D}$ and $\mathcal{G}$ as a naive estimator of $\mathbf{v}$:
$\hat{\mathbf{v}}_{\text{naive}}:=\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{D}}\{\mathbf{u}_{d}\},$ (5)
$\mathrm{E}_{\mathcal{D}}\{\mathbf{u}_{d}\}=\frac{1}{|\mathcal{D}|}\sum_{d\in\mathcal{D}}\mathbf{u}_{d},\quad\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\}=\frac{1}{|\mathcal{G}|}\sum_{g\in\mathcal{G}}\mathbf{u}_{g}.$ (6)
In Eq. (6), $\mathrm{E}_{\mathcal{D}}\{\cdot\}$ and $\mathrm{E}_{\mathcal{G}}\{\cdot\}$ represent the sample means over the set of all drugs $\mathcal{D}$ and the set of all genes $\mathcal{G}$, respectively. However, the definition of $\hat{\mathbf{v}}_{\text{naive}}$ in Eq. (5) includes the embeddings of unrelated genes and drugs. Therefore, a better estimator may be defined by using only the pairs $(d,g)\in\mathcal{R}$. We consider the estimator $\hat{\mathbf{v}}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}$ as follows:
$\hat{\mathbf{v}}:=\mathrm{E}_{\mathcal{\mathcal{R}}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{|\mathcal{R}|}\sum_{(d,g)\in\mathcal{R}}(\mathbf{u}_{g}-\mathbf{u}_{d}),$ (7)
where $\mathrm{E}_{\mathcal{R}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}$. For easy comparison with Eq. (5), Eq. (7) is rewritten as the difference of mean vectors:
$\hat{\mathbf{v}}=\mathrm{E}_{\mathcal{R}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}}\{\mathbf{u}_{d}\},$ (8)
$\mathrm{E}_{\mathcal{R}}\{\mathbf{u}_{g}\}=\frac{1}{|\mathcal{R}|}\sum_{(d,g)\in\mathcal{R}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{\mathcal{R}}}\{\mathbf{u}_{d}\}=\frac{1}{|\mathcal{R}|}\sum_{(d,g)\in\mathcal{R}}\mathbf{u}_{d}.$ (9)
To measure the performance of the estimator $\hat{\mathbf{v}}$ in Eq. (7), we prepare the evaluation of the analogy tasks. We define $D$ and $G$ as the sets of drugs and genes contained in $\mathcal{R}$, respectively. Specifically, we define $D:=\pi_{\mathcal{D}}(\mathcal{R})$ as the set of drugs $d$ such that $(d,g)\in\mathcal{R}$ for some genes $g$, and $G:=\pi_{\mathcal{G}}(\mathcal{R})$ as the set of genes $g$ such that $(d,g)\in\mathcal{R}$ for some drugs $d$, where the projection operations $\pi_{\mathcal{D}}$ and $\pi_{\mathcal{G}}$ are defined as
$\pi_{\mathcal{D}}:\mathcal{R}\mapsto\{d\mid(d,g)\in\mathcal{R}\}\subset\mathcal{D},\quad\pi_{\mathcal{G}}:\mathcal{R}\mapsto\{g\mid(d,g)\in\mathcal{R}\}\subset\mathcal{G}.$ (10)
We also define $[d]\subset\mathcal{G}$ as the set of genes that have drug-gene relations with a drug $d\in D$, and $[g]\subset\mathcal{D}$ as the set of drugs that have drug-gene relations with a gene $g\in G$. These are formally defined as
$[d]:=\{g\mid(d,g)\in\mathcal{R}\}\subset\mathcal{G},\quad[g]:=\{d\mid(d,g)\in\mathcal{R}\}\subset\mathcal{D}.$ (11)
Given the above, we perform the analogy computation in the following setting.
Setting G.
In the analogy tasks, the set of answer genes for a query drug $d\in D$ is $[d]$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}},\mathbf{u}_{g})$ and if $\hat{g}_{d}\in[d]$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th ranked $g\in\mathcal{G}$ based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}},\mathbf{u}_{g})$. For the top-$k$ accuracy, if any of the top $k$ predictions $\hat{g}_{d}^{(1)},\ldots,\hat{g}_{d}^{(k)}\in[d]$, then the prediction is considered correct.
Unlike in the basic analogy task, there may be multiple target genes for a single drug. While the basic analogy task is one-to-one, we address one-to-many analogy tasks in this study, where a single source may correspond to multiple targets [32]. Analogy computation for setting G is illustrated in Fig. 1, where the set of answer genes $[d]$ consists of $g_{1}$, $g_{2}$, and $g_{3}$. If the predicted gene $\hat{g}_{d}$ is one of these three genes, then the prediction is considered correct. Note that the analogy tasks for predicting drugs from a query gene can also be defined similarly. See Supplementary Information 1.3 for details.
Pathway-wise setting
Set Size Symbol Elements of the set
$\mathcal{D}_{p}$ 19 $d$ Bosutinib, Canertinib dihydrochloride, Dabrafenib, Dacomitinib, Masoprocol, Neratinib, Nilotinib, Osimertinib, Pelitinib, Pimasertib, Poziotinib, Rociletinib, Saracatinib, Selumetinib, Temsirolimus, Tozasertib, Trametinib, Vandetanib, Varlitinib
$\mathcal{G}_{p}$ 84 $g$ ABL1, ABL2, AKT1, AKT2, AKT3, ARAF, AREG, BAD, BRAF, BTC, CAMK2A, CAMK2B, CAMK2D, CAMK2G, CBL, CBLB, CDKN1A, CDKN1B, CRK, CRKL, EGF, EGFR, EIF4EBP1, ELK1, ERBB2, ERBB3, ERBB4, EREG, GAB1, GRB2, GSK3B, HBEGF, HRAS, JUN, KRAS, MAP2K1, MAP2K2, MAP2K4, MAP2K7, MAPK1, MAPK10, MAPK3, MAPK8, MAPK9, MTOR, MYC, NCK1, NCK2, NRAS, NRG1, NRG2, NRG3, NRG4, PAK1, PAK2, PAK3, PAK4, PAK5, PAK6, PIK3CA, PIK3CB, PIK3CD, PIK3R1, PIK3R2, PIK3R3, PLCG1, PLCG2, PRKCA, PRKCB, PRKCG, PTK2, RAF1, RPS6KB1, RPS6KB2, SHC1, SHC2, SHC3, SHC4, SOS1, SOS2, SRC, STAT5A, STAT5B, TGFA
$\mathcal{R}_{p}$ 34 $(d,g)$ (Bosutinib, ABL1), (Bosutinib, SRC), (Canertinib dihydrochloride, EGFR), (Canertinib dihydrochloride, ERBB2), (Canertinib dihydrochloride, ERBB3), (Canertinib dihydrochloride, ERBB4), (Dabrafenib, BRAF), (Dacomitinib, EGFR), (Dacomitinib, ERBB2), (Dacomitinib, ERBB4), (Masoprocol, EGFR), (Neratinib, EGFR), (Neratinib, ERBB2), (Neratinib, ERBB4), (Nilotinib, ABL1), (Osimertinib, EGFR), (Pelitinib, EGFR), (Pimasertib, MAP2K1), (Pimasertib, MAP2K2), (Poziotinib, EGFR), (Poziotinib, ERBB2), (Rociletinib, EGFR), (Saracatinib, ABL1), (Saracatinib, SRC), (Selumetinib, MAP2K1), (Selumetinib, MAP2K2), (Temsirolimus, MTOR), (Tozasertib, ABL1), (Trametinib, MAP2K1), (Trametinib, MAP2K2), (Vandetanib, EGFR), (Varlitinib, EGFR), (Varlitinib, ERBB2), (Varlitinib, ERBB4)
Table 1: $\mathcal{D}_{p}$, $\mathcal{G}_{p}$, and $\mathcal{R}_{p}$ for the pathway $p$, where $p$ is the ErbB signaling pathway. Note that we only use concepts that exist in BioConceptVec and can be converted to names, so they do not exactly match the actual drugs and genes categorized in the biological pathway. In addition, we have changed the notation of some drugs from compounds to more readable names.
As a next step, based on the analogy tasks in setting G, we consider more detailed analogy tasks. To do this, we consider the analogy tasks in the pathway-wise setting, where drugs and genes are categorized using biological pathways. We define $\mathcal{P}$ as the set of pathways $p$, and $\mathcal{D}_{p}\subset\mathcal{D}$ and $\mathcal{G}_{p}\subset\mathcal{G}$ as the sets of drugs and genes categorized in each pathway $p\in\mathcal{P}$, respectively. We then restrict the set $\mathcal{R}$ to each pathway $p$ and define the subset of $\mathcal{R}$ as:
$\mathcal{R}_{p}:=\{(d,g)\in\mathcal{R}\mid d\in\mathcal{D}_{p},g\in\mathcal{G}_{p}\}\subset\mathcal{D}_{p}\times\mathcal{G}_{p}.$ (12)
A specific example of these sets for the ErbB signaling pathway is shown in Table 1, and their BioConceptVec skip-gram embeddings are illustrated in Fig. 2. The vector difference between the mean embeddings of drugs and genes roughly points in the same direction as the vector differences between the embeddings of each drug and its target gene in $\mathcal{R}_{p}$, although such a two-dimensional illustration should be interpreted with caution.
For drug-gene pairs $(d,g)\in\mathcal{R}_{p}$, we consider the analogy tasks for predicting the target genes $g$ from a drug $d$. To solve these analogy tasks, we use the relation vector $\mathbf{v}_{p}$, which represents the relation between drugs and target genes categorized in the same pathway $p$. We predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}_{p}$ to $\mathbf{u}_{d}$, expecting that
$\mathbf{u}_{d}+\mathbf{v}_{p}\approx\mathbf{u}_{g}.$ (13)
Equation (13) corresponds to Eq. (4). Therefore, similar to the estimator $\hat{\mathbf{v}}$ in Eq. (7), we define an estimator $\hat{\mathbf{v}}_{p}$ for the relation vector $\mathbf{v}_{p}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}_{p}$:
$\hat{\mathbf{v}}_{p}:=\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{|\mathcal{R}_{p}|}\sum_{(d,g)\in\mathcal{R}_{p}}(\mathbf{u}_{g}-\mathbf{u}_{d}),$ (14)
where $\mathrm{E}_{\mathcal{R}_{p}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}_{p}$. For easy comparison with Eq. (5), Eq. (14) is rewritten as the difference of mean vectors:
$\hat{\mathbf{v}}_{p}=\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{d}\},$ (15)
$\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{g}\}=\frac{1}{|\mathcal{R}_{p}|}\sum_{(d,g)\in\mathcal{R}_{p}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{d}\}=\frac{1}{|\mathcal{R}_{p}|}\sum_{(d,g)\in\mathcal{R}_{p}}\mathbf{u}_{d}.$ (16)
To measure the performance of the estimator $\hat{\mathbf{v}}_{p}$, we prepare the evaluation of the analogy tasks. Similar to $D$ and $G$, we define $D_{p}$ and $G_{p}$ as the sets of drugs and genes contained in $\mathcal{R}_{p}$, respectively. Using the operations of Eq. (10), we define $D_{p}:=\pi_{\mathcal{D}_{p}}(\mathcal{R}_{p})\subset\mathcal{D}_{p}$ as the set of drugs $d$ such that $(d,g)\in\mathcal{R}_{p}$ for some genes $g$, and $G_{p}:=\pi_{\mathcal{G}_{p}}(\mathcal{R}_{p})\subset\mathcal{G}_{p}$ as the set of genes $g$ such that $(d,g)\in\mathcal{R}_{p}$ for some drugs $d$. We also define $[d]_{p}\subset\mathcal{G}_{p}$ as the set of genes that have drug-gene relations with a drug $d\in D_{p}$, and $[g]_{p}\subset\mathcal{D}_{p}$ as the set of drugs that have drug-gene relations with a gene $g\in G_{p}$. Similar to Eq. (11), these sets are formally defined as
$[d]_{p}:=\{g\mid(d,g)\in\mathcal{R}_{p}\}\subset\mathcal{G}_{p},\quad[g]_{p}:=\{d\mid(d,g)\in\mathcal{R}_{p}\}\subset\mathcal{D}_{p}.$ (17)
Given the above, we perform the analogy computation in the following two settings.
Setting P1.
For the target genes that have drug-gene relations with a drug $d$, only genes categorized in the same pathway $p$ as the drug $d$ are considered correct. In other words, for a query drug $d\in D_{p}$, the set of answer genes is $[d]_{p}$. The search space is the set of all genes $\mathcal{G}$, not limited to $\mathcal{G}_{p}$, the set of genes categorized in the pathway $p$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]_{p}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th ranked $g\in\mathcal{G}$ based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p},\mathbf{u}_{g})$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]_{p}$, then the prediction is considered correct.
Setting P2.
The gene predictions $\hat{g}_{d}$ and $\hat{g}_{d}^{(k)}$ are defined exactly the same as those in setting P1, but the answer genes are defined the same as in setting G. That is, for the target genes that have drug-gene relations with a drug $d$, genes are considered correct regardless of whether they are categorized in the same pathway $p$ as the drug $d$ or not. In other words, for a query drug $d\in D$, the set of answer genes is $[d]$, and the prediction is considered correct if $\hat{g}_{d}\in[d]$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]$, then the prediction is considered correct. Note that the experiment is performed for $d\in\mathcal{D}_{p}\cap D$ for each $p$.
Fig. S1 in Supplementary Information 1.1.1 shows the differences between settings P1 and P2 using specific examples. Table S2 in Supplementary Information 1.3 summarizes the queries, answer sets, and search spaces in settings G, P1, and P2.
Analogy tasks for drug-gene pairs by year
In this section, based on analogy tasks for drug-gene pairs, we explain analogy tasks in the setting where datasets are divided by year. In this setting, we can separate the drug-gene relations into known or unknown relations based on the chronological order they appear within the PubMed abstracts. By training skip-gram embeddings with this dataset, we can test if embeddings of known relations have the ability to predict unknwon relations. Similar to analogy tasks for drug-gene pairs, we first consider the global setting by year, where all drugs and genes are used in datasets divided by year. Then we consider the pathway-wise setting by year, where drugs and genes are categorized based on pathways in datasets divided by year.
Global setting by year
As a first step, based on analogy tasks for drug-gene pairs, we consider analogy tasks in the global setting, where all drugs and genes are used in datasets divided by year. Specifically, drug-gene relations that appeared up to year $y$ are considered known, while those that appeared after year $y$ are considered unknown. We then train skip-gram embeddings using PubMed abstracts published before the specified year, redefine vectors representing known relations, and use these vectors to predict unknown relations. In preparation, we define $y_{d}$ as the year when a drug $d$ first appeared in a PubMed abstract, and $y_{g}$ as the year when a gene $g$ first appeared in a PubMed abstract. In addition, we define $y_{(d,g)}$ as the year when both drug $d$ and gene $g$ first appeared together in a PubMed abstract, and use $y_{(d,g)}$ as a substitute for the year when the relation $(d,g)$ was first identified. By definition, $\max\{y_{d},y_{g}\}\leq y_{(d,g)}$ holds.
Consider a fixed year $y$. When learning embeddings using PubMed abstracts up to year $y$ as training data, we define $\mathcal{D}^{y}:=\{d\in\mathcal{D}\mid y_{d}\leq y\}$ and $\mathcal{G}^{y}:=\{g\in\mathcal{G}\mid y_{g}\leq y\}$ as the sets of drugs and genes that appeared up to year $y$, respectively. The set of drug-gene pairs that have drug-gene relations and whose drugs and genes appeared up to year $y$ is expressed as
$\mathcal{R}^{y}:=\{(d,g)\in\mathcal{R}\mid d\in\mathcal{D}^{y},g\in\mathcal{G}^{y}\}\subset\mathcal{D}^{y}\times\mathcal{G}^{y}.$ (18)
Since the vocabulary set of the trained embeddings includes both $d$ and $g$ for $(d,g)\in\mathcal{R}^{y}$, we may discuss this relation $(d,g)$. The relation $(d,g)$ is interpreted as either known by year $y$ if $y_{(d,g)}\leq y$ or unknown by year $y$ if $y<y_{(d,g)}$.
We divide $\mathcal{R}^{y}$ into two subsets based on whether $y_{(d,g)}\leq y$ or $y<y_{(d,g)}$. Let us define two intervals, $L_{y}:=(-\infty,y]$ and $U_{y}:=(y,\infty)$. Considering the set for an interval $I$, consisting of all $(d,g)\in\mathcal{R}^{y}$ such that $y_{(d,g)}\in I$, i.e.,
$\mathcal{R}^{y\mid I}:=\{(d,g)\in\mathcal{R}^{y}\mid y_{(d,g)}\in I\}\subset\mathcal{R}^{y},$ (19)
the set of “known” relations is expressed as $\mathcal{R}^{y\mid L_{y}}$, and the set of “unknown” relations is expressed as $\mathcal{R}^{y\mid U_{y}}$. By definition, $\mathcal{R}^{y\mid L_{y}}\cap\mathcal{R}^{y\mid U_{y}}=\emptyset$ and $\mathcal{R}^{y\mid L_{y}}\cup\mathcal{R}^{y\mid U_{y}}=\mathcal{R}^{y\mid(-\infty,\infty)}=\mathcal{R}^{y}$.
In analogy tasks, we use “known” $\mathcal{R}^{y\mid L_{y}}$ and then predict the target genes $g$ from a drug $d$ for $(d,g)$ in “unknown” $\mathcal{R}^{y\mid U_{y}}$. Using the vector $\mathbf{v}^{y}$, which represents the drug-gene relations and is derived from $\mathcal{R}^{y\mid L_{y}}$, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}^{y}$ to $\mathbf{u}_{d}$:
$\mathbf{u}_{d}+\mathbf{v}^{y}\approx\mathbf{u}_{g}.$ (20)
Equation (20) corresponds to Eq. (4). Therefore, similar to the estimator $\hat{\mathbf{v}}$ in Eq. (7), we define an estimator $\hat{\mathbf{v}}^{y}$ for the relation vector $\mathbf{v}^{y}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}^{y\mid L_{y}}$:
$\hat{\mathbf{v}}^{y}:=\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}^{y\mid L_{y}}\right|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}(\mathbf{u}_{g}-\mathbf{u}_{d}),$ (21)
where $\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}^{y\mid L_{y}}$. For easy comparison with Eq. (5), Eq. (21) is rewritten as the difference of mean vectors:
$\hat{\mathbf{v}}^{y}=\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\},$ (22)
$\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{d}.$ (23)
Given the above, similar to setting G, we perform the analogy tasks in the following two settings.
Setting Y1.
For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct. In other words, only new discoveries are counted as correct.
Setting Y2.
For the target genes that have drug-gene relations with a drug $d$, these genes are considered correct regardless of whether their relations appeared up to or after year $y$. In other words, rediscoveries of known relations are also counted as correct.
See Supplementary Information 1.2.1 for more details on these settings. Note that $y_{d}$, $y_{g}$, and $y_{(d,g)}$ are defined based on the year the drugs, genes, and drug-gene relations appeared in PubMed abstracts. Thus, they do not fully correspond to their actual years of discovery, and we only consider the analogy tasks in these hypothetical settings.
Pathway-wise setting by year
As a next step, based on the analogy tasks in settings P1, P2, Y1, and Y2, we consider more detailed analogy tasks. To do this, we consider the analogy tasks in the pathway-wise setting by year, where drugs and genes are categorized based on pathways in datasets divided by year.
Similar to settings P1, P2, Y1, and Y2, we perform the analogy tasks in the following four settings.
Setting P1Y1.
For the target genes that have drug-gene relations with a drug $d$, only genes categorized in the same pathway $p$ as $d$, and whose relations appeared after year $y$, are considered correct.
Setting P1Y2.
For the target genes that have drug-gene relations with a drug $d$, only genes categorized in the same pathway $p$ as the drug $d$ are considered correct, regardless of whether their relations appeared up to or after year $y$.
Setting P2Y1.
For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct, regardless of whether they are categorized in the same pathway $p$ as the drug $d$ or not.
Setting P2Y2.
For the target genes that have drug-gene relations with a drug $d$, these genes are considered correct regardless of whether they are categorized in the same pathway or not, and whether their relations appeared up to or after year $y$.
See Supplementary Information 1.2.2 for details on the analogy tasks and these settings.
Embeddings
BioConceptVec Ours
$|\mathcal{D}|$ 117282 28284
$|\mathcal{G}|$ 144584 51057
$|\mathcal{R}|$ 6645 5968
$|D|$ 2262 1980
$|G|$ 664 634
$\mathrm{E}_{d\in D}\{|[d]|\}$ 2.938 3.014
$\mathrm{E}_{g\in G}\{|[g]|\}$ 10.008 9.413
Table 2: Statistics for setting G.
BioConceptVec [13] provides four pretrained 100-dimensional word embeddings: CBOW [7], skip-gram [7, 8], GloVe [33], and fastText [34]. Since CBOW is a simpler model than skip-gram and skip-gram performs better in analogy tasks compared to GloVe [9], we used BioConceptVec skip-gram embeddings for our experiments. Note that fastText can essentially be considered as a skip-gram using $n$-grams. To complement the pretrained BioConceptVec embeddings, we further trained 300-dimensional skip-gram embeddings on the publicly available PubMed abstracts. To train our embeddings, we used PubTator [14] to normalize six major biological concepts (genes, mutations, diseases, chemicals, cell lines, and species) in PubMed abstracts, and then tokenized them using NLTK [35]. Note that since widely used embeddings such as word2vec and GloVe typically have 300 dimensions, we set the dimensions of our embeddings to 300 instead of the original 100. The hyperparameters used to train our embeddings are shown in Table S4 in Supplementary Information 1.4.
For BioConceptVec and our skip-gram embeddings, Table 2 shows some basic statistics for setting G. Due to the difference in training data size and minimal word occurrence, the sizes of certain sets such as $|\mathcal{D}|$ and $|\mathcal{G}|$ differ significantly between BioConceptVec and our skip-gram embeddings, but the sizes of other sets show somewhat similar trends. Table S1 in Supplementary Information 1.1.1 shows the statistics for settings P1 and P2. The hyperparameters used to train our skip-gram are shown in Table S4 in Supplementary Information 1.4.
Datasets
Corpus
Following Chen et al. [13], we used PubMed (https://pubmed.ncbi.nlm.nih.gov/) abstracts to train our skip-gram embeddings. We used about 35 million abstracts up to the year 2023, while they used about 30 million abstracts.
Drug-gene relations
For drug-gene relations, we used publicly available data from AsuratDB [36], which collects information from various databases including the KEGG [17] database. KEGG (Kyoto Encyclopedia of Genes and Genomes) is a comprehensive database system that integrates a wide range of bioinformatics information such as genomics, chemical reactions, and biological pathways.
Biological Pathways
We obtained a list of human pathways from the KEGG API (https://rest.kegg.jp/list/pathway/hsa) for use in our experiments. For each pathway, we again used the KEGG API to define sets of drugs and genes. To avoid oversimplification of analogy tasks, we excluded the pathways where only one type of drug or gene had drug-gene relations.
For more details, see Data availability section and Supplementary Information 1.5.
Baseline
To evaluate the performance of predicting target genes by adding the relation vector to a drug embedding, we compare it to a simple baseline method. As a baseline, we randomly select one gene from the search space for each query. We refer to this baseline as the random baseline. In the random baseline, we repeated the experiments 10 times and used the average score as the final result.
Evaluation metrics
As evaluation metrics, we use the top-$k$ accuracy explained in each setting, especially the top-1 and top-10 accuracies. We also use Mean Reciprocal Rank (MRR) as another evaluation metric. MRR is a statistical measure to evaluate search performance, expressed as the sum of the reciprocals of the ranks at which the predicted results first appear in the answer set.
When evaluating performance using top-1 accuracy, top-10 accuracy, and MRR, the embeddings of all genes $\mathcal{G}$ (or $\mathcal{G}^{y}$ in settings by year) are centered using the mean embedding $\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\}$ in Eq. (6) (or $\mathrm{E}_{\mathcal{G}^{y}}\{\mathbf{u}_{g}\}$). Since cosine similarity is affected by the origin, centering mitigates this effect. In the case of setting G, for example, we actually calculate the value of $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}-\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\},\mathbf{u}_{g}-\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\})$ instead of $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}},\mathbf{u}_{g})$.
Results
Metric
Embeddings Setting Method Top1 Top10 MRR
BioConceptVec G Random 0.000 0.000 0.000
$\hat{\mathbf{v}}$ 0.304 0.645 0.419
P1 Random 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}$ 0.499 0.790 0.602
P2 Random 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}$ 0.522 0.807 0.624
Ours G Random 0.000 0.000 0.001
$\hat{\mathbf{v}}$ 0.300 0.686 0.426
P1 Random 0.000 0.001 0.001
$\hat{\mathbf{v}}_{p}$ 0.589 0.862 0.685
P2 Random 0.000 0.001 0.001
$\hat{\mathbf{v}}_{p}$ 0.600 0.880 0.700
Table 3: Gene prediction performance in settings G, P1, and P2.
(a) Y1
(b) Y2
(c) P1Y1
(d) P1Y2
Figure 3: Gene prediction performance and the number of queries for each year in settings Y1, Y2, P1Y1, and P1Y2.
Gene prediction performance in settings G, P1, and P2.
Table 3 shows the results of experiments with BioConceptVec and our skip-gram embeddings in settings G, P1, and P2. In setting G, the prediction by simply adding the global relation vector $\hat{\mathbf{v}}$ resulted in a top-1 accuracy of about 0.3, a top-10 accuracy of over 0.6, and an MRR also over 0.4. These results show that the embeddings have the ability to interpret the drug-gene relations. Similar to the basic analogy task, these relations were not explicitly provided during the training of the embeddings. Previous research [37] has shown that skip-gram achieves a top-1 accuracy of 0.56 on the Microsoft Research Syntactic Analogies Dataset [38]. Given this, BioConceptVec and our skip-gram embeddings in the analogy tasks for drug-gene pairs show comparable performance to skip-gram in the basic analogy tasks.
In settings P1 and P2, prediction by adding the pathway-wise relation vector $\hat{\mathbf{v}}_{p}$ showed better performance than in setting G. For example, our skip-gram embeddings achieved a top-1 accuracy of over 0.5 in both settings P1 and P2. This is probably because the queries are specific drugs categorized in some pathways, and we use the pathway information to calculate the relation vectors. Since the analogy tasks are considered for each pathway, these settings are also likely to make the tasks easier than those in setting G.
In contrast, the random baseline in settings G, P1, and P2 showed nearly zero scores for all evaluation metrics, suggesting that random prediction is nearly impossible in these analogy tasks. This also confirms the superior performance of prediction by adding relation vectors.
In addition, our skip-gram outperformed BioConceptVec except for the top-1 accuracy in setting G. This is probably because BioConceptVec has a dimensionality of 100, while our skip-gram has a dimensionality of 300. Note that, as shown in Table 2, the settings of the analogy tasks are not identical for these two embeddings.
For the results in Table 3, we used the better estimators in equations (7) and (14), not the naive estimators in equations (5) and (S1), and calculated the cosine similarities after centering the embeddings of all genes $\mathcal{G}$ (or $\mathcal{G}^{y}$). The results of the naive estimators and the centering ablation study are shown in Table S5 in Supplementary Information 2.1.1. The results of the analogy tasks for predicting drugs from a query gene are shown in Table S6 in Supplementary Information 2.1.2.
Gene prediction performance in settings Y1 and Y2
Figures 3(a) and 3(b) show the results of experiments with our skip-gram embeddings in settings Y1 and Y2. BioConceptVec embeddings are not appropriate for the settings by year, since they are pretrained on the entire dataset without year-specific divisions.
In setting Y1, since only genes whose relations appeared after year $y$ are included in the answer set, the tasks are more challenging. Therefore, the evaluation metric scores are lower compared to those in setting Y2, where genes whose relations appeared up to year $y$ are also included in the answer set. Yet, in setting Y1, the top-1 accuracy from $y=1985$ to $y=2015$ is about 0.1, the top-10 accuracy is over 0.3, and the MRR is about 0.2. Although in setting Y1 the relations to be predicted appeared after year $y$ and are not used to calculate $\hat{\mathbf{v}}^{y}$, these results show that adding $\hat{\mathbf{v}}^{y}$ to $\mathbf{u}_{d}$ can predict these relations. Setting Y2 is easier than setting Y1, and using the relation vectors results in high evaluation metric scores. For example, from $y=1985$ to $y=2015$, the top-1 accuracy reaches about 0.3, and the top-10 accuracy is about 0.7.
In setting Y1, the number of queries increases initially and then decreases over time, while in setting Y2, the number of queries increases steadily from older to more recent years. In addition, the size of the search space is small for older years (e.g., $|\mathcal{G}^{1970}|=353$) and increases for later years (e.g., $|\mathcal{G}^{2020}|=47509$). As a result, in both settings Y1 and Y2, the datasets for older years such as $y=1970$, $1975$, and $1980$ have fewer queries and smaller search spaces, resulting in unusually high evaluation metric scores. In setting Y1, the datasets for later years have fewer queries and larger search spaces, which tends to produce lower evaluation metric scores. In both settings, the evaluation metric scores for the random baseline are close to zero. For more detailed results for settings Y1 and Y2, see Table S8 in Supplementary Information 2.2.1.
Gene prediction performance in settings P1Y1, P1Y2, P2Y1 and P2Y2.
Figures 3(c) and 3(d) show the results of experiments with our skip-gram embeddings in settings P1Y1 and P1Y2. The results in settings P1Y1 and P1Y2 not only follow similar trends to those in settings Y1 and Y2 but also consistently show higher evaluation metric scores. The analogy tasks in settings P1Y1 and P1Y2 use detailed pathway information. Therefore, predicting the relations that appeared after year $y$ is easier in settings P1Y1 and P1Y2 than in settings Y1 and Y2. In addition, the results for the older years show extremely high evaluation metric scores in settings P1Y1 and P1Y2, similar to those in settings Y1 and Y2. The evaluation metric scores of the random baseline are close to zero.
For the results in settings P2Y1 and P2Y2, see Fig. S2 in Supplementary Information 2.2.2. They are very similar to those in settings P1Y1 and P1Y2. For detailed results for settings P1Y1, P1Y2, P2Y1, and P2Y2, see Table S10 in Supplementary Information 2.2.2.
Discussion
Gene Prediction
Drug Setting Top1 Top2 Top3 Top4 Top5 Top6 Top7 Top8 Top9 Top10 Answer Set
[…]
Table 4: Gene prediction results for the drugs categorized in the ErbB signaling pathway. Shown only for drugs where the size of $[d]$ is two or more. For each drug, the top ten predicted genes and their answer sets are shown in settings G, P1, and P2. Predicted genes included in the answer sets are underlined. Genes whose IDs could not be converted to gene names are indicated by “–”.
In this section, we investigate the predictions made by adding the relation vector in settings G, P1, and P2. For this purpose, we focus on the results using genes and drugs categorized in the ErbB signaling pathway as shown in Table 1 and Fig. 2. Furthermore, since our skip-gram embeddings showed better performance compared to the BioConceptVec skip-gram embeddings (Table 3), we adopted our embeddings for this analysis. $\mathcal{D}_{p}$, $\mathcal{G}_{p}$, and $\mathcal{R}_{p}$ of our embeddings are identical to those of the BioConceptVec embeddings in Table 1.
In Table 4, we list the predicted target genes for drugs that are categorized in the ErbB signaling pathway and whose answer set size in setting G is two or more. For all drugs meeting these criteria, at least one answer gene was included in the top 10 predicted target genes for either of the settings. In the following, we show that genes that were ranked high in the predictions, even though they were not in the answer set, can be interpreted biologically.
First, the target genes of Bosutinib are ABL1 and SRC, which are both known to be non-receptor tyrosine kinases (non-RTKs) [39]. Within the top 10 predicted target genes, TXK [40] and JAK2 [41] can also be categorized as non-RTKs, sharing some properties with ABL1 and SRC. This implies that the structural and biochemical similarities of these genes may have been reflected in their high-dimensional representations.
For Masoprocol, we were able to predict the correct target gene EGFR as the fifth prediction in settings P1 and P2. However, although EGFR is the sole target gene of Masoprocol according to the information deposited in KEGG, it is also reported to have inhibitory effects towards lipoxygenase activity [42]. Since ALOX5 is a gene that codes a lipoxygenase, treating this prediction as inaccurate should be considered incorrect. The diseases in which Masoprocol is used for its treatment include actinic keratosis [43], and it has been previously known that the mutation of TP53 is involved in the onset of this disease [44, 45]. Also, the gene that was ranked the highest in this setting, ALOX5, is reported to be one of the transcription targets of TP53 [46]. Taken together, this information suggests that TP53 is deeply related to the target gene and symptoms of Masoprocol, and therefore, although it may not be a direct target, it was highly ranked in the prediction.
When we set Poziotinib as the query, EGFR and ERBB2 were correctly ranked the highest in settings P1 and P2. However, in setting G, ROS1, EML4, and ALK were also found among the highly ranked targets. Since Poziotinib was initially developed as an effective drug to treat lung cancer with HER2 mutation [47], it can be interpreted that such context was reflected in the embeddings, as ROS1 and EML4-ALK fusion genes are also characteristic mutations and therapeutic targets in lung cancer [48, 49].
Selumetinib and Trametinib both target MAP2K1 and MAP2K2, and in both cases these genes were found within the top 10 predictions. Aside from these target genes, we observed BRAF and PI3KCA placed in higher rank among the predicted genes. Considering the history that these drugs were both developed for treating cancers with BRAF V600 mutation and that combination treatment with inhibition of the PI3K-AKT pathway have been explored [50, 51, 52], we assumed that such background was reflected in these results.
As we explored in this section, it was suggested that the high-dimensional representations of genes and drugs calculated from biomedical texts not only capture the simple drug-gene relations we sought to predict in this study, but also integrate broader higher-order information. In this specific analysis, predicted genes are closely related to the genes or symptoms that have direct molecular interactions with the target genes, or those that are alternative therapeutic targets for the target disease.
Conclusions
In this study, we used the embeddings learned from biological texts and performed analogy tasks for predicting drug-gene relations. We defined vectors representing these relations and showed that these vectors can accurately predict the target genes for given drugs. Furthermore, we categorized drugs and genes based on biological pathways and then defined vectors representing drug-gene relations for each pathway. Analogy computation using these vectors showed more performance improvement. In addition, we divided datasets by year and redefined the vectors representing known relations to predict unknown relations. The experimental results show that our approach can predict unknown relations to some extent.
Data availability
Corpus
PubMed abstracts is available at https://ftp.ncbi.nlm.nih.gov/pubmed/.
Embeddings
word2vec is available at https://code.google.com/archive/p/word2vec/. BioConceptVec [13] is available at https://github.com/ncbi/BioConceptVec.
Drug-gene relations
The published data from AsuratDB [36], which is based on the KEGG database [53], is available at https://github.com/keita-iida/ASURATDB/blob/main/genes2bioterm/20221102_human_KEGG_drug.rda. The drug-gene relations used in this study are published with our code.
Pathways
A list of pathways is available at https://rest.kegg.jp/list/pathway/hsa. For each pathway, drug and gene sets are available. For example, the sets of drugs and genes categorized in the ErbB signaling pathway, whose pathway ID is hsa04012, are available at https://rest.kegg.jp/get/hsa04012.
Code availability
Acknowledgements
This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3, JST SPRING JPMJSP2110.
Author contributions statement
H.Y., R.H., K.A., Ma.O. and H.S. conceived the experiments, H.Y., R.H., K.A., and K.M. conducted the experiments, H.Y., R.H., K.A., K.M., S.S., Ma.O., and H.S. analysed the results, R.H. and Mo.O. surveyed existing research. All authors reviewed the manuscript.
Supplementary Information
1 Details of analogy tasks
1.1 Analogy tasks for drug-gene pairs.
1.1.1 Pathway-wise setting
In Eq. (14), the estimator $\hat{\mathbf{v}}_{p}$ for the relation vector $\mathbf{v}_{p}$ is calculated from the embeddings of the drug-gene pairs $(d,g)\in\mathcal{R}_{p}$, while, similar to Eq. (5), the estimator for $\mathbf{v}_{p}$ can also be calculated from the embeddings of $d\in\mathcal{D}_{p}$ and $g\in\mathcal{G}_{p}$. The following equation gives a naive estimator for $\mathbf{v}_{p}$:
$\hat{\mathbf{v}}_{p,\text{naive}}:=\mathrm{E}_{\mathcal{G}_{p}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{D}_{p}}\{\mathbf{u}_{d}\},$ (S1)
$\mathrm{E}_{\mathcal{D}_{p}}\{\mathbf{u}_{d}\}=\frac{1}{|\mathcal{D}_{p}|}\sum_{d\in\mathcal{D}_{p}}\mathbf{u}_{d},\quad\mathrm{E}_{\mathcal{G}_{p}}\{\mathbf{u}_{g}\}=\frac{1}{|\mathcal{G}_{p}|}\sum_{g\in\mathcal{G}_{p}}\mathbf{u}_{g}.$ (S2)
In Eq. (S2), $\mathrm{E}_{\mathcal{D}_{p}}\{\cdot\}$ and $\mathrm{E}_{\mathcal{G}_{p}}\{\cdot\}$ are the means over the set of all drugs $\mathcal{D}$ and the set of all genes $\mathcal{G}$, respectively. Equation (S1) represents the vector difference between the mean vectors of $\mathcal{D}_{p}$ and $\mathcal{G}_{p}$. However, similar to $\hat{\mathbf{v}}_{\text{naive}}$ in Eq. (5), the definition of $\hat{\mathbf{v}}_{p,\text{naive}}$ in Eq. (S1) includes the embeddings of unrelated genes and drugs. Therefore, in Supplementary Information 2.1.1, we compare $\hat{\mathbf{v}}$ and $\hat{\mathbf{v}}_{\text{naive}}$, as well as $\hat{\mathbf{v}}_{p}$ and $\hat{\mathbf{v}}_{p,\text{naive}}$.
Note the following in setting P2:
  • Since $D_{p}\subset\mathcal{D}_{p}$ and $D_{p}=\pi_{\mathcal{D}_{p}}(\mathcal{R}_{p})=\pi_{\mathcal{D}}(\mathcal{R}_{p})\subset\pi_{\mathcal{D}}(\mathcal{R})=D$, it follows that $D_{p}\subset\mathcal{D}_{p}\cap D$. For $d\in\mathcal{D}_{p}\cap D$, there may be $g\in[d]\setminus[d]_{p}$. Therefore, in Eq. (13), we consider $d\in\mathcal{D}_{p}\cap D$ and $g\in[d]$, instead of $(d,g)\in\mathcal{R}_{p}$.
Fig. S1 shows a specific example to illustrate the differences between settings P1 and P2. For drug-gene pairs with drug-gene relations, setting P1 focuses on drugs and genes categorized in the same pathway, while setting P2 considers those genes as well as genes not categorized in the pathway.
1.2 Analogy tasks for drug-gene pairs by year
1.2.1 Global setting by year
Similar to $\hat{\mathbf{v}}$ in Eq. (7), to measure the performance of the estimator $\hat{\mathbf{v}}^{y}$ in Eq. (21), we prepare the evaluation of the analogy tasks. For $I\in\{(-\infty,\infty),L_{y},U_{y}\}$, using the projection operations in Eq. (10), we define $D^{y\mid I}:=\pi_{\mathcal{D}^{y}}(\mathcal{R}^{y\mid I})\subset\mathcal{D}^{y}$ and $G^{y\mid I}:=\pi_{\mathcal{G}^{y}}(\mathcal{R}^{y\mid I})\subset\mathcal{G}^{y}$. Also we define $D^{y}:=D^{y\mid(-\infty,\infty)}$ and $G^{y}:=G^{y\mid(-\infty,\infty)}$.
Similar to Eq. (11), for $I\in\{(-\infty,\infty),L_{y},U_{y}\}$, we define $[d]^{y\mid I}\subset\mathcal{G}^{y}$ as the set of genes that have drug-gene relations with a drug $d\in D^{y\mid I}$, and $[g]^{y\mid I}\subset\mathcal{D}^{y}$ as the set of drugs that have drug-gene relations with a gene $g\in G^{y\mid I}$. These are formally defined as follows:
$[d]^{y\mid I}:=\{g\mid(d,g)\in\mathcal{R}^{y|I}\}\subset\mathcal{G}^{y},\quad[g]^{y\mid I}:=\{d\mid(d,g)\in\mathcal{R}^{y|I}\}\subset\mathcal{D}^{y}.$ (S3)
In particular, we define $[d]^{y}:=[d]^{y\mid(-\infty,\infty)}$ and $[g]^{y}:=[g]^{y\mid(-\infty,\infty)}$. Given the above, we perform the analogy tasks in the following setting.
Setting Y1.
For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct. In other words, for a query drug $d\in D^{y\mid U_{y}}$, the set of answer genes is $[d]^{y\mid U_{y}}$. The search space is not the set of all genes $\mathcal{G}^{y}$, but the gene set $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct.
Setting Y2.
For the target genes that have drug-gene relations with a drug $d$, these genes are considered correct regardless of whether their relations appeared up to or after year $y$. In other words, for a query drug $d\in D^{y}$, the set of answer genes is $[d]^{y}$, where $[d]^{y}=[d]^{y\mid(-\infty,\infty)}$. The search space is the set of all genes $\mathcal{G}^{y}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y}$, then the prediction is considered correct.
Note the following in settings Y1 and Y2:
  • In setting Y1, the search space is $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. This ensures that predicted target genes have the relations that appeared only after year $y$.
  • In setting Y2, to evaluate the ability to predict target genes regardless of whether their relations appeared up to or after year $y$, we consider $(d,g)\in\mathcal{R}^{y}$ instead of $(d,g)\in\mathcal{R}^{y\mid U_{y}}$ in Eq. (20).
BCV Ours
$|\mathcal{P}|$ 136 129
$\sum_{p\in\mathcal{P}}|D_{p}|$ 4087 3612
$\sum_{p\in\mathcal{P}}|\mathcal{D}_{p}\cap D|$ 4091 3614
$\sum_{p\in\mathcal{P}}|G_{p}|$ 1251 1178
$\sum_{p\in\mathcal{P}}|\mathcal{G}_{p}\cap G|$ 3463 3186
$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in D_{p}}\{|[d]_{p}|\}\}$ 1.965 1.990
$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in G_{p}}\{|[g]_{p}|\}\}$ 5.050 4.847
Table S1: Statistics for settings P1 and P2.
(a) P1
(b) P2
Figure S1: Differences between settings (a) P1 and (b) P2 using a specific example. We set $\mathcal{D}=\{d_{1},d_{2},d_{3},d_{4},d_{5},d_{6}\}$, $\mathcal{G}=\{g_{1},g_{2},g_{3},g_{4},g_{5},g_{6},g_{7},g_{8}\}$, and =$\mathcal{R}=\{(d_{1},g_{1}),(d_{1},g_{2}),(d_{1},g_{3}),(d_{2},g_{4}),(d_{3},g_{5}),(d_{4},g_{6})\}$. According to the definitions of $D$ and $G$, $D=\{d_{1},d_{2},d_{3},d_{4}\}$ and $G=\{g_{1},g_{2},g_{3},g_{4},g_{5},g_{6}\}$. For the pathway $p$, we set $\mathcal{D}_{p}=\{d_{1},d_{2},d_{3},d_{5}\}$ and $\mathcal{G}_{p}=\{g_{1},g_{2},g_{4},g_{7}\}$. According to the definition of $\mathcal{R}_{p}$, $\mathcal{R}_{p}=\{(d_{1},g_{1}),(d_{1},g_{2}),(d_{2},g_{4})\}$. The definitions of $D_{p}$ and $G_{p}$ then give $D_{p}=\{d_{1},d_{2}\}$ and $G_{p}=\{g_{1},g_{2},g_{4}\}$, based on $\mathcal{R}_{p}$. In setting P1, the query set is $D_{p}=\{d_{1},d_{2}\}$, where the set of answer genes for $d_{1}$ is $[d_{1}]_{p}=\{g_{1},g_{2}\}$ and for $d_{2}$ is $[d_{2}]_{p}=\{g_{4}\}$. In setting P2, the query set is $\mathcal{D}_{p}\cap D=\{d_{1},d_{2},d_{3}\}$, where the set of answer genes for $d_{1}$ is $[d_{1}]=\{g_{1},g_{2},g_{3}\}$, for $d_{2}$ is $[d_{2}]=\{g_{4}\}$, and for $d_{3}$ is $[d_{3}]=\{g_{5}\}$.
1.2.2 Pathway-wise setting by year
Based on the analogy tasks in settings P1, P2, Y1, and Y2, we consider the analogy tasks in the pathway-wise setting by year, where drugs and genes are categorized based on pathways in datasets divided by year.
Consider a fixed year $y$. When learning embeddings using PubMed abstracts up to year $y$ as training data, we define $\mathcal{D}_{p}^{y}\subset\mathcal{D}^{y}$ and $\mathcal{G}_{p}^{y}\subset\mathcal{G}^{y}$ as the sets of drugs and genes that are categorized in each pathway $p\in\mathcal{P}$ and appeared up to year $y$, respectively. We then restrict the set $\mathcal{R}^{y}$ in Eq. (18) to each pathway $p$ and define the subset of the set $\mathcal{R}_{p}^{y}$ as follows:
$\mathcal{R}_{p}^{y}:=\{(d,g)\in\mathcal{R}^{y}\mid d\in\mathcal{D}_{p}^{y},\,g\in\mathcal{G}_{p}^{y}\}\subset\mathcal{D}_{p}^{y}\times\mathcal{G}_{p}^{y}.$ (S4)
For $(d,g)\in\mathcal{R}_{p}^{y}$, we divide $\mathcal{R}_{p}^{y}$ into two subsets based on whether $y_{(d,g)}\leq y$ or $y<y_{(d,g)}$. Similar to $\mathcal{R}^{y\mid I}$ in Eq. (19), for an interval $I$, we define the set $\mathcal{R}_{p}^{y\mid I}$, which consists of all $(d,g)\in\mathcal{R}_{p}^{y}$ such that $y_{(d,g)}\in I$, as follows:
$\mathcal{R}^{y|I}_{p}:=\{(d,g)\in\mathcal{R}_{p}^{y}\mid y_{(d,g)}\in I\}.$ (S5)
Using $L_{y}=(-\infty,y]$ and $U_{y}=(y,\infty)$, the set of drug-gene pairs that satisfies $y_{(d,g)}\leq y$ is $\mathcal{R}_{p}^{y\mid L_{y}}$, and the set of drug-gene pairs that satisfies $y<y_{(d,g)}$ is $\mathcal{R}_{p}^{y\mid U_{y}}$. By definition, $\mathcal{R}_{p}^{y\mid L_{y}}\cap\mathcal{R}_{p}^{y\mid U_{y}}=\emptyset$ and $\mathcal{R}_{p}^{y\mid L_{y}}\cup\mathcal{R}_{p}^{y\mid U_{y}}=\mathcal{R}_{p}^{y\mid(-\infty,\infty)}=\mathcal{R}_{p}^{y}$.
In analogy tasks, we use “known” $\mathcal{R}_{p}^{y\mid L_{y}}$ and then predict the target genes $g$ from a drug $d$ for $(d,g)$ in “unknown” $\mathcal{R}_{p}^{y\mid U_{y}}$. Using the vector $\mathbf{v}_{p}^{y}$, which represents the drug-gene relations and is derived from $\mathcal{R}_{p}^{y\mid L_{y}}$, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}_{p}^{y}$ to $\mathbf{u}_{d}$:
$\mathbf{u}_{d}+\mathbf{v}_{p}^{y}\approx\mathbf{u}_{g}.$ (S6)
Similar to equations (13) and (20), Eq. (S6) corresponds to Eq. (4). Therefore, an estimator $\hat{\mathbf{v}}_{p}^{y}$ for the relation vector $\mathbf{v}_{p}^{y}$ in Eq. (S6) is defined as follows:
$\hat{\mathbf{v}}_{p}^{y}:=\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}_{p}^{y\mid L_{y}}\right|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}(\mathbf{u}_{g}-\mathbf{u}_{d}).$ (S7)
where $\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}_{p}^{y\mid L_{y}}$. Similar to Eq. (7), Eq. (S7) defines the estimator $\hat{\mathbf{v}}_{p}^{y}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}$. For easy comparison with Eq. (5), Eq. (S7) is rewritten as the difference of mean vectors:
$\hat{\mathbf{v}}_{p}^{y}=\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{d}\},$ (S8)
$\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}\mathbf{u}_{d}.$ (S9)
Similar to $\hat{\mathbf{v}}$ in Eq. (7), to measure the performance of the estimator $\hat{\mathbf{v}}_{p}^{y}$ in Eq. (S7), we prepare the evaluation of the analogy tasks. For $I\in\{(-\infty,\infty),L_{y},U_{y}\}$, using the projection operations in Eq. (10), we define $D_{p}^{y\mid I}:=\pi_{\mathcal{D}_{p}^{y}}(\mathcal{R}_{p}^{y\mid I})$ and $G_{p}^{y\mid I}:=\pi_{\mathcal{G}_{p}^{y}}(\mathcal{R}_{p}^{y\mid I})$. Also we define $D_{p}^{y}:=D_{p}^{y\mid(-\infty,\infty)}$ and $G_{p}^{y}:=G_{p}^{y\mid(-\infty,\infty)}$.
Similar to Eq. (11), for $I\in\{(-\infty,\infty),L_{y},U_{y}\}$, we define $[d]_{p}^{y\mid I}\subset\mathcal{G}_{p}^{y}$ as the set of genes that have drug-gene relations with a drug $d\in D_{p}^{y\mid I}$, and $[g]_{p}^{y\mid I}\subset\mathcal{D}_{p}^{y}$ as the set of drugs that have drug-gene relations with a gene $g\in G_{p}^{y\mid I}$. These are formally defined as follows:
$[d]_{p}^{y\mid I}:=\{g\mid(d,g)\in\mathcal{R}_{p}^{y\mid I}\}\subset\mathcal{G}_{p}^{y},\quad[g]_{p}^{y\mid I}:=\{d\mid(d,g)\in\mathcal{R}_{p}^{y\mid I}\}\subset\mathcal{D}_{p}^{y}.$ (S10)
In particular, we define $[d]_{p}^{y}:=[d]_{p}^{y\mid(-\infty,\infty)}$ and $[g]_{p}^{y}:=[g]_{p}^{y\mid(-\infty,\infty)}$. Similar to settings P1, P2, Y1, and Y2, we perform the analogy tasks in the following four settings.
Setting P1Y1.
For the target genes that have drug-gene relations with a drug $d$, only genes that are categorized in the same pathway $p$ as $d$ and whose relations appeared after year $y$ are considered correct. In other words, for a query drug $d\in D_{p}^{y\mid U_{y}}$, the set of answer genes is $[d]_{p}^{y\mid U_{y}}$. The search space is not the set of all genes $\mathcal{G}^{y}$, but the gene set $\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]_{p}^{y\mid U_{y}}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]_{p}^{y\mid U_{y}}$, then the prediction is considered correct.
Setting P2Y1.
For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct, regardless of whether they are categorized in the same pathway $p$ as the drug $d$ or not. In other words, for a query drug $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$, the set of answer genes is $[d]^{y\mid U_{y}}$. The search space is not the set of all genes $\mathcal{G}^{y}$, but the gene set $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct.
Setting P1Y2.
For the target genes that have drug-gene relations with a drug $d$, only genes categorized in the same pathway as the drug $d$ are considered correct, regardless of whether their relations appeared up to or after year $y$. In other words, for a query drug $d\in D_{p}^{y}$, the set of answer genes is $[d]_{p}^{y}$. The search space is the set of all genes $\mathcal{G}^{y}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]_{p}^{y}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]_{p}^{y}$, then the prediction is considered correct.
Setting P2Y2.
For the target genes that have drug-gene relations with a drug $d$, these genes are considered correct regardless of whether they are categorized in the same pathway or not and whether their relations appeared up to or after year $y$. In other words, for a query drug $d\in\mathcal{D}_{p}^{y}\cap D^{y}$, the set of answer genes is $[d]^{y}$. The search space is the set of all genes $\mathcal{G}^{y}$. The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y}$, then the prediction is considered correct. We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}$. For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y}$, then the prediction is considered correct.
Note the following in settings P1Y1, P1Y2, P2Y1 and P2Y2:
  • In setting P1Y1, the search space is $\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$. This ensures that predicted target genes, which are categorized in the same pathway $p$ as the drug $d$, have the relations that appeared only after year $y$.
  • In setting P2Y1, since $D_{p}^{y\mid U_{y}}\subset\mathcal{D}_{p}^{y}$ and $D_{p}^{y\mid U_{y}}=\pi_{\mathcal{D}_{p}^{y}}\left(\mathcal{R}_{p}^{y\mid U_{y}}\right)=\pi_{\mathcal{D}^{y}}\left(\mathcal{R}_{p}^{y\mid U_{y}}\right)\subset\pi_{\mathcal{D}^{y}}\left(\mathcal{R}^{y\mid U_{y}}\right)=D^{y\mid U_{y}}$, it follows that $D_{p}^{y\mid U_{y}}\subset\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$. For $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$, there may be $g\in[d]^{y\mid U_{y}}\setminus[d]_{p}^{y\mid U_{y}}$. Therefore, in Eq. (S6), we consider $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ and $g\in[d]^{y\mid U_{y}}$, instead of $(d,g)\in\mathcal{R}_{p}^{y\mid U_{y}}$. The search space is $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. This ensure that predicted target genes have the relations that appeared only after year $y$.
  • In setting P1Y2, for a query drug $d\in D_{p}^{y}$, to evaluate the ability to predict target genes categorized in the same pathway $p$ as the drug $d$, regardless of whether their relations appeared up to or after year $y$, we consider $(d,g)\in\mathcal{R}_{p}^{y}$ instead of $(d,g)\in\mathcal{R}_{p}^{y\mid U_{y}}$ in Eq. (20).
  • In setting P2Y2, since $D_{p}^{y}\subset\mathcal{D}_{p}^{y}$ and $D_{p}^{y}=\pi_{\mathcal{D}_{p}^{y}}\left(\mathcal{R}_{p}^{y}\right)=\pi_{\mathcal{D}^{y}}\left(\mathcal{R}_{p}^{y}\right)\subset\pi_{\mathcal{D}^{y}}\left(\mathcal{R}^{y}\right)=D^{y}$, it follows that $D_{p}^{y}\subset\mathcal{D}_{p}^{y}\cap D^{y}$. For $d\in\mathcal{D}_{p}^{y}\cap D^{y}$, there may be $g\in[d]^{y}\setminus[d]_{p}^{y}$. Therefore, in Eq. (S6), we consider $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ and $g\in[d]^{y}$, instead of $(d,g)\in\mathcal{R}_{p}^{y\mid U_{y}}$.
1.3 Comparison of Experimental Settings
Setting Query Answer Set Search Space
G $d\in D$ $[d]$ $\mathcal{G}$
P1 $d\in D_{p}$ $[d]_{p}$ $\mathcal{G}$
P2 $d\in\mathcal{D}_{p}\cap D$ $[d]$ $\mathcal{G}$
Y1 $d\in D^{y\mid U_{y}}$ $[d]^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$
Y2 $d\in D^{y}$ $[d]^{y}$ $\mathcal{G}^{y}$
P1Y1 $d\in D_{p}^{y\mid U_{y}}$ $[d]_{p}^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$
P2Y1 $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ $[d]^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$
P1Y2 $d\in D_{p}^{y}$ $[d]_{p}^{y}$ $\mathcal{G}^{y}$
P2Y2 $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ $[d]^{y}$ $\mathcal{G}^{y}$
Table S2: Query, answer set, and search space for each setting for predicting genes from drugs.
Setting Query Answer Set Search Space
${}^{{}^{\prime}}$ $g\in G$ $[g]$ $\mathcal{D}$
${}^{{}^{\prime}}$ $g\in G_{p}$ $[g]_{p}$ $\mathcal{D}$
${}^{{}^{\prime}}$ $g\in\mathcal{G}_{p}\cap G$ $[g]$ $\mathcal{D}$
${}^{{}^{\prime}}$ $g\in G^{y\mid U_{y}}$ $[g]^{y\mid U_{y}}$ $\mathcal{D}^{y}\setminus[g]^{y\mid L_{y}}$
${}^{{}^{\prime}}$ $g\in G^{y}$ $[g]^{y}$ $\mathcal{D}^{y}$
Table S3: Query, answer set, and search space for each setting for predicting drugs from genes.
Hyperparameter Values
Training epochs $10$
Down-sampling threshold $10^{-5}$
Learning rate $0.025$
Window size $5$
Negative samples $5$
Minimal word occurrence $30$
Dimension $300$
Table S4: Hyperparameter for our skip-gram.
Table S2 shows the settings for predicting genes from drugs, as used in the main text. Similarly, Table S3 shows the settings for predicting drugs from genes. In Table S3, we adapted the notations used for predicting genes from drugs in Table S2 to those used for predicting drugs from genes, denoting them as ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, and ${}^{{}^{\prime}}$.
1.4 Embeddings
The hyperparameters used to train our skip-gram are shown in Table S4.
1.5 Datasets
In the BioConceptVec vocabulary, genes are represented by gene IDs[54] and drugs are represented by MeSH (Medical Subject Headings, https://www.nlm.nih.gov/mesh/meshhome.html) IDs. Therefore, a conversion from IDs to names is necessary when performing experiments. In addition, in the data obtained from AsuratDB and the KEGG API, genes are represented by gene IDs and drugs are represented by KEGG IDs. Therefore, a conversion from KEGG IDs to MeSH IDs is also required for use with BioConceptVec. The procedures for converting these IDs are described in the following sections.
1.5.1 Conversion from MeSH ID to drug name
The BioConceptVec vocabulary registers drugs using MeSH (Medical Subject Headings, https://www.nlm.nih.gov/mesh/meshhome.html) IDs. This is due to the normalization of drug names using MeSH IDs in PubTator. Therefore, we explain the procedure for converting MeSH IDs to drug names. We used MeSH SPARQL (https://hhs.github.io/meshrdf/sparql-and-uri-requests) to obtain the MeSH headings corresponding to the MeSH IDs in the BioConceptVec vocabulary. We used these headings to convert MeSH IDs to drug names. We also applied this conversion to drugs in the vocabulary of our trained skip-gram model.
Embeddings Setting Method Centering Top1 Top10 MRR
Ours G Random 0.000 0.000 0.001
$\hat{\mathbf{v}}_{\text{naive}}$ 0.067 0.197 0.111
$\checkmark$ 0.125 0.381 0.209
$\hat{\mathbf{v}}$ 0.206 0.455 0.289
$\checkmark$ 0.300 0.686 0.426
P1 Random 0.000 0.001 0.001
$\hat{\mathbf{v}}_{p,\text{naive}}$ 0.364 0.661 0.470
$\checkmark$ 0.455 0.810 0.581
$\hat{\mathbf{v}}_{p}$ 0.521 0.788 0.615
$\checkmark$ 0.589 0.862 0.685
P2 Random 0.000 0.001 0.001
$\hat{\mathbf{v}}_{p,\text{naive}}$ 0.376 0.682 0.484
$\checkmark$ 0.473 0.837 0.601
$\hat{\mathbf{v}}_{p}$ 0.530 0.803 0.626
$\checkmark$ 0.600 0.880 0.700
Table S5: Results of the naive estimators and the centering ablation study for settings G, P1, and P2.
Metric
Embed. Setting Method Top1 Top10 MRR
BCV ${}^{{}^{\prime}}$ Random 0.000 0.001 0.001
$\hat{\mathbf{v}}^{\prime}$ 0.233 0.560 0.345
${}^{{}^{\prime}}$ Random 0.000 0.001 0.001
$\hat{\mathbf{v}}_{p}^{\prime}$ 0.426 0.679 0.515
${}^{{}^{\prime}}$ Random 0.000 0.001 0.001
$\hat{\mathbf{v}}_{p}^{\prime}$ 0.248 0.507 0.337
Ours ${}^{{}^{\prime}}$ Random 0.000 0.003 0.002
$\hat{\mathbf{v}}^{\prime}$ 0.240 0.577 0.351
${}^{{}^{\prime}}$ Random 0.000 0.003 0.002
$\hat{\mathbf{v}}_{p}^{\prime}$ 0.478 0.751 0.571
${}^{{}^{\prime}}$ Random 0.000 0.003 0.002
$\hat{\mathbf{v}}_{p}^{\prime}$ 0.279 0.516 0.359
Table S6: Gene prediction performance in settings G, P1, and P2.
Year
1970 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
$|\mathcal{D}^{y}|$ 2268 3275 5059 7125 9387 11748 14044 16827 19758 22717 25431
$|\mathcal{G}^{y}|$ 353 725 1782 3413 5976 10533 16949 24784 32264 39961 47509
$\left|\mathcal{R}^{y\mid L_{y}}\right|$ 12 37 117 256 499 859 1269 1637 1978 2354 2603
$\left|\mathcal{R}^{y\mid U_{y}}\right|$ 13 53 183 296 344 515 473 379 281 189 61
$\left|D^{y\mid U_{y}}\right|$ 12 48 154 272 288 369 325 256 189 131 48
$|D^{y}|$ 22 79 246 475 663 893 1047 1151 1269 1423 1495
$\left|G^{y\mid U_{y}}\right|$ 6 14 34 55 89 137 173 153 140 115 44
$|G^{y}|$ 8 20 46 74 122 215 325 398 477 540 570
$\mathrm{E}_{d\in D^{y\mid L_{y}}}\left\{\left|[d]^{y\mid L_{y}}\right|\right\}$ 1.091 1.028 1.104 1.133 1.188 1.372 1.500 1.613 1.665 1.722 1.760
$\mathrm{E}_{d\in D^{y\mid U_{y}}}\left\{\left|[d]^{y\mid U_{y}}\right|\right\}$ 1.083 1.104 1.188 1.088 1.194 1.396 1.455 1.480 1.487 1.443 1.271
$\mathrm{E}_{g\in G^{y\mid L_{y}}}\left\{\left|[g]^{y\mid L_{y}}\right|\right\}$ 3.000 2.467 4.179 5.020 5.940 5.335 4.789 4.547 4.425 4.467 4.591
$\mathrm{E}_{g\in G^{y\mid U_{y}}}\left\{\left|[g]^{y\mid U_{y}}\right|\right\}$ 2.167 3.786 5.382 5.382 3.865 3.759 2.734 2.477 2.007 1.643 1.386
Table S7: Statistics for settings Y1 and Y2.
1.5.2 Conversion from gene ID to gene name
The BioConceptVec vocabulary registers genes using gene IDs. Therefore, we explain the procedure for converting gene IDs to gene names. We used the KEGG API to obtain the names corresponding to the gene IDs in the BioConceptVec vocabulary. This API allows for batch requests; for example, to process gene IDs 20, 21, 22, 23, 24, the data is available at https://rest.kegg.jp/list/hsa:20+hsa:21+hsa:22+hsa:23+hsa:24. If multiple names corresponded to a single gene ID, we chose the first name. We also applied this conversion to genes in the vocabulary of our trained skip-gram model.
1.5.3 Conversion from KEGG ID to MeSH ID
In AsuratDB and the KEGG API, which is used to obtain drug-gene relations, genes are represented by gene IDs, while drugs are represented by KEGG IDs. Therefore, to use the data in BioConceptVec, we need to convert the KEGG IDs to MeSH IDs. We followed a four-step procedure (I), (II), (III), and (IV) to convert KEGG IDs to MeSH IDs.
(I) KEGG ID to PubChem SID or ChEBI ID
Some drugs in KEGG are manually linked to external databases such as PubChem (https://pubchem.ncbi.nlm.nih.gov/) and ChEBI (https://www.ebi.ac.uk/chebi/), which are larger databases than KEGG. Therefore, we used these databases to link their IDs to MeSH IDs.
Since PubChem is maintained by the NCBI (National Center for Biotechnology Information, https://www.ncbi.nlm.nih.gov/) that also maintains the MeSH database, we prioritized the conversion of KEGG IDs to PubChem Substance IDs (SIDs). If we could not convert KEGG IDs directly to PubChem SIDs, we converted them to ChEBI IDs. The conversion from KEGG IDs to PubChem SIDs is available at https://rest.kegg.jp/conv/pubchem/drug, and the conversion to ChEBI IDs is available at https://rest.kegg.jp/conv/chebi/drug.
(II) PubChem SID or ChEBI ID to PubChem CID
Next, we converted PubChem SIDs and ChEBI IDs to PubChem Compound IDs (CIDs), which are associated with MeSH IDs. The conversion from PubChem SIDs to PubChem CIDs is available at https://pubchem.ncbi.nlm.nih.gov/rest/pug/substance/sourceall/KEGG/cids/json. To convert ChEBI IDs to PubChem CIDs, we used the PubChem API. For example, for ChEBI ID 39112, the corresponding data can be obtained at https://pubchem.ncbi.nlm.nih.gov/rest/pug//compound/xref/RegistryID/chebi:39112/cids/json.
(III) Convert CIDs to MeSH Headings
We used the Entrez (https://www.ncbi.nlm.nih.gov/Web/Search/entrezfs.html) System API provided by NCBI, and obtained the MeSH headings associated with each CID. This API allows for batch requests. For example, for CIDs 5328940 and 156413, the corresponding data can be obtained at https://eutils.ncbi.nlm.nih.gov/entrez/eutils/esummary.fcgi?db=pccompound&id=5328940,156413&retmode=json.
(IV) MeSH Headings to MeSH ID
Since we had already obtained the correspondence between MeSH IDs and MeSH headings in Supplementary Information 1.5.1, we could match MeSH IDs with the MeSH headings obtained through the Entrez system API. In this way, we converted the original KEGG IDs to MeSH IDs. Note that multiple KEGG IDs may correspond to the same MeSH ID. In such cases, we chose the smallest KEGG ID to associate with the MeSH ID.
Year
Metric Setting Method 1970 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
Top1 Y1 Random 0.000 0.000 0.003 0.000 0.000 0.001 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}^{y}$ 0.250 0.333 0.208 0.165 0.132 0.076 0.092 0.062 0.090 0.107 0.021
Y2 Random 0.005 0.005 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}^{y}$ 0.500 0.481 0.350 0.318 0.305 0.299 0.328 0.326 0.318 0.376 0.318
Top10 Y1 Random 0.008 0.010 0.008 0.002 0.003 0.001 0.000 0.001 0.001 0.000 0.000
$\hat{\mathbf{v}}^{y}$ 0.500 0.625 0.487 0.438 0.410 0.347 0.397 0.352 0.344 0.328 0.271
Y2 Random 0.041 0.022 0.007 0.003 0.002 0.001 0.001 0.001 0.000 0.000 0.000
$\hat{\mathbf{v}}^{y}$ 0.727 0.785 0.699 0.657 0.665 0.679 0.735 0.744 0.702 0.773 0.764
MRR Y1 Random 0.012 0.007 0.005 0.003 0.002 0.002 0.001 0.001 0.000 0.000 0.000
$\hat{\mathbf{v}}^{y}$ 0.331 0.438 0.304 0.253 0.214 0.163 0.182 0.147 0.173 0.188 0.103
Y2 Random 0.017 0.010 0.006 0.003 0.002 0.002 0.001 0.001 0.001 0.000 0.000
$\hat{\mathbf{v}}^{y}$ 0.590 0.592 0.476 0.434 0.425 0.429 0.468 0.463 0.451 0.513 0.469
Table S8: Gene prediction performance in settings Y1 and Y2.
Year
1970 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
$\sum_{p\in\mathcal{P}}\left|D_{p}^{y\mid U_{y}}\right|$ 14 104 314 507 521 601 526 439 312 201 67
$\sum_{p\in\mathcal{P}}|\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}|$ 16 115 363 572 573 732 644 529 401 264 98
$\sum_{p\in\mathcal{P}}|D_{p}^{y}|$ 26 156 513 959 1363 1728 1999 2205 2419 2711 2821
$\sum_{p\in\mathcal{P}}|\mathcal{D}_{p}^{y}\cap D^{y}|$ 30 174 566 1028 1419 1794 2068 2263 2480 2773 2886
$\sum_{p\in\mathcal{P}}\left|G_{p}^{y\mid U_{y}}\right|$ 6 29 88 129 195 305 326 269 228 168 62
$\sum_{p\in\mathcal{P}}|\mathcal{G}_{p}^{y}\cap G^{y\mid U_{y}}|$ 15 50 177 281 520 935 1087 859 774 682 275
$\sum_{p\in\mathcal{P}}|G_{p}^{y}|$ 6 38 110 170 281 484 658 772 925 1066 1124
$\sum_{p\in\mathcal{P}}|\mathcal{G}_{p}^{y}\cap G^{y}|$ 18 68 207 367 652 1356 2015 2359 3087 3518 3627
$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{d\in D_{p}}\left\{\left|[d]_{p}^{y\mid L_{y}}\right|\right\}\right\}$ 4.000 2.500 5.921 8.148 11.407 12.667 14.319 14.649 14.631 15.465 15.977
$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{d\in D_{p}}\left\{|[d]_{p}^{y\mid U_{y}}|\right\}\right\}$ 2.333 4.160 6.038 7.243 6.432 7.805 6.188 5.776 5.032 4.467 3.190
$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{g\in G_{p}}\left\{\left|[g]_{p}^{y\mid L_{y}}\right|\right\}\right\}$ 1.000 1.136 1.789 1.918 2.444 3.441 4.483 5.216 5.529 6.171 6.400
$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{g\in G_{p}}\left\{\left|[g]_{p}^{y\mid U_{y}}\right|\right\}\right\}$ 1.000 1.160 1.692 1.843 2.407 3.961 3.835 3.539 3.677 3.733 2.952
Table S9: Statistics for settings P1Y1, P1Y2, P2Y1, and P2Y2.
2 Details of experimental results
2.1 Analogy tasks for drug-gene pairs
2.1.1 Comparison of estimators and centering ablation study
For settings G, P1, and P2, Table S5 shows the results of the naive estimators defined in equations (5) and (S1) and the centering ablation study, using our skip-gram. In all settings, the results of the estimators $\hat{\mathbf{v}}$ and $\hat{\mathbf{v}}_{p}$ calculated from drug-gene relations outperformed those of the naive estimators $\hat{\mathbf{v}}_{\text{naive}}$ and $\hat{\mathbf{v}}_{p,\text{naive}}$. We also observed performance improvements due to centering. Based on this, we showed the results using the estimators calculated from the drug-gene relations with centering applied in Table 3.
2.1.2 Prediction of drugs from genes
As seen in Supplementary Information 1.3, Table S3 defines settings ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, and ${}^{{}^{\prime}}$ for the analogy tasks for predicting drugs from genes. We then define estimators to perform the analogy tasks in these settings. Similar to the estimator $\hat{\mathbf{v}}$ in Eq. (7) for setting G, we define the estimator $\hat{\mathbf{v}}^{\prime}$ for setting ${}^{{}^{\prime}}$:
$\hat{\mathbf{v}}^{\prime}:=-\hat{\mathbf{v}}=\mathrm{E}_{\mathcal{\mathcal{R}}}\{\mathbf{u}_{d}-\mathbf{u}_{g}\}=\frac{1}{|\mathcal{R}|}\sum_{(d,g)\in\mathcal{R}}(\mathbf{u}_{d}-\mathbf{u}_{g}).$ (S11)
Similar to the estimator $\hat{\mathbf{v}}_{p}$ in Eq. (14) for settings P1 and P2, we also define the estimator $\hat{\mathbf{v}}_{p}^{\prime}$ for settings ${}^{{}^{\prime}}$ and ${}^{{}^{\prime}}$:
$\hat{\mathbf{v}}_{p}^{\prime}:=-\hat{\mathbf{v}}_{p}=\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{d}-\mathbf{u}_{g}\}=\frac{1}{|\mathcal{R}_{p}|}\sum_{(d,g)\in\mathcal{R}_{p}}(\mathbf{u}_{d}-\mathbf{u}_{g}).$ (S12)
For settings ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, and ${}^{{}^{\prime}}$, we performed the analogy tasks using $\hat{\mathbf{v}}^{\prime}$ and $\hat{\mathbf{v}}_{p}^{\prime}$, and Table S6 shows the results. Similar to Table 3, the estimators calculated from the drug-gene relations outperformed the random baseline.
While Table 3 shows higher scores for setting P2 compared to setting P1, Table S6 shows higher scores for setting P1 compared to setting P2. To explain these results, we focus on the sizes of the query and answer sets, using actual values from BioConceptVec. As shown in Table S1, the sizes of the query sets for settings P1 and P2 are $\sum_{p\in\mathcal{P}}|D_{p}|=4087$ and $\sum_{p\in\mathcal{P}}|\mathcal{D}_{p}\cap D|=4091$, respectively. The ratio is $\sum_{p\in\mathcal{P}}|\mathcal{D}_{p}\cap D|/\sum_{p\in\mathcal{P}}|D_{p}|\approx 1.001$. On the other hand, for settings ${}^{{}^{\prime}}$ and ${}^{{}^{\prime}}$, the sizes of the query sets are $\sum_{p\in\mathcal{P}}|G_{p}|=1251$ and $\sum_{p\in\mathcal{P}}|\mathcal{G}_{p}\cap G|=3463$, respectively, with a ratio of $\sum_{p\in\mathcal{P}}|\mathcal{G}_{p}\cap G|/\sum_{p\in\mathcal{P}}|G_{p}|\approx 2.768$. For settings P1, P2, P1, and P2, the answer sets defined in Tables S2 and S3 are $[d]_{p}$, $[d]$, and $[g]_{p}$, $[g]$, respectively. By definition, the sizes of the answer sets satisfy $|[d]_{p}|\leq|[d]|$ and $|[g]_{p}|\leq|[g]|$. In fact, the expected values of them are $\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in D_{p}}\{|[d]_{p}|\}\}=1.965$, $\mathrm{E}_{d\in D}\{|[d]|\}=2.938$, $\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in G_{p}}\{|[g]_{p}|\}\}=5.050$, and $\mathrm{E}_{g\in G}\{|[g]|\}=10.008$, respectively. In settings P1 and P2, the sizes of the query sets are nearly identical, but the larger answer sets in setting P2 probably make the task easier. On the other hand, the sizes of the query sets in setting ${}^{{}^{\prime}}$ are more than twice as large as those in ${}^{{}^{\prime}}$, which likely contributes to the lower evaluation metric scores in setting ${}^{{}^{\prime}}$.
2.2 Analogy tasks for drug-gene pairs by year
(a) P2Y1
(b) P2Y2
Figure S2: Gene prediction performance and the number of queries for each year in settings P2Y1 and P2Y2.
Year
Metric Setting Method 1970 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
Top1 P1Y1 Random 0.014 0.000 0.001 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.653 0.515 0.484 0.394 0.303 0.275 0.263 0.256 0.285 0.269
P1Y2 Random 0.000 0.001 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.780 0.674 0.656 0.651 0.633 0.636 0.627 0.630 0.659 0.623
P2Y1 Random 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.625 0.467 0.437 0.358 0.250 0.226 0.221 0.209 0.236 0.184
P2Y2 Random 0.005 0.003 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.744 0.633 0.620 0.627 0.613 0.616 0.619 0.621 0.653 0.617
Top10 P1Y1 Random 0.029 0.020 0.009 0.003 0.003 0.001 0.002 0.001 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.800 0.670 0.727 0.638 0.597 0.573 0.627 0.566 0.665 0.612
P1Y2 Random 0.026 0.012 0.005 0.004 0.002 0.002 0.001 0.001 0.001 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.906 0.842 0.863 0.863 0.879 0.896 0.913 0.899 0.925 0.911
P2Y1 Random 0.000 0.013 0.006 0.002 0.002 0.001 0.001 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.762 0.627 0.674 0.590 0.533 0.502 0.567 0.513 0.563 0.480
P2Y2 Random 0.032 0.013 0.007 0.004 0.002 0.002 0.001 0.001 0.000 0.000 0.001
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.872 0.814 0.832 0.845 0.866 0.883 0.909 0.896 0.922 0.912
MRR P1Y1 Random 0.032 0.008 0.005 0.003 0.002 0.001 0.001 0.001 0.001 0.000 0.001
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.727 0.583 0.582 0.483 0.404 0.378 0.392 0.373 0.410 0.391
P1Y2 Random 0.023 0.011 0.005 0.003 0.002 0.001 0.001 0.001 0.001 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.843 0.749 0.743 0.731 0.722 0.734 0.737 0.732 0.758 0.731
P2Y1 Random 0.012 0.008 0.006 0.003 0.002 0.001 0.001 0.001 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.695 0.532 0.529 0.444 0.345 0.323 0.341 0.318 0.340 0.282
P2Y2 Random 0.019 0.010 0.006 0.003 0.002 0.001 0.001 0.001 0.001 0.001 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.808 0.711 0.708 0.709 0.704 0.716 0.729 0.725 0.752 0.727
Table S10: Gene prediction performance in settings P1Y1, P1Y2, P2Y1, and P2Y2.
2.2.1 Global setting by year
Table S7 shows the statistics for settings Y1 and Y2. As seen in Table S2, in setting Y1, the query is $d\in D^{y\mid U_{y}}$, and the search space is $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. In setting Y2, the query is $d\in D^{y}$, and the search space is $\mathcal{G}^{y}$. Table S7 shows that $D^{y\mid U_{y}}$ has increased since 1970 and started to decrease after 1995. On the other hand, $D^{y}$ and $\mathcal{G}^{y}$ show a monotonic increase over the years.
We showed the plots of the results for settings Y1 and Y2 in Fig. 3. For settings Y1 and Y2, Table S8 shows detailed results used in Fig. 3.
2.2.2 Pahtway-wise setting by year
Table S9 shows the statistics for settings P1Y1, P1Y2, P2Y1, and P2Y2. As seen in Table S2, the query is $d\in D_{p}^{y\mid U_{y}}$ in setting P1Y1, $d\in D_{p}^{y}$ in setting P1Y2, $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ in P2Y1, and $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ in setting P2Y2. Table S9 shows that $D_{p}^{y\mid U_{y}}$ and $\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ have increased since 1970 and started to decrease after 1995. On the other hand, $D_{p}^{y}$ and $\mathcal{D}_{p}^{y}\cap D^{y}$ show a monotonic increase over the years.
We showed the plots of the results for settings P1Y1 and P1Y2 in Fig. 3. Fig. S2 shows the plots of the results for settings P2Y1 and P2Y2. For settings P1Y1, P1Y2, P2Y1, and P2Y2, Table S10 shows detailed results used in Fig. 3 and Fig. S2. As shown in Table 3, settings P1 and P2 show roughly the same trends. Similarly, in Table S10 , settings P1Y1 and P2Y1, as well as P1Y2 and P2Y2, each show similar trends.

### v2

@@
 In the following sections, we perform experiments for each setting and evaluate the performance using metrics such as top-1 accuracy.
-Finally, we discuss specific prediction results and then conclude.
+Finally, we discuss the prediction results and then conclude.
 Related work
 […]
-In the biomedical field, capturing the relations between biomedical concepts such as drug-drug and protein-protein interaction is an important issue.
+In the biomedical field, capturing the relations between biomedical concepts, such as drug-drug and protein-protein interactions, is an important issue.
 Word embeddings trained on biomedical domain corpora have been used to predict drug-drug interactions for new drugs [19] and to construct networks of gene-gene interactions [20].
-Also, neural networks have been used in a number of studies to predict these relations [21, 22, 23, 24, 25].
-Technologies developed in NLP, not limited to word embedding, have been applied to a wide variety of problems in the biomedical domain: Müller et al.
+While word embeddings have shown promise in various biomedical applications, their potential for predicting drug-gene relations remains relatively unexplored.
+To the best of our knowledge, there is limited research specifically focused on utilizing analogy computation of word embeddings to predict drug-gene relations in the biological domain.
+Our study aims to address this gap by investigating the application of analogy computation of word embeddings to drug-gene relationship prediction.
+In addition, neural networks have been used in a number of studies to predict these relations [21, 22, 23, 24, 25].
+However, it is important to note that these relation extraction approaches differ from our proposed task.
+While they are useful for identifying explicitly stated relationships in text, our focus is on leveraging the latent semantic information captured by word embeddings to predict potential drug-gene relations, even when such relationships are not explicitly mentioned in the literature.
+The technologies developed in NLP, not limited to word embeddings, have been applied to a wide variety of problems in the biomedical domain, demonstrating their utility in this field.
+Müller et al.
 [26] have created an online literature search and curation system using an “ontology dictionary” obtained by text mining, and Friedman et al.
@@
 In this section, based on analogy tasks for drug-gene pairs, we explain analogy tasks in the setting where datasets are divided by year.
-In this setting, we can separate the drug-gene relations into known or unknown relations based on the chronological order they appear within the PubMed abstracts.
-By training skip-gram embeddings with this dataset, we can test if embeddings of known relations have the ability to predict unknwon relations.
-Similar to analogy tasks for drug-gene pairs, we first consider the global setting by year, where all drugs and genes are used in datasets divided by year.
-Then we consider the pathway-wise setting by year, where drugs and genes are categorized based on pathways in datasets divided by year.
+To do so, we first use embeddings trained on PubMed abstracts up to year $y$ and consider analogy tasks in a global setting such as setting G.
+Next, we separate the drug-gene relations into “known” or “unknown” based on their chronological appearance in the PubMed abstracts.
+Using the embeddings from the datasets divided by year, we test whether embeddings of “known” relations have the ability to predict “unknwon” relations.
+We then consider the pathway-wise settings by year, where drugs and genes are categorized based on pathways in the datasets divided by year.
+In these settings, we use the year-specific embeddings and consider analogy tasks in settings such as P1 and P2.
+Finally, we evaluate whether “unknown” relations can be predicted by “known” relations.
 Global setting by year
-As a first step, based on analogy tasks for drug-gene pairs, we consider analogy tasks in the global setting, where all drugs and genes are used in datasets divided by year.
-Specifically, drug-gene relations that appeared up to year $y$ are considered known, while those that appeared after year $y$ are considered unknown.
-We then train skip-gram embeddings using PubMed abstracts published before the specified year, redefine vectors representing known relations, and use these vectors to predict unknown relations.
-In preparation, we define $y_{d}$ as the year when a drug $d$ first appeared in a PubMed abstract, and $y_{g}$ as the year when a gene $g$ first appeared in a PubMed abstract.
-In addition, we define $y_{(d,g)}$ as the year when both drug $d$ and gene $g$ first appeared together in a PubMed abstract, and use $y_{(d,g)}$ as a substitute for the year when the relation $(d,g)$ was first identified.
-By definition, $\max\{y_{d},y_{g}\}\leq y_{(d,g)}$ holds.
+First, using embeddings trained on PubMed abstracts up to year $y$, we consider analogy tasks in a global setting.
+In preparation, we define $y_{d}$ as the year when a drug $d$ first appeared in a PubMed abstract and $y_{g}$ as the year when a gene $g$ first appeared in a PubMed abstract.
 Consider a fixed year $y$.
@@
 $\mathcal{R}^{y}:=\{(d,g)\in\mathcal{R}\mid d\in\mathcal{D}^{y},g\in\mathcal{G}^{y}\}\subset\mathcal{D}^{y}\times\mathcal{G}^{y}.$ (18)
-Since the vocabulary set of the trained embeddings includes both $d$ and $g$ for $(d,g)\in\mathcal{R}^{y}$, we may discuss this relation $(d,g)$.
-The relation $(d,g)$ is interpreted as either known by year $y$ if $y_{(d,g)}\leq y$ or unknown by year $y$ if $y<y_{(d,g)}$.
-We divide $\mathcal{R}^{y}$ into two subsets based on whether $y_{(d,g)}\leq y$ or $y<y_{(d,g)}$.
-Let us define two intervals, $L_{y}:=(-\infty,y]$ and $U_{y}:=(y,\infty)$.
-Considering the set for an interval $I$, consisting of all $(d,g)\in\mathcal{R}^{y}$ such that $y_{(d,g)}\in I$, i.e.,
-$\mathcal{R}^{y\mid I}:=\{(d,g)\in\mathcal{R}^{y}\mid y_{(d,g)}\in I\}\subset\mathcal{R}^{y},$ (19)
-the set of “known” relations is expressed as $\mathcal{R}^{y\mid L_{y}}$, and the set of “unknown” relations is expressed as $\mathcal{R}^{y\mid U_{y}}$.
-By definition, $\mathcal{R}^{y\mid L_{y}}\cap\mathcal{R}^{y\mid U_{y}}=\emptyset$ and $\mathcal{R}^{y\mid L_{y}}\cup\mathcal{R}^{y\mid U_{y}}=\mathcal{R}^{y\mid(-\infty,\infty)}=\mathcal{R}^{y}$.
-In analogy tasks, we use “known” $\mathcal{R}^{y\mid L_{y}}$ and then predict the target genes $g$ from a drug $d$ for $(d,g)$ in “unknown” $\mathcal{R}^{y\mid U_{y}}$.
-Using the vector $\mathbf{v}^{y}$, which represents the drug-gene relations and is derived from $\mathcal{R}^{y\mid L_{y}}$, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}^{y}$ to $\mathbf{u}_{d}$:
+Similar to the global setting, for drug-gene pairs $(d,g)\in\mathcal{R}^{y}$, we consider the analogy tasks for predicting the target genes $g$ from a drug $d$.
+To solve these analogy tasks, we use the relation vector $\mathbf{v}^{y}$.
+We predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}^{y}$ to $\mathbf{u}_{d}$:
 $\mathbf{u}_{d}+\mathbf{v}^{y}\approx\mathbf{u}_{g}.$ (19)
@@
 Therefore, similar to the estimator $\hat{\mathbf{v}}$ in Eq.
-(7), we define an estimator $\hat{\mathbf{v}}^{y}$ for the relation vector $\mathbf{v}^{y}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}^{y\mid L_{y}}$:
-$\hat{\mathbf{v}}^{y}:=\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}^{y\mid L_{y}}\right|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}(\mathbf{u}_{g}-\mathbf{u}_{d}),$ (21)
-where $\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}^{y\mid L_{y}}$.
+(7), we define an estimator $\hat{\mathbf{v}}^{y}$ for the relation vector $\mathbf{v}^{y}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}^{y}$:
+$\hat{\mathbf{v}}^{y}:=\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}^{y}\right|}\sum_{(d,g)\in\mathcal{R}^{y}}(\mathbf{u}_{g}-\mathbf{u}_{d}),$ (20)
+where $\mathrm{E}_{\mathcal{R}^{y}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}^{y}$.
 For easy comparison with Eq.
@@
 (20) is rewritten as the difference of mean vectors:
-$\hat{\mathbf{v}}^{y}=\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\},$ (22)
-$\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{d}.$ (23)
-Given the above, similar to setting G, we perform the analogy tasks in the following two settings.
+$\hat{\mathbf{v}}^{y}=\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{d}\},$ (21)
+$\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y}|}\sum_{(d,g)\in\mathcal{R}^{y}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y}|}\sum_{(d,g)\in\mathcal{R}^{y}}\mathbf{u}_{d}.$ (22)
+Given the above, similar to setting G, we perform the analogy tasks in the following setting.
 Setting Y1.
+Using embeddings trained on PubMed abstracts up to year $y$, we consider analogy tasks in setting G. Thus, if $y$ is the most recent, it simply corresponds to setting G.
+See Supplementary Information 1.2.1 for more details on the setting.
+Global setting to predict unknown relations by year
+Based on analogy tasks in the global setting by year, we consider analogy tasks to predict unknown relations using known relations.
+Specifically, drug-gene relations that appeared up to year $y$ are considered known, while those that appeared after year $y$ are considered unknown.
+We then use embeddings trained on PubMed abstracts up to year $y$, redefine vectors representing known relations, and use these vectors to predict unknown relations.
+In preparation, we define $y_{(d,g)}$ as the year when both drug $d$ and gene $g$ first appeared together in a PubMed abstract.
+We consider $y_{(d,g)}$ as a substitute for the year when the relation $(d,g)$ was first identified.
+By definition, $\max\{y_{d},y_{g}\}\leq y_{(d,g)}$ holds.
+The relation $(d,g)$ is interpreted as either known by year $y$ if $y_{(d,g)}\leq y$ or unknown by year $y$ if $y<y_{(d,g)}$.
+We define two subsets of $\mathcal{R}^{y}$ based on whether $y_{(d,g)}\leq y$ or $y<y_{(d,g)}$.
+To do this, we define two intervals, $L_{y}:=(-\infty,y]$ and $U_{y}:=(y,\infty)$.
+Using $L_{y}$ and $U_{y}$, we define the subsets $\mathcal{R}^{y\mid L_{y}}$ and $\mathcal{R}^{y\mid U_{y}}$ of $\mathcal{R}^{y}$ as follows:
+$\mathcal{R}^{y\mid L_{y}}:=\{(d,g)\in\mathcal{R}^{y}\mid y_{(d,g)}\in L_{y}\},\,\mathcal{R}^{y\mid U_{y}}:=\{(d,g)\in\mathcal{R}^{y}\mid y_{(d,g)}\in U_{y}\}\subset\mathcal{R}^{y}.$ (23)
+The set of “known” relations is expressed as $\mathcal{R}^{y\mid L_{y}}$, and the set of “unknown” relations is expressed as $\mathcal{R}^{y\mid U_{y}}$.
+By definition, $\mathcal{R}^{y\mid L_{y}}\cap\mathcal{R}^{y\mid U_{y}}=\emptyset$ and $\mathcal{R}^{y\mid L_{y}}\cup\mathcal{R}^{y\mid U_{y}}\subset\mathcal{R}^{y}$.
+In analogy tasks, we use “known” $\mathcal{R}^{y\mid L_{y}}$ and then predict the target genes $g$ from a drug $d$ for $(d,g)$ in “unknown” $\mathcal{R}^{y\mid U_{y}}$.
+Using the relation vector $\mathbf{v}^{y\mid L_{y}}$, which represents the drug-gene relations in $\mathcal{R}^{y\mid L_{y}}$, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}^{y}$ to $\mathbf{u}_{d}$:
+$\mathbf{u}_{d}+\mathbf{v}^{y\mid L_{y}}\approx\mathbf{u}_{g}.$ (24)
+Equation (24) corresponds to Eq.
+(19).
+We define the estimator $\hat{\mathbf{v}}^{y\mid L_{y}}$ for the relation vector $\mathbf{v}^{y\mid L_{y}}$ by using $\mathcal{R}^{y\mid L_{y}}$ instead of $\mathcal{R}^{y}$ in the estimator $\hat{\mathbf{v}}^{y}$ in Eq.
+(20).
+Given the above, we perform the analogy tasks in the following setting.
+Setting Y2.
+Using embeddings trained on PubMed abstracts up to year $y$, we consider analogy tasks with a reduced answer set.
 For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct.
 In other words, only new discoveries are counted as correct.
-Setting Y2.
-For the target genes that have drug-gene relations with a drug $d$, these genes are considered correct regardless of whether their relations appeared up to or after year $y$.
-In other words, rediscoveries of known relations are also counted as correct.
-See Supplementary Information 1.2.1 for more details on these settings.
+See Supplementary Information 1.2.2 for more details on the setting.
 Note that $y_{d}$, $y_{g}$, and $y_{(d,g)}$ are defined based on the year the drugs, genes, and drug-gene relations appeared in PubMed abstracts.
@@
 Pathway-wise setting by year
-As a next step, based on the analogy tasks in settings P1, P2, Y1, and Y2, we consider more detailed analogy tasks.
-To do this, we consider the analogy tasks in the pathway-wise setting by year, where drugs and genes are categorized based on pathways in datasets divided by year.
-Similar to settings P1, P2, Y1, and Y2, we perform the analogy tasks in the following four settings.
+As a next step, based on the analogy tasks in settings P1, P2, and Y1, we consider the analogy tasks in the pathway-wise setting by year, where drugs and genes are categorized based on pathways in datasets divided by year.
+To do this, we perform the analogy tasks in the following two settings.
 Setting P1Y1.
+Using embeddings trained on PubMed abstracts up to year $y$, we consider analogy tasks in setting P1. Thus, if $y$ is the most recent, it corresponds to setting P1.
+Setting P2Y1.
+Using embeddings trained on PubMed abstracts up to year $y$, we consider analogy tasks in setting P2. Thus, if $y$ is the most recent, it corresponds to setting P2.
+See Supplementary Information 1.2.3 for details on the analogy tasks and these settings.
+Pathway-wise setting to predict unknown relations by year
+Furthermore, based on the analogy tasks in settings P1, P2, and Y2, we consider analogy tasks to predict unknown relations using known relations in the pathway-wise setting by year.
+To do this, we perform the analogy tasks in the following two settings.
+Setting P1Y2.
+Using embeddings trained on PubMed abstracts up to year $y$, we consider analogy tasks in setting P1 with a reduced answer set.
 For the target genes that have drug-gene relations with a drug $d$, only genes categorized in the same pathway $p$ as $d$, and whose relations appeared after year $y$, are considered correct.
-Setting P1Y2.
-For the target genes that have drug-gene relations with a drug $d$, only genes categorized in the same pathway $p$ as the drug $d$ are considered correct, regardless of whether their relations appeared up to or after year $y$.
-Setting P2Y1.
+Setting P2Y2.
+Using embeddings trained on PubMed abstracts up to year $y$, we consider analogy tasks in setting P2 with a reduced answer set.
 For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct, regardless of whether they are categorized in the same pathway $p$ as the drug $d$ or not.
-Setting P2Y2.
-For the target genes that have drug-gene relations with a drug $d$, these genes are considered correct regardless of whether they are categorized in the same pathway or not, and whether their relations appeared up to or after year $y$.
-See Supplementary Information 1.2.2 for details on the analogy tasks and these settings.
+See Supplementary Information 1.2.4 for details on the analogy tasks and these settings.
 Embeddings
-BioConceptVec Ours
+BioConceptVec Our embeddings
 $|\mathcal{D}|$ 117282 28284
@@
 Datasets
-Corpus
+Corpus.
 Following Chen et al.
@@
 We used about 35 million abstracts up to the year 2023, while they used about 30 million abstracts.
-Drug-gene relations
+Drug-gene relations.
 For drug-gene relations, we used publicly available data from AsuratDB [36], which collects information from various databases including the KEGG [17] database.
 KEGG (Kyoto Encyclopedia of Genes and Genomes) is a comprehensive database system that integrates a wide range of bioinformatics information such as genomics, chemical reactions, and biological pathways.
-Biological Pathways
+Biological Pathways.
 We obtained a list of human pathways from the KEGG API (https://rest.kegg.jp/list/pathway/hsa) for use in our experiments.
@@
 For more details, see Data availability section and Supplementary Information 1.5.
-Baseline
-To evaluate the performance of predicting target genes by adding the relation vector to a drug embedding, we compare it to a simple baseline method.
-As a baseline, we randomly select one gene from the search space for each query.
+Baselines
+To evaluate the performance of predicting target genes by adding the relation vector to a drug embedding, we compare it to baseline methods.
+As a simple baseline, we randomly select one gene from the search space for each query.
 We refer to this baseline as the random baseline.
 In the random baseline, we repeated the experiments 10 times and used the average score as the final result.
+In addition to this, as strong generative model baselines, we used the Generative Pre-trained Transformer (GPT) [37] series to predict the top 10 target genes from a query drug by zero-shot.
+We used the GPT-3.5 [38], GPT-4 [39], and GPT-4o (https://platform.openai.com/docs/models/gpt-4o) models with the temperature hyperparameter set to 00.
+See Supplementary Information 1.6 for more details on the models and prompt template for the predictions.
 Evaluation metrics
@@
 MRR is a statistical measure to evaluate search performance, expressed as the sum of the reciprocals of the ranks at which the predicted results first appear in the answer set.
+Note that since MRR cannot be calculated for the GPT models due to the limited number of prediction candidates output by GPT, only top-1 and top-10 accuracies are calculated.
 When evaluating performance using top-1 accuracy, top-10 accuracy, and MRR, the embeddings of all genes $\mathcal{G}$ (or $\mathcal{G}^{y}$ in settings by year) are centered using the mean embedding $\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\}$ in Eq.
@@
 Since cosine similarity is affected by the origin, centering mitigates this effect.
-In the case of setting G, for example, we actually calculate the value of $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}-\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\},\mathbf{u}_{g}-\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\})$ instead of $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}},\mathbf{u}_{g})$.
+In the case of setting G, for example, we actually calculate the value of $\cos(\mathbf{u}_{d}-\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\}+\hat{\mathbf{v}},\mathbf{u}_{g}-\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\})$ instead of $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}},\mathbf{u}_{g})$.
 Results
 Metric
-Embeddings Setting Method Top1 Top10 MRR
+Model Setting Method Top1 Top10 MRR
 BioConceptVec G Random 0.000 0.000 0.000
@@
 $\hat{\mathbf{v}}_{p}$ 0.522 0.807 0.624
-Ours G Random 0.000 0.000 0.001
+Our embeddings G Random 0.000 0.000 0.001
 $\hat{\mathbf{v}}$ 0.300 0.686 0.426
@@
 $\hat{\mathbf{v}}_{p}$ 0.600 0.880 0.700
+GPT-3.5 Best of all zero-shot 0.653 0.821
+GPT-4 Best of all zero-shot 0.710 0.832
+GPT-4o Best of all zero-shot 0.760 0.887
 Table 3: Gene prediction performance in settings G, P1, and P2.
@@
 Figure 3: Gene prediction performance and the number of queries for each year in settings Y1, Y2, P1Y1, and P1Y2.
-Gene prediction performance in settings G, P1, and P2.
+Gene prediction performance in settings G, P1, and P2
 Table 3 shows the results of experiments with BioConceptVec and our skip-gram embeddings in settings G, P1, and P2.
+For GPT-3.5, GPT-4, and GPT-4o, only the best results are shown; see Supplementary Information 2.1.3 for detailed results.
 In setting G, the prediction by simply adding the global relation vector $\hat{\mathbf{v}}$ resulted in a top-1 accuracy of about 0.3, a top-10 accuracy of over 0.6, and an MRR also over 0.4.
 These results show that the embeddings have the ability to interpret the drug-gene relations.
-Similar to the basic analogy task, these relations were not explicitly provided during the training of the embeddings.
-Previous research [37] has shown that skip-gram achieves a top-1 accuracy of 0.56 on the Microsoft Research Syntactic Analogies Dataset [38].
-Given this, BioConceptVec and our skip-gram embeddings in the analogy tasks for drug-gene pairs show comparable performance to skip-gram in the basic analogy tasks.
+As with the basic analogy task, these relations were not explicitly provided during the training of the embeddings.
 In settings P1 and P2, prediction by adding the pathway-wise relation vector $\hat{\mathbf{v}}_{p}$ showed better performance than in setting G.
@@
 Since the analogy tasks are considered for each pathway, these settings are also likely to make the tasks easier than those in setting G.
-In contrast, the random baseline in settings G, P1, and P2 showed nearly zero scores for all evaluation metrics, suggesting that random prediction is nearly impossible in these analogy tasks.
+The random baseline in settings G, P1, and P2 showed nearly zero scores for all evaluation metrics, suggesting that random prediction is nearly impossible in these analogy tasks.
 This also confirms the superior performance of prediction by adding relation vectors.
+Among the GPT series, GPT-4o performed the best, followed by GPT-4, which outperformed GPT-3.5.
+In top-1 accuracy, the GPT series performed better than predictions made by adding relation vectors.
+However, in top-10 accuracy, the difference becomes smaller.
+Notably, predictions using $\mathbf{v}_{p}$ with our skip-gram embeddings in settings P1 and P2 performed comparably to GPT-4o and outperformed both GPT-3.5 and GPT-4.
 In addition, our skip-gram outperformed BioConceptVec except for the top-1 accuracy in setting G.
-This is probably because BioConceptVec has a dimensionality of 100, while our skip-gram has a dimensionality of 300.
-Note that, as shown in Table 2, the settings of the analogy tasks are not identical for these two embeddings.
+This is probably because BioConceptVec has a dimensionality of 100 and the size of the search space is $|\mathcal{G}|=117282$, while our skip-gram has a dimensionality of 300 and the size of the search space is $|\mathcal{G}|=28284$, as shown in Table 2.
 For the results in Table 3, we used the better estimators in equations (7) and (14), not the naive estimators in equations (5) and (S1), and calculated the cosine similarities after centering the embeddings of all genes $\mathcal{G}$ (or $\mathcal{G}^{y}$).
@@
 Gene prediction performance in settings Y1 and Y2
-Figures 3(a) and 3(b) show the results of experiments with our skip-gram embeddings in settings Y1 and Y2.
+Figures 3(a) and 3(b) show the results of experiments in settings Y1 and Y2 with our skip-gram embeddings trained on the year-specific divisions up to year $y$.
 BioConceptVec embeddings are not appropriate for the settings by year, since they are pretrained on the entire dataset without year-specific divisions.
-In setting Y1, since only genes whose relations appeared after year $y$ are included in the answer set, the tasks are more challenging.
-Therefore, the evaluation metric scores are lower compared to those in setting Y2, where genes whose relations appeared up to year $y$ are also included in the answer set.
-Yet, in setting Y1, the top-1 accuracy from $y=1985$ to $y=2015$ is about 0.1, the top-10 accuracy is over 0.3, and the MRR is about 0.2.
-Although in setting Y1 the relations to be predicted appeared after year $y$ and are not used to calculate $\hat{\mathbf{v}}^{y}$, these results show that adding $\hat{\mathbf{v}}^{y}$ to $\mathbf{u}_{d}$ can predict these relations.
-Setting Y2 is easier than setting Y1, and using the relation vectors results in high evaluation metric scores.
-For example, from $y=1985$ to $y=2015$, the top-1 accuracy reaches about 0.3, and the top-10 accuracy is about 0.7.
-In setting Y1, the number of queries increases initially and then decreases over time, while in setting Y2, the number of queries increases steadily from older to more recent years.
-In addition, the size of the search space is small for older years (e.g., $|\mathcal{G}^{1970}|=353$) and increases for later years (e.g., $|\mathcal{G}^{2020}|=47509$).
-As a result, in both settings Y1 and Y2, the datasets for older years such as $y=1970$, $1975$, and $1980$ have fewer queries and smaller search spaces, resulting in unusually high evaluation metric scores.
-In setting Y1, the datasets for later years have fewer queries and larger search spaces, which tends to produce lower evaluation metric scores.
+Setting Y1 is easier than setting Y2, and results in high evaluation metric scores.
+For example, from $y=1980$ to $y=2023$, the top-1 accuracy reaches about 0.3, and the top-10 accuracy is about 0.6.
+In setting Y1, as in setting G, all relations contained in $\cal{R}$ are considered correct, so all the scores for $y=2023$ are identical to those of setting G.
+For example, a top-10 accuracy of 0.686 is confirmed in both Fig.
+3(a) and Table 3.
+In setting Y2, since only genes whose relations appeared after year $y$ are included in the answer set, the tasks are more challenging.
+Therefore, the evaluation metric scores are lower compared to those in setting Y1, where genes whose relations appeared up to year $y$ are also included in the answer set.
+Yet, in setting Y2, the top-1 accuracy from $y=1985$ to $y=2015$ is about 0.1, the top-10 accuracy is over 0.3, and the MRR is about 0.2.
+Although in setting Y2 the relations to be predicted appeared after year $y$ and are not used to calculate $\hat{\mathbf{v}}^{y\mid L_{y}}$, these results show that adding $\hat{\mathbf{v}}^{y\mid L_{y}}$ to $\mathbf{u}_{d}$ can predict these relations.
+In setting Y1, the number of queries increases steadily from older to more recent years, while in setting Y2, the number of queries increases initially and then decreases over time.
+In addition, the size of the search space is small for older years (e.g., $|\mathcal{G}^{1975}|=725$) and increases for later years (e.g., $|\mathcal{G}^{2020}|=47509$).
+As a result, in both settings Y1 and Y2, the datasets for older years such as $y=1975$, and $1980$ have fewer queries and smaller search spaces, resulting in unusually high evaluation metric scores.
+In setting Y2, the datasets for later years have fewer queries and larger search spaces, which tends to produce lower evaluation metric scores.
 In both settings, the evaluation metric scores for the random baseline are close to zero.
 For more detailed results for settings Y1 and Y2, see Table S12 in Supplementary Information 2.2.1.
-Gene prediction performance in settings P1Y1, P1Y2, P2Y1 and P2Y2.
+Gene prediction performance in settings P1Y1, P2Y1, P1Y2, and P2Y2
 Figures 3(c) and 3(d) show the results of experiments with our skip-gram embeddings in settings P1Y1 and P1Y2.
 The results in settings P1Y1 and P1Y2 not only follow similar trends to those in settings Y1 and Y2 but also consistently show higher evaluation metric scores.
+From the definition of setting P1Y1, the scores for $y=2023$ are identical to those of setting P1.
+For example, a top-10 accuracy of 0.862 is confirmed in both Fig.
+3(c) and Table 3.
 The analogy tasks in settings P1Y1 and P1Y2 use detailed pathway information.
 Therefore, predicting the relations that appeared after year $y$ is easier in settings P1Y1 and P1Y2 than in settings Y1 and Y2.
-In addition, the results for the older years show extremely high evaluation metric scores in settings P1Y1 and P1Y2, similar to those in settings Y1 and Y2.
+In addition, the results for the older years show high evaluation metric scores in settings P1Y1 and P1Y2, similar to those in settings Y1 and Y2.
 The evaluation metric scores of the random baseline are close to zero.
@@
 They are very similar to those in settings P1Y1 and P1Y2.
-For detailed results for settings P1Y1, P1Y2, P2Y1, and P2Y2, see Table S10 in Supplementary Information 2.2.2.
-Discussion
+For detailed results for settings P1Y1, P2Y1, P1Y2, and P2Y2, see Table S15 in Supplementary Information 2.2.2.
+Biological Insights from Predicting Drug-Gene Relations
 Gene Prediction
@@
 Genes whose IDs could not be converted to gene names are indicated by “–”.
-In this section, we investigate the predictions made by adding the relation vector in settings G, P1, and P2.
-For this purpose, we focus on the results using genes and drugs categorized in the ErbB signaling pathway as shown in Table 1 and Fig.
+In this section, we investigate the biological insights obtained through analogy tasks by adding the relation vector in settings G, P1, and P2.
+For this purpose, we focus on genes and drugs categorized in the ErbB signaling pathway as shown in Table 1 and Fig.
 2.
@@
 $\mathcal{D}_{p}$, $\mathcal{G}_{p}$, and $\mathcal{R}_{p}$ of our embeddings are identical to those of the BioConceptVec embeddings in Table 1.
-In Table 4, we list the predicted target genes for drugs that are categorized in the ErbB signaling pathway and whose answer set size in setting G is two or more.
-For all drugs meeting these criteria, at least one answer gene was included in the top 10 predicted target genes for either of the settings.
-In the following, we show that genes that were ranked high in the predictions, even though they were not in the answer set, can be interpreted biologically.
+In Table 4, we have listed the predicted target genes for drugs that are categorized in the ErbB signaling pathway and whose answer set size in setting G is two or more.
+For all drugs meeting these criteria, at least one answer gene was included in the top 10 predicted target genes for any of the settings.
+Therefore, it can be said that the prediction of drug-gene relations through analogy tasks functioned appropriately.
+Additionally, we provide several examples demonstrating that some genes ranked high in the predictions, even though they were not in the answer set, can still be interpreted biologically.
 First, the target genes of Bosutinib are ABL1 and SRC, which are both known to be non-receptor tyrosine kinases (non-RTKs) [40].
-Within the top 10 predicted target genes, TXK [40] and JAK2 [41] can also be categorized as non-RTKs, sharing some properties with ABL1 and SRC.
+Among the top 10 predicted target genes, TXK [41] and JAK2 [42] can also be categorized as non-RTKs, sharing several properties with ABL1 and SRC.
 This implies that the structural and biochemical similarities of these genes may have been reflected in their high-dimensional representations.
 For Masoprocol, we were able to predict the correct target gene EGFR as the fifth prediction in settings P1 and P2.
-However, although EGFR is the sole target gene of Masoprocol according to the information deposited in KEGG, it is also reported to have inhibitory effects towards lipoxygenase activity [42].
-Since ALOX5 is a gene that codes a lipoxygenase, treating this prediction as inaccurate should be considered incorrect.
+However, although EGFR is the sole target gene of Masoprocol according to the information deposited in KEGG, it has also been reported to inhibit lipoxygenase activity [43].
+Since ALOX5 codes for a lipoxygenase, this prediction should not be considered inaccurate.
 The diseases in which Masoprocol is used for its treatment include actinic keratosis [44], and it has been previously known that the mutation of TP53 is involved in the onset of this disease [45, 46].
 Also, the gene that was ranked the highest in this setting, ALOX5, is reported to be one of the transcription targets of TP53 [47].
-Taken together, this information suggests that TP53 is deeply related to the target gene and symptoms of Masoprocol, and therefore, although it may not be a direct target, it was highly ranked in the prediction.
+Together, this information suggests that TP53 is deeply related to the target gene and symptoms of Masoprocol.
+Therefore, although it may not be a direct target, its high rank in the prediction is justified.
 When we set Poziotinib as the query, EGFR and ERBB2 were correctly ranked the highest in settings P1 and P2.
 However, in setting G, ROS1, EML4, and ALK were also found among the highly ranked targets.
-Since Poziotinib was initially developed as an effective drug to treat lung cancer with HER2 mutation [47], it can be interpreted that such context was reflected in the embeddings, as ROS1 and EML4-ALK fusion genes are also characteristic mutations and therapeutic targets in lung cancer [48, 49].
+Since Poziotinib was initially developed as an effective drug to treat lung cancer with HER2 mutation [48], this context appears to be reflected in the embeddings, as ROS1 and EML4-ALK fusion genes are also characteristic mutations and therapeutic targets in lung cancer [49, 50].
 Selumetinib and Trametinib both target MAP2K1 and MAP2K2, and in both cases these genes were found within the top 10 predictions.
 Aside from these target genes, we observed BRAF and PI3KCA placed in higher rank among the predicted genes.
-Considering the history that these drugs were both developed for treating cancers with BRAF V600 mutation and that combination treatment with inhibition of the PI3K-AKT pathway have been explored [50, 51, 52], we assumed that such background was reflected in these results.
-As we explored in this section, it was suggested that the high-dimensional representations of genes and drugs calculated from biomedical texts not only capture the simple drug-gene relations we sought to predict in this study, but also integrate broader higher-order information.
-In this specific analysis, predicted genes are closely related to the genes or symptoms that have direct molecular interactions with the target genes, or those that are alternative therapeutic targets for the target disease.
+Given the history that these drugs were both developed to treat cancers with BRAF V600 mutation and that combination treatment with inhibition of the PI3K-AKT pathway has been explored [51, 52, 53], we assumed that such background was reflected in these results.
+Discussion
+Mikolov et al.
+[54] has shown that skip-gram achieves a top-1 accuracy of 0.56 on the Microsoft Research Syntactic Analogies Dataset [55].
+Given this, the performance of BioConceptVec and our skip-gram embeddings in the analogy tasks for drug-gene pairs, as presented in Table 3, is comparable to that of skip-gram in the basic analogy tasks.
+Also, Tshitoyan et al.
+[18] reported an overall accuracy of 60.1% when using analogy computation of word embedding to predict 29,046 relationships across various concepts in materials science, including element names, crystal symmetries, and magnetic properties.
+Our results for drug-gene relationship prediction show a similar level of accuracy, suggesting that the approach of using analogy computation of word embedding can be effective in the biomedical domain as well.
+This demonstrates the potential of this method across different scientific disciplines.
+However, it’s important to note that the accuracy varies significantly across different types of relationships in the materials science study.
+For instance, predictions for chemical element names showed high accuracy (71.4%), while crystal structure names had lower accuracy (18.7%).
+This variability suggests that the effectiveness of word embedding analogies may depend on the specific type of relationship being predicted.
+In Table 3, predictions using $\mathbf{v}_{p}$ with our skip-gram embeddings in settings P1 and P2 performed on par with GPT-4o and outperformed both GPT-3.5 and GPT-4.
+It is rather surprising that simple vector addition can achieve performance comparable to large language models.
+Given that the GPT series is trained on vast amounts of text data, using relation vector addition is an efficient method.
+Furthermore, the results showing that target genes can be predicted with performance comparable to or better than large language models imply that high-quality information about drug-gene relationships is encoded in the embeddings.
+This also suggests that the computation of analogies may be a fundamental inference principle within large language models.
+In addition, we divided the datasets by year and redefined the vectors representing known relations to predict unknown relations.
+The experimental results presented in Fig.
+3 demonstrate that our approach can predict unknown future relations to some extent.
+The experiments conducted with year-split datasets are a novel attempt not found in existing research, and confirmed the effectiveness of predicting unknown future events.
+In Table 4, predicted genes in the ErbB signaling pathway are closely related to the genes or symptoms that have direct molecular interactions with the target genes, or alternative therapeutic targets for the target disease.
+This may suggest that the high-dimensional representations of genes and drugs calculated from biomedical texts not only capture the simple drug-gene relations we aimed to predict in this study but also integrate broader, higher-order information.
 Conclusions
-In this study, we used the embeddings learned from biological texts and performed analogy tasks for predicting drug-gene relations.
+In this study, we used embeddings learned from biological texts and performed analogy tasks to predict drug-gene relations.
 We defined vectors representing these relations and showed that these vectors can accurately predict the target genes for given drugs.
-Furthermore, we categorized drugs and genes based on biological pathways and then defined vectors representing drug-gene relations for each pathway.
-Analogy computation using these vectors showed more performance improvement.
-In addition, we divided datasets by year and redefined the vectors representing known relations to predict unknown relations.
-The experimental results show that our approach can predict unknown relations to some extent.
+Additionally, we categorized drugs and genes based on biological pathways and defined vectors representing drug-gene relations for each pathway.
+Analogy computations with these vectors showed performance improvement.
+Our analogy computations demonstrated performance comparable to analogy tasks in other fields and predictions by state-of-the-art large language models, reinforcing the effectiveness of relationship prediction through simple vector addition.
+Moreover, the experiments with year-split datasets demonstrated that it is possible to predict unknown future relations.
+Not only were the predictions highly accurate, but the top genes predicted from drugs by our analogy computations were also confirmed to be reasonable from the perspective of biological expertise.
 Data availability
@@
 All authors reviewed the manuscript.
+Additional information
+The authors declare that they have no competing interests.
 Supplementary Information
@@
 (20), we prepare the evaluation of the analogy tasks.
-For $I\in\{(-\infty,\infty),L_{y},U_{y}\}$, using the projection operations in Eq.
+Using the projection operations in Eq.
+(10), we define $D^{y}:=\pi_{\mathcal{D}^{y}}(\mathcal{R}^{y})\subset\mathcal{D}^{y}$ and $G^{y}:=\pi_{\mathcal{G}^{y}}(\mathcal{R}^{y})\subset\mathcal{G}^{y}$.
+Similar to Eq.
+(11), we define $[d]^{y}\subset\mathcal{G}^{y}$ as the set of genes that have drug-gene relations with a drug $d\in D^{y}$, and $[g]^{y}\subset\mathcal{D}^{y}$ as the set of drugs that have drug-gene relations with a gene $g\in G^{y}$.
+These are formally defined as follows:
+$[d]^{y}:=\{g\mid(d,g)\in\mathcal{R}^{y}\}\subset\mathcal{G}^{y},\quad[g]^{y}:=\{d\mid(d,g)\in\mathcal{R}^{y}\}\subset\mathcal{D}^{y}.$ (S3)
+Given the above, we perform the analogy tasks in the following setting.
+Setting Y1.
+In the analogy tasks, the set of answer genes for a query drug $d\in D^{y}$ is $[d]^{y}$.
+The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$ and if $\hat{g}_{d}\in[d]^{y}$, then the prediction is considered correct.
+We define $\hat{g}_{d}^{(k)}$ as the $k$-th ranked $g\in\mathcal{G}^{y}$ based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$.
+For the top-$k$ accuracy, if any of the top $k$ predictions $\hat{g}_{d}^{(1)},\ldots,\hat{g}_{d}^{(k)}\in[d]^{y}$, then the prediction is considered correct.
+1.2.2 Global setting to predict unknown relations by year
+Similar to the estimator $\hat{\mathbf{v}}$ in Eq.
+(7), we define an estimator $\hat{\mathbf{v}}^{y\mid L_{y}}$ for the relation vector $\mathbf{v}^{y\mid L_{y}}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}^{y\mid L_{y}}$:
+$\hat{\mathbf{v}}^{y\mid L_{y}}:=\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}^{y\mid L_{y}}\right|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}(\mathbf{u}_{g}-\mathbf{u}_{d}),$ (S4)
+where $\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}^{y\mid L_{y}}$.
+We define the estimator $\hat{\mathbf{v}}^{y\mid L_{y}}$ in Eq.
+(S4) by using $\mathcal{R}^{y\mid L_{y}}$ instead of $\mathcal{R}^{y}$ in the estimator $\hat{\mathbf{v}}^{y}$ in Eq.
+(20).
+For easy comparison with Eq.
+(5), Eq.
+(S4) is rewritten as the difference of mean vectors:
+$\hat{\mathbf{v}}^{y\mid L_{y}}=\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\},$ (S5)
+$\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{d}.$ (S6)
+Similar to $\hat{\mathbf{v}}$ in Eq.
+(7), to measure the performance of the estimator $\hat{\mathbf{v}}^{y\mid L_{y}}$ in Eq.
+(S4), we prepare the evaluation of the analogy tasks.
+For $I\in\{L_{y},U_{y}\}$, using the projection operations in Eq.
 (10), we define $D^{y\mid I}:=\pi_{\mathcal{D}^{y}}(\mathcal{R}^{y\mid I})\subset\mathcal{D}^{y}$ and $G^{y\mid I}:=\pi_{\mathcal{G}^{y}}(\mathcal{R}^{y\mid I})\subset\mathcal{G}^{y}$.
-Also we define $D^{y}:=D^{y\mid(-\infty,\infty)}$ and $G^{y}:=G^{y\mid(-\infty,\infty)}$.
 Similar to Eq.
-(11), for $I\in\{(-\infty,\infty),L_{y},U_{y}\}$, we define $[d]^{y\mid I}\subset\mathcal{G}^{y}$ as the set of genes that have drug-gene relations with a drug $d\in D^{y\mid I}$, and $[g]^{y\mid I}\subset\mathcal{D}^{y}$ as the set of drugs that have drug-gene relations with a gene $g\in G^{y\mid I}$.
+(11), for $I\in\{L_{y},U_{y}\}$, we define $[d]^{y\mid I}\subset\mathcal{G}^{y}$ as the set of genes that have drug-gene relations with a drug $d\in D^{y\mid I}$, and $[g]^{y\mid I}\subset\mathcal{D}^{y}$ as the set of drugs that have drug-gene relations with a gene $g\in G^{y\mid I}$.
 These are formally defined as follows:
 $[d]^{y\mid I}:=\{g\mid(d,g)\in\mathcal{R}^{y|I}\}\subset\mathcal{G}^{y},\quad[g]^{y\mid I}:=\{d\mid(d,g)\in\mathcal{R}^{y|I}\}\subset\mathcal{D}^{y}.$ (S7)
-In particular, we define $[d]^{y}:=[d]^{y\mid(-\infty,\infty)}$ and $[g]^{y}:=[g]^{y\mid(-\infty,\infty)}$. Given the above, we perform the analogy tasks in the following setting.
-Setting Y1.
+Given the above, we perform the analogy tasks in the following setting.
+Setting Y2.
 For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct.
@@
 The search space is not the set of all genes $\mathcal{G}^{y}$, but the gene set $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$.
-The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct.
-We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$.
+The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y\mid L_{y}},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct.
+We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y\mid L_{y}},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$.
 For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct.
-Setting Y2.
-For the target genes that have drug-gene relations with a drug $d$, these genes are considered correct regardless of whether their relations appeared up to or after year $y$.
-In other words, for a query drug $d\in D^{y}$, the set of answer genes is $[d]^{y}$, where $[d]^{y}=[d]^{y\mid(-\infty,\infty)}$.
-The search space is the set of all genes $\mathcal{G}^{y}$.
-The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y}$, then the prediction is considered correct.
-We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}$.
-For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y}$, then the prediction is considered correct.
-Note the following in settings Y1 and Y2:
-  • In setting Y1, the search space is $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. This ensures that predicted target genes have the relations that appeared only after year $y$.
-  • In setting Y2, to evaluate the ability to predict target genes regardless of whether their relations appeared up to or after year $y$, we consider $(d,g)\in\mathcal{R}^{y}$ instead of $(d,g)\in\mathcal{R}^{y\mid U_{y}}$ in Eq.
-(20).
-BCV Ours
+Note the following in setting Y2:
+  • In setting Y2, the search space is $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. This ensures that predicted target genes have the relations that appeared only after year $y$.
+BioConceptVec Our embeddings
 $|\mathcal{P}|$ 136 129
 $\sum_{p\in\mathcal{P}}|D_{p}|$ 4087 3612
+$\sum_{p\in\mathcal{P}}|G_{p}|$ 1251 1178
 $\sum_{p\in\mathcal{P}}|\mathcal{D}_{p}\cap D|$ 4091 3614
-$\sum_{p\in\mathcal{P}}|G_{p}|$ 1251 1178
 $\sum_{p\in\mathcal{P}}|\mathcal{G}_{p}\cap G|$ 3463 3186
@@
 $\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in G_{p}}\{|[g]_{p}|\}\}$ 5.050 4.847
+$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in\mathcal{D}_{p}\cap D}\{|[d]|\}\}$ 16.346 16.302
+$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in\mathcal{G}_{p}\cap G}\{|[g]|\}\}$ 101.765 91.488
 Table S1: Statistics for settings P1 and P2.
@@
 $\mathcal{R}_{p}^{y}:=\{(d,g)\in\mathcal{R}^{y}\mid d\in\mathcal{D}_{p}^{y},\,g\in\mathcal{G}_{p}^{y}\}\subset\mathcal{D}_{p}^{y}\times\mathcal{G}_{p}^{y}.$ (S8)
-For $(d,g)\in\mathcal{R}_{p}^{y}$, we divide $\mathcal{R}_{p}^{y}$ into two subsets based on whether $y_{(d,g)}\leq y$ or $y<y_{(d,g)}$.
-Similar to $\mathcal{R}^{y\mid I}$ in Eq.
-(19), for an interval $I$, we define the set $\mathcal{R}_{p}^{y\mid I}$, which consists of all $(d,g)\in\mathcal{R}_{p}^{y}$ such that $y_{(d,g)}\in I$, as follows:
-$\mathcal{R}^{y|I}_{p}:=\{(d,g)\in\mathcal{R}_{p}^{y}\mid y_{(d,g)}\in I\}.$ (S5)
-Using $L_{y}=(-\infty,y]$ and $U_{y}=(y,\infty)$, the set of drug-gene pairs that satisfies $y_{(d,g)}\leq y$ is $\mathcal{R}_{p}^{y\mid L_{y}}$, and the set of drug-gene pairs that satisfies $y<y_{(d,g)}$ is $\mathcal{R}_{p}^{y\mid U_{y}}$.
-By definition, $\mathcal{R}_{p}^{y\mid L_{y}}\cap\mathcal{R}_{p}^{y\mid U_{y}}=\emptyset$ and $\mathcal{R}_{p}^{y\mid L_{y}}\cup\mathcal{R}_{p}^{y\mid U_{y}}=\mathcal{R}_{p}^{y\mid(-\infty,\infty)}=\mathcal{R}_{p}^{y}$.
-In analogy tasks, we use “known” $\mathcal{R}_{p}^{y\mid L_{y}}$ and then predict the target genes $g$ from a drug $d$ for $(d,g)$ in “unknown” $\mathcal{R}_{p}^{y\mid U_{y}}$.
-Using the vector $\mathbf{v}_{p}^{y}$, which represents the drug-gene relations and is derived from $\mathcal{R}_{p}^{y\mid L_{y}}$, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}_{p}^{y}$ to $\mathbf{u}_{d}$:
+For drug-gene pairs $(d,g)\in\mathcal{R}_{p}^{y}$, we consider the analogy tasks for predicting the target genes $g$ from a drug $d$.
+To solve these analogy tasks, we use the relation vector $\mathbf{v}_{p}^{y}$, which represents the relation between drugs and target genes categorized in the same pathway $p$.
+We predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}_{p}^{y}$ to $\mathbf{u}_{d}$:
 $\mathbf{u}_{d}+\mathbf{v}_{p}^{y}\approx\mathbf{u}_{g}.$ (S9)
 Similar to equations (13) and (19), Eq. (S9) corresponds to Eq. (4). Therefore, an estimator $\hat{\mathbf{v}}_{p}^{y}$ for the relation vector $\mathbf{v}_{p}^{y}$ in Eq. (S9) is defined as follows:
-$\hat{\mathbf{v}}_{p}^{y}:=\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}_{p}^{y\mid L_{y}}\right|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}(\mathbf{u}_{g}-\mathbf{u}_{d}).$ (S7)
+$\hat{\mathbf{v}}_{p}^{y}:=\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}_{p}^{y}\right|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}(\mathbf{u}_{g}-\mathbf{u}_{d}).$ (S10)
+where $\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}_{p}^{y}$.
+Similar to Eq.
+(7), Eq.
+(S10) defines the estimator $\hat{\mathbf{v}}_{p}^{y}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}_{p}^{y}$.
+For easy comparison with Eq.
+(5), Eq.
+(S10) is rewritten as the difference of mean vectors:
+$\hat{\mathbf{v}}_{p}^{y}=\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{d}\},$ (S11)
+$\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}\mathbf{u}_{d}.$ (S12)
+Similar to $\hat{\mathbf{v}}$ in Eq.
+(7), to measure the performance of the estimator $\hat{\mathbf{v}}_{p}^{y}$ in Eq.
+(S10), we prepare the evaluation of the analogy tasks.
+Using the projection operations in Eq.
+(10), we define $D_{p}^{y}:=\pi_{\mathcal{D}_{p}^{y}}(\mathcal{R}_{p}^{y})$ and $G_{p}^{y}:=\pi_{\mathcal{G}_{p}^{y}}(\mathcal{R}_{p}^{y})$.
+Similar to Eq.
+(11), we define $[d]_{p}^{y}\subset\mathcal{G}_{p}^{y}$ as the set of genes that have drug-gene relations with a drug $d\in D_{p}^{y}$, and $[g]_{p}^{y}\subset\mathcal{D}_{p}^{y}$ as the set of drugs that have drug-gene relations with a gene $g\in G_{p}^{y}$.
+These are formally defined as follows:
+$[d]_{p}^{y}:=\{g\mid(d,g)\in\mathcal{R}_{p}^{y}\}\subset\mathcal{G}_{p}^{y},\quad[g]_{p}^{y}:=\{d\mid(d,g)\in\mathcal{R}_{p}^{y}\}\subset\mathcal{D}_{p}^{y}.$ (S13)
+Similar to settings P1, P2, Y1, we perform the analogy tasks in the following four settings.
+Setting P1Y1.
+For the target genes that have drug-gene relations with a drug $d$, only genes categorized in the same pathway $p$ as the drug $d$ are considered correct.
+In other words, for a query drug $d\in D_{p}^{y}$, the set of answer genes is $[d]_{p}^{y}$.
+The search space is the set of all genes $\mathcal{G}^{y}$, not limited to $\mathcal{G}_{p}^{y}$, the set of genes categorized in the pathway $p$.
+The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]_{p}^{y}$, then the prediction is considered correct.
+We define $\hat{g}_{d}^{(k)}$ as the $k$-th ranked $g\in\mathcal{G}^{y}$ based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$.
+For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]_{p}^{y}$, then the prediction is considered correct.
+Setting P2Y1.
+The gene predictions $\hat{g}_{d}$ and $\hat{g}_{d}^{(k)}$ are defined exactly the same as those in setting P1Y1, but the answer genes are defined the same as in setting Y1.
+That is, for the target genes that have drug-gene relations with a drug $d$, genes are considered correct regardless of whether they are categorized in the same pathway $p$ as the drug $d$ or not.
+In other words, for a query drug $d\in D^{y}$, the set of answer genes is $[d]^{y}$, and the prediction is considered correct if $\hat{g}_{d}\in[d]^{y}$.
+For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y}$, then the prediction is considered correct.
+Note that the experiment is performed for $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ for each $p$.
+1.2.4 Pathway-wise setting to predict unknown relations by year
+For $(d,g)\in\mathcal{R}_{p}^{y}$, we define two subsets of $\mathcal{R}_{p}^{y}$ based on whether $y_{(d,g)}\leq y$ or $y<y_{(d,g)}$.
+Similar to $\mathcal{R}^{y\mid L_{y}}$ and $\mathcal{R}^{y\mid U_{y}}$ in Eq.
+(23), using $L_{y}=(-\infty,y]$ and $U_{y}=(y,\infty)$, we define the set $\mathcal{R}_{p}^{y\mid L_{y}}$ and $\mathcal{R}_{p}^{y\mid U_{y}}$ as follows:
+$\mathcal{R}^{y|L_{y}}_{p}:=\{(d,g)\in\mathcal{R}_{p}^{y}\mid y_{(d,g)}\in L_{y}\},\mathcal{R}^{y|U_{y}}_{p}:=\{(d,g)\in\mathcal{R}_{p}^{y}\mid y_{(d,g)}\in U_{y}\}\subset\mathcal{R}_{p}^{y}.$ (S14)
+The set of drug-gene pairs that satisfies $y_{(d,g)}\leq y$ is $\mathcal{R}_{p}^{y\mid L_{y}}$, and the set of drug-gene pairs that satisfies $y<y_{(d,g)}$ is $\mathcal{R}_{p}^{y\mid U_{y}}$.
+By definition, $\mathcal{R}_{p}^{y\mid L_{y}}\cap\mathcal{R}_{p}^{y\mid U_{y}}=\emptyset$ and $\mathcal{R}^{y|L_{y}}_{p}\cup\mathcal{R}^{y|U_{y}}_{p}\subset\mathcal{R}_{p}^{y}$.
+In analogy tasks, we use “known” $\mathcal{R}_{p}^{y\mid L_{y}}$ and then predict the target genes $g$ from a drug $d$ for $(d,g)$ in “unknown” $\mathcal{R}_{p}^{y\mid U_{y}}$.
+Using the vector $\mathbf{v}_{p}^{y\mid L_{y}}$, which represents the drug-gene relations and is derived from $\mathcal{R}_{p}^{y\mid L_{y}}$, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}_{p}^{y\mid L_{y}}$ to $\mathbf{u}_{d}$:
+$\mathbf{u}_{d}+\mathbf{v}_{p}^{y\mid L_{y}}\approx\mathbf{u}_{g}.$ (S15)
+Similar to equations (13) and (24), Eq.
+(S15) corresponds to Eq.
+(4).
+Therefore, an estimator $\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ for the relation vector $\mathbf{v}_{p}^{y\mid L_{y}}$ in Eq.
+(S15) is defined as follows:
+$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}:=\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}_{p}^{y\mid L_{y}}\right|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}(\mathbf{u}_{g}-\mathbf{u}_{d}).$ (S16)
 where $\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}_{p}^{y\mid L_{y}}$.
@@
 (7), Eq.
-(S7) defines the estimator $\hat{\mathbf{v}}_{p}^{y}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}$.
+(S16) defines the estimator $\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}$.
 For easy comparison with Eq.
@@
 (S16) is rewritten as the difference of mean vectors:
-$\hat{\mathbf{v}}_{p}^{y}=\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{d}\},$ (S8)
+$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}=\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{d}\},$ (S17)
 $\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}\mathbf{u}_{d}.$ (S18)
 Similar to $\hat{\mathbf{v}}$ in Eq.
-(7), to measure the performance of the estimator $\hat{\mathbf{v}}_{p}^{y}$ in Eq.
+(7), to measure the performance of the estimator $\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ in Eq.
 (S16), we prepare the evaluation of the analogy tasks.
-For $I\in\{(-\infty,\infty),L_{y},U_{y}\}$, using the projection operations in Eq.
+For $I\in\{L_{y},U_{y}\}$, using the projection operations in Eq.
 (10), we define $D_{p}^{y\mid I}:=\pi_{\mathcal{D}_{p}^{y}}(\mathcal{R}_{p}^{y\mid I})$ and $G_{p}^{y\mid I}:=\pi_{\mathcal{G}_{p}^{y}}(\mathcal{R}_{p}^{y\mid I})$.
-Also we define $D_{p}^{y}:=D_{p}^{y\mid(-\infty,\infty)}$ and $G_{p}^{y}:=G_{p}^{y\mid(-\infty,\infty)}$.
 Similar to Eq.
-(11), for $I\in\{(-\infty,\infty),L_{y},U_{y}\}$, we define $[d]_{p}^{y\mid I}\subset\mathcal{G}_{p}^{y}$ as the set of genes that have drug-gene relations with a drug $d\in D_{p}^{y\mid I}$, and $[g]_{p}^{y\mid I}\subset\mathcal{D}_{p}^{y}$ as the set of drugs that have drug-gene relations with a gene $g\in G_{p}^{y\mid I}$.
+(11), for $I\in\{L_{y},U_{y}\}$, we define $[d]_{p}^{y\mid I}\subset\mathcal{G}_{p}^{y}$ as the set of genes that have drug-gene relations with a drug $d\in D_{p}^{y\mid I}$, and $[g]_{p}^{y\mid I}\subset\mathcal{D}_{p}^{y}$ as the set of drugs that have drug-gene relations with a gene $g\in G_{p}^{y\mid I}$.
 These are formally defined as follows:
 $[d]_{p}^{y\mid I}:=\{g\mid(d,g)\in\mathcal{R}_{p}^{y\mid I}\}\subset\mathcal{G}_{p}^{y},\quad[g]_{p}^{y\mid I}:=\{d\mid(d,g)\in\mathcal{R}_{p}^{y\mid I}\}\subset\mathcal{D}_{p}^{y}.$ (S19)
-In particular, we define $[d]_{p}^{y}:=[d]_{p}^{y\mid(-\infty,\infty)}$ and $[g]_{p}^{y}:=[g]_{p}^{y\mid(-\infty,\infty)}$.
-Similar to settings P1, P2, Y1, and Y2, we perform the analogy tasks in the following four settings.
-Setting P1Y1.
+Similar to settings P1, P2, Y2, we perform the analogy tasks in the following four settings.
+Setting P1Y2.
 For the target genes that have drug-gene relations with a drug $d$, only genes that are categorized in the same pathway $p$ as $d$ and whose relations appeared after year $y$ are considered correct.
@@
 The search space is not the set of all genes $\mathcal{G}^{y}$, but the gene set $\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$.
-The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]_{p}^{y\mid U_{y}}$, then the prediction is considered correct.
-We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$.
+The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y\mid L_{y}},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]_{p}^{y\mid U_{y}}$, then the prediction is considered correct.
+We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y\mid L_{y}},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$.
 For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]_{p}^{y\mid U_{y}}$, then the prediction is considered correct.
-Setting P2Y1.
+Setting P2Y2.
 For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct, regardless of whether they are categorized in the same pathway $p$ as the drug $d$ or not.
@@
 The search space is not the set of all genes $\mathcal{G}^{y}$, but the gene set $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$.
-The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct.
-We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$.
+The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y\mid L_{y}},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct.
+We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y\mid L_{y}},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$.
 For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y\mid U_{y}}$, then the prediction is considered correct.
-Setting P1Y2.
-For the target genes that have drug-gene relations with a drug $d$, only genes categorized in the same pathway as the drug $d$ are considered correct, regardless of whether their relations appeared up to or after year $y$.
-In other words, for a query drug $d\in D_{p}^{y}$, the set of answer genes is $[d]_{p}^{y}$.
-The search space is the set of all genes $\mathcal{G}^{y}$.
-The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]_{p}^{y}$, then the prediction is considered correct.
-We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}$.
-For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]_{p}^{y}$, then the prediction is considered correct.
-Setting P2Y2.
-For the target genes that have drug-gene relations with a drug $d$, these genes are considered correct regardless of whether they are categorized in the same pathway or not and whether their relations appeared up to or after year $y$.
-In other words, for a query drug $d\in\mathcal{D}_{p}^{y}\cap D^{y}$, the set of answer genes is $[d]^{y}$.
-The search space is the set of all genes $\mathcal{G}^{y}$.
-The predicted gene is $\hat{g}_{d}=\operatorname{argmax}_{g\in\mathcal{G}^{y}}\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$, and if $\hat{g}_{d}\in[d]^{y}$, then the prediction is considered correct.
-We define $\hat{g}_{d}^{(k)}$ as the $k$-th predicted gene, based on $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}_{p}^{y},\mathbf{u}_{g})$ for $g\in\mathcal{G}^{y}$.
-For the top-$k$ accuracy, if $\hat{g}_{d}^{(k)}\in[d]^{y}$, then the prediction is considered correct.
-Note the following in settings P1Y1, P1Y2, P2Y1 and P2Y2:
-  • In setting P1Y1, the search space is $\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$.
+Note the following in settings P1Y2 and P2Y2:
+  • In setting P1Y2, the search space is $\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$.
 This ensures that predicted target genes, which are categorized in the same pathway $p$ as the drug $d$, have the relations that appeared only after year $y$.
-  • In setting P2Y1, since $D_{p}^{y\mid U_{y}}\subset\mathcal{D}_{p}^{y}$ and $D_{p}^{y\mid U_{y}}=\pi_{\mathcal{D}_{p}^{y}}\left(\mathcal{R}_{p}^{y\mid U_{y}}\right)=\pi_{\mathcal{D}^{y}}\left(\mathcal{R}_{p}^{y\mid U_{y}}\right)\subset\pi_{\mathcal{D}^{y}}\left(\mathcal{R}^{y\mid U_{y}}\right)=D^{y\mid U_{y}}$, it follows that $D_{p}^{y\mid U_{y}}\subset\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$.
+  • In setting P2Y2, since $D_{p}^{y\mid U_{y}}\subset\mathcal{D}_{p}^{y}$ and $D_{p}^{y\mid U_{y}}=\pi_{\mathcal{D}_{p}^{y}}\left(\mathcal{R}_{p}^{y\mid U_{y}}\right)=\pi_{\mathcal{D}^{y}}\left(\mathcal{R}_{p}^{y\mid U_{y}}\right)\subset\pi_{\mathcal{D}^{y}}\left(\mathcal{R}^{y\mid U_{y}}\right)=D^{y\mid U_{y}}$, it follows that $D_{p}^{y\mid U_{y}}\subset\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$.
 For $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$, there may be $g\in[d]^{y\mid U_{y}}\setminus[d]_{p}^{y\mid U_{y}}$.
@@
 This ensure that predicted target genes have the relations that appeared only after year $y$.
-  • In setting P1Y2, for a query drug $d\in D_{p}^{y}$, to evaluate the ability to predict target genes categorized in the same pathway $p$ as the drug $d$, regardless of whether their relations appeared up to or after year $y$, we consider $(d,g)\in\mathcal{R}_{p}^{y}$ instead of $(d,g)\in\mathcal{R}_{p}^{y\mid U_{y}}$ in Eq.
-(20).
-  • In setting P2Y2, since $D_{p}^{y}\subset\mathcal{D}_{p}^{y}$ and $D_{p}^{y}=\pi_{\mathcal{D}_{p}^{y}}\left(\mathcal{R}_{p}^{y}\right)=\pi_{\mathcal{D}^{y}}\left(\mathcal{R}_{p}^{y}\right)\subset\pi_{\mathcal{D}^{y}}\left(\mathcal{R}^{y}\right)=D^{y}$, it follows that $D_{p}^{y}\subset\mathcal{D}_{p}^{y}\cap D^{y}$.
-For $d\in\mathcal{D}_{p}^{y}\cap D^{y}$, there may be $g\in[d]^{y}\setminus[d]_{p}^{y}$.
-Therefore, in Eq.
-(S6), we consider $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ and $g\in[d]^{y}$, instead of $(d,g)\in\mathcal{R}_{p}^{y\mid U_{y}}$.
 1.3 Comparison of experimental settings
@@
 P2 $d\in\mathcal{D}_{p}\cap D$ $[d]$ $\mathcal{G}$
-Y1 $d\in D^{y\mid U_{y}}$ $[d]^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$
-Y2 $d\in D^{y}$ $[d]^{y}$ $\mathcal{G}^{y}$
-P1Y1 $d\in D_{p}^{y\mid U_{y}}$ $[d]_{p}^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$
-P2Y1 $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ $[d]^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$
-P1Y2 $d\in D_{p}^{y}$ $[d]_{p}^{y}$ $\mathcal{G}^{y}$
-P2Y2 $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ $[d]^{y}$ $\mathcal{G}^{y}$
+Y1 $d\in D^{y}$ $[d]^{y}$ $\mathcal{G}^{y}$
+Y2 $d\in D^{y\mid U_{y}}$ $[d]^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$
+P1Y1 $d\in D_{p}^{y}$ $[d]_{p}^{y}$ $\mathcal{G}^{y}$
+P2Y1 $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ $[d]^{y}$ $\mathcal{G}^{y}$
+P1Y2 $d\in D_{p}^{y\mid U_{y}}$ $[d]_{p}^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]_{p}^{y\mid L_{y}}$
+P2Y2 $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ $[d]^{y\mid U_{y}}$ $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$
 Table S2: Query, answer set, and search space for each setting for predicting genes from drugs.
@@
 ${}^{{}^{\prime}}$ $g\in\mathcal{G}_{p}\cap G$ $[g]$ $\mathcal{D}$
-${}^{{}^{\prime}}$ $g\in G^{y\mid U_{y}}$ $[g]^{y\mid U_{y}}$ $\mathcal{D}^{y}\setminus[g]^{y\mid L_{y}}$
-${}^{{}^{\prime}}$ $g\in G^{y}$ $[g]^{y}$ $\mathcal{D}^{y}$
 Table S3: Query, answer set, and search space for each setting for predicting drugs from genes.
@@
 Table S4: Hyperparameter for our skip-gram.
+Model Version
+GPT-3.5 gpt-3.5-turbo-0125
+GPT-4 gpt-4-turbo-2024-04-09
+GPT-4o gpt-4o-2024-05-13
+Table S5: Version of each GPT model.
+Role Prompt
+system You are a biologist specializing in drug-target interactions. Your responses should be formatted as JSON, with data listed under the key ‘target_genes’.
+user List the top $10$ potential target genes for the drug drug name.
+Table S6: Prompt template used for the GPT series to predict the top 10 target genes for each query drug. Here, drug name is a specific name such as Bosutinib.
 Table S2 shows the settings for predicting genes from drugs, as used in the main text.
 Similarly, Table S3 shows the settings for predicting drugs from genes.
-In Table S3, we adapted the notations used for predicting genes from drugs in Table S2 to those used for predicting drugs from genes, denoting them as ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, and ${}^{{}^{\prime}}$.
+In Table S3, we adapted the notations used for predicting genes from drugs in Table S2 to those used for predicting drugs from genes, denoting them as ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, and ${}^{{}^{\prime}}$.
 1.4 Embeddings
@@
 We also applied this conversion to drugs in the vocabulary of our trained skip-gram model.
-Embeddings Setting Method Centering Top1 Top10 MRR
-Ours G Random 0.000 0.000 0.001
+Model Setting Method Centering Top1 Top10 MRR
+Our embeddings G Random 0.000 0.000 0.001
 $\hat{\mathbf{v}}_{\text{naive}}$ 0.067 0.197 0.111
@@
 Metric
-Embed. Setting Method Top1 Top10 MRR
-BCV ${}^{{}^{\prime}}$ Random 0.000 0.001 0.001
+Model Setting Method Top1 Top10 MRR
+BioConceptVec ${}^{{}^{\prime}}$ Random 0.000 0.001 0.001
 $\hat{\mathbf{v}}^{\prime}$ 0.233 0.560 0.345
@@
 $\hat{\mathbf{v}}_{p}^{\prime}$ 0.248 0.507 0.337
-Ours ${}^{{}^{\prime}}$ Random 0.000 0.003 0.002
+Our embeddings ${}^{{}^{\prime}}$ Random 0.000 0.003 0.002
 $\hat{\mathbf{v}}^{\prime}$ 0.240 0.577 0.351
@@
 Table S8: Gene prediction performance in settings G, P1, and P2.
+Metric
+Vocabulary Setting LLM Top1 Top10
+BioConceptVec G GPT-3.5 0.611 0.785
+GPT-4 0.665 0.795
+GPT-4o 0.718 0.848
+P1 GPT-3.5 0.581 0.780
+GPT-4 0.633 0.792
+GPT-4o 0.689 0.855
+P2 GPT-3.5 0.627 0.803
+GPT-4 0.685 0.813
+GPT-4o 0.742 0.873
+Our embeddings G GPT-3.5 0.637 0.804
+GPT-4 0.693 0.818
+GPT-4o 0.741 0.867
+P1 GPT-3.5 0.604 0.796
+GPT-4 0.654 0.808
+GPT-4o 0.703 0.868
+P2 GPT-3.5 0.653 0.821
+GPT-4 0.710 0.832
+GPT-4o 0.760 0.887
+Table S9: Gene prediction performance for GPT models.
 Year
-1970 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
-$|\mathcal{D}^{y}|$ 2268 3275 5059 7125 9387 11748 14044 16827 19758 22717 25431
-$|\mathcal{G}^{y}|$ 353 725 1782 3413 5976 10533 16949 24784 32264 39961 47509
-$\left|\mathcal{R}^{y\mid L_{y}}\right|$ 12 37 117 256 499 859 1269 1637 1978 2354 2603
-$\left|\mathcal{R}^{y\mid U_{y}}\right|$ 13 53 183 296 344 515 473 379 281 189 61
-$\left|D^{y\mid U_{y}}\right|$ 12 48 154 272 288 369 325 256 189 131 48
-$|D^{y}|$ 22 79 246 475 663 893 1047 1151 1269 1423 1495
-$\left|G^{y\mid U_{y}}\right|$ 6 14 34 55 89 137 173 153 140 115 44
-$|G^{y}|$ 8 20 46 74 122 215 325 398 477 540 570
-$\mathrm{E}_{d\in D^{y\mid L_{y}}}\left\{\left|[d]^{y\mid L_{y}}\right|\right\}$ 1.091 1.028 1.104 1.133 1.188 1.372 1.500 1.613 1.665 1.722 1.760
-$\mathrm{E}_{d\in D^{y\mid U_{y}}}\left\{\left|[d]^{y\mid U_{y}}\right|\right\}$ 1.083 1.104 1.188 1.088 1.194 1.396 1.455 1.480 1.487 1.443 1.271
-$\mathrm{E}_{g\in G^{y\mid L_{y}}}\left\{\left|[g]^{y\mid L_{y}}\right|\right\}$ 3.000 2.467 4.179 5.020 5.940 5.335 4.789 4.547 4.425 4.467 4.591
-$\mathrm{E}_{g\in G^{y\mid U_{y}}}\left\{\left|[g]^{y\mid U_{y}}\right|\right\}$ 2.167 3.786 5.382 5.382 3.865 3.759 2.734 2.477 2.007 1.643 1.386
-Table S7: Statistics for settings Y1 and Y2.
+1975 1980 1985 1990 1995 2000 2005 2010 2015 2020 2023
+$|\mathcal{D}^{y}|$ 3275 5059 7125 9387 11748 14044 16827 19758 22717 25431 28284
+$|\mathcal{G}^{y}|$ 725 1782 3413 5976 10533 16949 24784 32264 39961 47509 51057
+$|\mathcal{R}^{y}|$ 128 429 807 1251 2337 3282 4054 4849 5551 5909 5968
+$|D^{y}|$ 104 343 666 904 1253 1436 1566 1708 1875 1961 1980
+$|G^{y}|$ 21 47 75 125 228 348 430 521 590 630 634
+$\mathrm{E}_{d\in D^{y}}\{|[d]|\}$ 1.231 1.251 1.212 1.384 1.865 2.286 2.589 2.839 2.961 3.013 3.014
+$\mathrm{E}_{g\in G^{y}}\{|[g]|\}$ 6.095 9.128 10.760 10.008 10.250 9.431 9.428 9.307 9.408 9.379 9.413
+Table S10: Statistics for setting Y1.
+Year
+1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
+$\left|\mathcal{R}^{y\mid L_{y}}\right|$ 37 117 256 499 859 1269 1637 1978 2354 2603
+$\left|\mathcal{R}^{y\mid U_{y}}\right|$ 53 183 296 344 515 473 379 281 189 61
+$\left|D^{y\mid U_{y}}\right|$ 48 154 272 288 369 325 256 189 131 48
+$\left|G^{y\mid U_{y}}\right|$ 14 34 55 89 137 173 153 140 115 44
+$\mathrm{E}_{d\in D^{y\mid L_{y}}}\left\{\left|[d]^{y\mid L_{y}}\right|\right\}$ 1.028 1.104 1.133 1.188 1.372 1.500 1.613 1.665 1.722 1.760
+$\mathrm{E}_{d\in D^{y\mid U_{y}}}\left\{\left|[d]^{y\mid U_{y}}\right|\right\}$ 1.104 1.188 1.088 1.194 1.396 1.455 1.480 1.487 1.443 1.271
+$\mathrm{E}_{g\in G^{y\mid L_{y}}}\left\{\left|[g]^{y\mid L_{y}}\right|\right\}$ 2.467 4.179 5.020 5.940 5.335 4.789 4.547 4.425 4.467 4.591
+$\mathrm{E}_{g\in G^{y\mid U_{y}}}\left\{\left|[g]^{y\mid U_{y}}\right|\right\}$ 3.786 5.382 5.382 3.865 3.759 2.734 2.477 2.007 1.643 1.386
+Table S11: Statistics for setting Y2.
 1.5.2 Conversion from gene ID to gene name
@@
 Year
-Metric Setting Method 1970 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
-Top1 Y1 Random 0.000 0.000 0.003 0.000 0.000 0.001 0.000 0.000 0.000 0.000 0.000
-$\hat{\mathbf{v}}^{y}$ 0.250 0.333 0.208 0.165 0.132 0.076 0.092 0.062 0.090 0.107 0.021
-Y2 Random 0.005 0.005 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
-$\hat{\mathbf{v}}^{y}$ 0.500 0.481 0.350 0.318 0.305 0.299 0.328 0.326 0.318 0.376 0.318
-Top10 Y1 Random 0.008 0.010 0.008 0.002 0.003 0.001 0.000 0.001 0.001 0.000 0.000
-$\hat{\mathbf{v}}^{y}$ 0.500 0.625 0.487 0.438 0.410 0.347 0.397 0.352 0.344 0.328 0.271
-Y2 Random 0.041 0.022 0.007 0.003 0.002 0.001 0.001 0.001 0.000 0.000 0.000
-$\hat{\mathbf{v}}^{y}$ 0.727 0.785 0.699 0.657 0.665 0.679 0.735 0.744 0.702 0.773 0.764
-MRR Y1 Random 0.012 0.007 0.005 0.003 0.002 0.002 0.001 0.001 0.000 0.000 0.000
-$\hat{\mathbf{v}}^{y}$ 0.331 0.438 0.304 0.253 0.214 0.163 0.182 0.147 0.173 0.188 0.103
-Y2 Random 0.017 0.010 0.006 0.003 0.002 0.002 0.001 0.001 0.001 0.000 0.000
-$\hat{\mathbf{v}}^{y}$ 0.590 0.592 0.476 0.434 0.425 0.429 0.468 0.463 0.451 0.513 0.469
+Metric Setting Method 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020 2023
+Top1 Y1 Random 0.001 0.000 0.000 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+$\hat{\mathbf{v}}^{y}$ 0.365 0.262 0.249 0.236 0.247 0.272 0.275 0.287 0.323 0.301 0.300
+Y2 Random 0.000 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+$\hat{\mathbf{v}}^{y\mid L_{y}}$ 0.333 0.208 0.165 0.132 0.076 0.092 0.062 0.090 0.107 0.021
+Top10 Y1 Random 0.016 0.007 0.003 0.002 0.002 0.001 0.001 0.001 0.001 0.001 0.000
+$\hat{\mathbf{v}}^{y}$ 0.837 0.577 0.563 0.562 0.578 0.652 0.691 0.651 0.714 0.693 0.686
+Y2 Random 0.015 0.003 0.002 0.003 0.001 0.001 0.000 0.001 0.000 0.000
+$\hat{\mathbf{v}}^{y\mid L_{y}}$ 0.625 0.487 0.438 0.410 0.347 0.397 0.352 0.344 0.328 0.271
+MRR Y1 Random 0.011 0.005 0.003 0.002 0.001 0.001 0.001 0.001 0.001 0.001 0.000
+$\hat{\mathbf{v}}^{y}$ 0.534 0.370 0.353 0.346 0.363 0.399 0.410 0.410 0.455 0.430 0.426
+Y2 Random 0.008 0.006 0.003 0.002 0.001 0.001 0.000 0.000 0.000 0.000
+$\hat{\mathbf{v}}^{y\mid L_{y}}$ 0.438 0.304 0.253 0.214 0.163 0.182 0.147 0.173 0.188 0.103
 Table S12: Gene prediction performance in settings Y1 and Y2.
 Year
-1970 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
-$\sum_{p\in\mathcal{P}}\left|D_{p}^{y\mid U_{y}}\right|$ 14 104 314 507 521 601 526 439 312 201 67
-$\sum_{p\in\mathcal{P}}|\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}|$ 16 115 363 572 573 732 644 529 401 264 98
-$\sum_{p\in\mathcal{P}}|D_{p}^{y}|$ 26 156 513 959 1363 1728 1999 2205 2419 2711 2821
-$\sum_{p\in\mathcal{P}}|\mathcal{D}_{p}^{y}\cap D^{y}|$ 30 174 566 1028 1419 1794 2068 2263 2480 2773 2886
-$\sum_{p\in\mathcal{P}}\left|G_{p}^{y\mid U_{y}}\right|$ 6 29 88 129 195 305 326 269 228 168 62
-$\sum_{p\in\mathcal{P}}|\mathcal{G}_{p}^{y}\cap G^{y\mid U_{y}}|$ 15 50 177 281 520 935 1087 859 774 682 275
-$\sum_{p\in\mathcal{P}}|G_{p}^{y}|$ 6 38 110 170 281 484 658 772 925 1066 1124
-$\sum_{p\in\mathcal{P}}|\mathcal{G}_{p}^{y}\cap G^{y}|$ 18 68 207 367 652 1356 2015 2359 3087 3518 3627
-$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{d\in D_{p}}\left\{\left|[d]_{p}^{y\mid L_{y}}\right|\right\}\right\}$ 4.000 2.500 5.921 8.148 11.407 12.667 14.319 14.649 14.631 15.465 15.977
-$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{d\in D_{p}}\left\{|[d]_{p}^{y\mid U_{y}}|\right\}\right\}$ 2.333 4.160 6.038 7.243 6.432 7.805 6.188 5.776 5.032 4.467 3.190
-$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{g\in G_{p}}\left\{\left|[g]_{p}^{y\mid L_{y}}\right|\right\}\right\}$ 1.000 1.136 1.789 1.918 2.444 3.441 4.483 5.216 5.529 6.171 6.400
-$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{g\in G_{p}}\left\{\left|[g]_{p}^{y\mid U_{y}}\right|\right\}\right\}$ 1.000 1.160 1.692 1.843 2.407 3.961 3.835 3.539 3.677 3.733 2.952
-Table S9: Statistics for settings P1Y1, P1Y2, P2Y1, and P2Y2.
+1975 1980 1985 1990 1995 2000 2005 2010 2015 2020 2023
+$\sum_{p\in\mathcal{P}}|D_{p}^{y}|$ 85 536 1092 1668 2271 2609 2897 3122 3447 3578 3612
+$\sum_{p\in\mathcal{P}}|G_{p}^{y}|$ 15 76 127 234 444 634 765 916 1088 1167 1178
+$\sum_{p\in\mathcal{P}}|\mathcal{D}_{p}^{y}\cap D^{y}|$ 397 1015 1559 2055 2394 2690 2909 3132 3458 3581 3614
+$\sum_{p\in\mathcal{P}}|\mathcal{G}_{p}^{y}\cap G^{y}|$ 25 136 271 553 1242 1896 2325 2570 2952 3174 3186
+$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in D_{p}^{y}}\{|[d]_{p}^{y}|\}\}$ 1.078 1.446 1.289 1.341 1.603 1.803 1.841 2.027 1.983 1.979 1.990
+$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in G_{p}^{y}}\{|[g]_{p}^{y}|\}\}$ 6.143 8.485 8.084 7.485 7.027 6.032 5.607 5.529 4.973 4.830 4.847
+$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in\mathcal{D}_{p}^{y}\cap D^{y}}\{|[d]^{y}|\}\}$ 1.078 1.430 1.356 1.616 2.211 2.532 2.724 2.902 2.749 2.759 2.776
+$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in\mathcal{G}_{p}^{y}\cap G^{y}}\{|[g]^{y}|\}\}$ 10.000 13.496 12.787 11.633 8.815 7.163 6.968 6.505 6.525 6.692 6.714
+Table S13: Statistics for settings P1Y1 and P2Y1.
+Year
+1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
+$\sum_{p\in\mathcal{P}}\left|D_{p}^{y\mid U_{y}}\right|$ 40 253 419 448 565 500 426 303 195 66
+$\sum_{p\in\mathcal{P}}\left|G_{p}^{y\mid U_{y}}\right|$ 10 58 90 154 272 295 250 218 163 61
+$\sum_{p\in\mathcal{P}}|\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}|$ 42 275 464 492 681 612 511 387 257 96
+$\sum_{p\in\mathcal{P}}|\mathcal{G}_{p}^{y}\cap G^{y\mid U_{y}}|$ 14 68 121 250 531 607 589 549 496 206
+$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{d\in D_{p}^{y\mid U_{y}}}\left\{\left|[d]_{p}^{y\mid U_{y}}\right|\right\}\right\}$ 1.034 1.445 1.247 1.285 1.477 1.571 1.596 1.724 1.648 1.614
+$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{g\in G_{p}^{y\mid U_{y}}}\left\{\left|[g]_{p}^{y\mid U_{y}}\right|\right\}\right\}$ 4.194 7.157 6.560 5.761 5.107 4.329 3.924 3.884 3.538 3.443
+$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}}\left\{\left|[d]^{y\mid U_{y}}\right|\right\}\right\}$ 1.034 1.427 1.301 1.497 1.925 2.126 2.288 2.376 2.176 2.113
+$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{g\in\mathcal{G}_{p}^{y}\cap G^{y\mid U_{y}}}\left\{\left|[g]^{y\mid U_{y}}\right|\right\}\right\}$ 7.000 10.525 9.027 7.890 5.667 4.729 4.392 4.155 4.096 4.249
+Table S14: Statistics for settings P1Y2 and P2Y2.
+1.6 Details of predictions using the GPT series
+In this study, we used GPT-3.5, GPT-4, and GPT-4o as baselines for generative models. The details of the models are shown in Table S5. The prompts are shown in Table S6.
 2 Details of experimental results
@@
 On the other hand, the sizes of the query sets in setting ${}^{{}^{\prime}}$ are more than twice as large as those in ${}^{{}^{\prime}}$, which likely contributes to the lower evaluation metric scores in setting ${}^{{}^{\prime}}$.
+2.1.3 Predictions by GPT models
+The prediction results from the GPT models are shown in Table S9.
+The set of query drugs and the set of correct target genes depend not only on the settings G, P1, and P2 but also on the vocabulary set.
+Here, we conducted experiments using vocabularies obtained from BioConceptVec or our skip-gram embeddings.
+The best accuracies for all combinations of vocabularies and settings are summarized in the Table 3 in the main text.
 2.2 Analogy tasks for drug-gene pairs by year
@@
 Year
-Metric Setting Method 1970 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
-Top1 P1Y1 Random 0.014 0.000 0.001 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000
-$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.653 0.515 0.484 0.394 0.303 0.275 0.263 0.256 0.285 0.269
-P1Y2 Random 0.000 0.001 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
-$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.780 0.674 0.656 0.651 0.633 0.636 0.627 0.630 0.659 0.623
-P2Y1 Random 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
-$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.625 0.467 0.437 0.358 0.250 0.226 0.221 0.209 0.236 0.184
-P2Y2 Random 0.005 0.003 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
-$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.744 0.633 0.620 0.627 0.613 0.616 0.619 0.621 0.653 0.617
-Top10 P1Y1 Random 0.029 0.020 0.009 0.003 0.003 0.001 0.002 0.001 0.000 0.000 0.000
-$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.800 0.670 0.727 0.638 0.597 0.573 0.627 0.566 0.665 0.612
-P1Y2 Random 0.026 0.012 0.005 0.004 0.002 0.002 0.001 0.001 0.001 0.000 0.000
-$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.906 0.842 0.863 0.863 0.879 0.896 0.913 0.899 0.925 0.911
-P2Y1 Random 0.000 0.013 0.006 0.002 0.002 0.001 0.001 0.000 0.000 0.000 0.000
-$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.762 0.627 0.674 0.590 0.533 0.502 0.567 0.513 0.563 0.480
-P2Y2 Random 0.032 0.013 0.007 0.004 0.002 0.002 0.001 0.001 0.000 0.000 0.001
-$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.872 0.814 0.832 0.845 0.866 0.883 0.909 0.896 0.922 0.912
-MRR P1Y1 Random 0.032 0.008 0.005 0.003 0.002 0.001 0.001 0.001 0.001 0.000 0.001
-$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.727 0.583 0.582 0.483 0.404 0.378 0.392 0.373 0.410 0.391
-P1Y2 Random 0.023 0.011 0.005 0.003 0.002 0.001 0.001 0.001 0.001 0.000 0.000
-$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.843 0.749 0.743 0.731 0.722 0.734 0.737 0.732 0.758 0.731
-P2Y1 Random 0.012 0.008 0.006 0.003 0.002 0.001 0.001 0.001 0.000 0.000 0.000
-$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.695 0.532 0.529 0.444 0.345 0.323 0.341 0.318 0.340 0.282
-P2Y2 Random 0.019 0.010 0.006 0.003 0.002 0.001 0.001 0.001 0.001 0.001 0.000
-$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.808 0.711 0.708 0.709 0.704 0.716 0.729 0.725 0.752 0.727
-Table S10: Gene prediction performance in settings P1Y1, P1Y2, P2Y1, and P2Y2.
+Metric Setting Method 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020 2023
+Top1 P1Y1 Random 0.001 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+$\hat{\mathbf{v}}_{p}^{y}$ 0.706 0.621 0.613 0.592 0.589 0.588 0.588 0.587 0.614 0.582 0.589
+P2Y1 Random 0.002 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+$\hat{\mathbf{v}}_{p}^{y}$ 0.638 0.562 0.568 0.570 0.581 0.581 0.596 0.597 0.626 0.591 0.600
+P1Y2 Random 0.003 0.001 0.001 0.000 0.000 0.001 0.000 0.000 0.000 0.000
+$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.350 0.482 0.442 0.366 0.290 0.266 0.254 0.244 0.272 0.258
+P2Y2 Random 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.333 0.444 0.399 0.333 0.244 0.221 0.213 0.202 0.226 0.177
+Top10 P1Y1 Random 0.014 0.009 0.004 0.002 0.001 0.001 0.001 0.001 0.000 0.001 0.000
+$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.871 0.862 0.835 0.847 0.881 0.889 0.867 0.886 0.869 0.862
+P2Y1 Random 0.015 0.006 0.004 0.003 0.002 0.002 0.001 0.001 0.001 0.001 0.001
+$\hat{\mathbf{v}}_{p}^{y}$ 0.915 0.797 0.812 0.811 0.842 0.877 0.899 0.881 0.900 0.884 0.880
+P1Y2 Random 0.018 0.006 0.004 0.002 0.002 0.002 0.001 0.001 0.000 0.002
+$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.625 0.648 0.704 0.621 0.586 0.564 0.620 0.558 0.662 0.606
+P2Y2 Random 0.012 0.007 0.003 0.002 0.001 0.001 0.001 0.001 0.001 0.001
+$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.595 0.615 0.655 0.575 0.530 0.495 0.562 0.506 0.560 0.479
+MRR P1Y1 Random 0.012 0.006 0.003 0.002 0.002 0.001 0.001 0.001 0.001 0.000 0.000
+$\hat{\mathbf{v}}_{p}^{y}$ 0.849 0.724 0.712 0.684 0.680 0.690 0.694 0.688 0.713 0.684 0.685
+P2Y1 Random 0.008 0.006 0.003 0.002 0.002 0.001 0.001 0.001 0.001 0.001 0.001
+$\hat{\mathbf{v}}_{p}^{y}$ 0.772 0.658 0.664 0.661 0.673 0.684 0.704 0.700 0.726 0.696 0.700
+P1Y2 Random 0.008 0.007 0.002 0.002 0.001 0.001 0.001 0.000 0.000 0.000
+$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.488 0.555 0.547 0.459 0.392 0.370 0.384 0.362 0.400 0.382
+P2Y2 Random 0.012 0.005 0.003 0.003 0.001 0.001 0.001 0.000 0.000 0.000
+$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.466 0.514 0.499 0.423 0.340 0.316 0.333 0.312 0.333 0.277
+Table S15: Gene prediction performance in settings P1Y1, P2Y1, P1Y2, and P2Y2.
 2.2.1 Global setting by year
-Table S7 shows the statistics for settings Y1 and Y2.
-As seen in Table S2, in setting Y1, the query is $d\in D^{y\mid U_{y}}$, and the search space is $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$.
-In setting Y2, the query is $d\in D^{y}$, and the search space is $\mathcal{G}^{y}$.
-Table S7 shows that $D^{y\mid U_{y}}$ has increased since 1970 and started to decrease after 1995.
-On the other hand, $D^{y}$ and $\mathcal{G}^{y}$ show a monotonic increase over the years.
+Tables S10 and S11 show the statistics for settings Y1 and Y2.
+As seen in Table S2, in setting Y1, the query is $d\in D^{y}$, and the search space is $\mathcal{G}^{y}$.
+In setting Y2, the query is $d\in D^{y\mid U_{y}}$, and the search space is $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$.
+In Table S10, $D^{y}$ and $\mathcal{G}^{y}$ show a monotonic increase over the years.
+On the other hand, Table S11 shows that $D^{y\mid U_{y}}$ has increased since 1975 and started to decrease after 1995.
 We showed the plots of the results for settings Y1 and Y2 in Fig. 3. For settings Y1 and Y2, Table S12 shows detailed results used in Fig. 3.
 2.2.2 Pahtway-wise setting by year
-Table S9 shows the statistics for settings P1Y1, P1Y2, P2Y1, and P2Y2.
-As seen in Table S2, the query is $d\in D_{p}^{y\mid U_{y}}$ in setting P1Y1, $d\in D_{p}^{y}$ in setting P1Y2, $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ in P2Y1, and $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ in setting P2Y2.
-Table S9 shows that $D_{p}^{y\mid U_{y}}$ and $\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ have increased since 1970 and started to decrease after 1995.
-On the other hand, $D_{p}^{y}$ and $\mathcal{D}_{p}^{y}\cap D^{y}$ show a monotonic increase over the years.
+Tables S13 and S14 show the statistics for settings P1Y1, P2Y1, P1Y2, and P2Y2.
+As seen in Table S2, the query is $d\in D_{p}^{y}$ in setting P1Y1, $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ in setting P2Y1, $d\in D_{p}^{y\mid U_{y}}$ in setting P1Y2, and $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ in P2Y2.
+In Table S13, $D_{p}^{y}$ and $\mathcal{D}_{p}^{y}\cap D^{y}$ show a monotonic increase over the years.
+On the other hand, Table S14 shows that $D_{p}^{y\mid U_{y}}$ and $\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ have increased since 1975 and started to decrease after 1995.
 We showed the plots of the results for settings P1Y1 and P1Y2 in Fig.
@@
 S2 shows the plots of the results for settings P2Y1 and P2Y2.
-For settings P1Y1, P1Y2, P2Y1, and P2Y2, Table S10 shows detailed results used in Fig.
+For settings P1Y1, P2Y1, P1Y2, and P2Y2, Table S15 shows detailed results used in Fig.
 3 and Fig.

### v4

@@
 To the best of our knowledge, there is limited research specifically focused on utilizing analogy computation of word embeddings to predict drug-gene relations in the biological domain.
-Our study aims to address this gap by investigating the application of analogy computation of word embeddings to drug-gene relationship prediction.
+Our study aims to address this gap by investigating the application of analogy computation of word embeddings to drug-gene relation prediction.
 In addition, neural networks have been used in a number of studies to predict these relations [21, 22, 23, 24, 25].
@@
 The mean embeddings are computed by Eq.
-(16).
+(S2) in Supplementary Information1.1.1.
 As a first step, based on the basic analogy task, we consider analogy tasks in the global setting, where all drugs and genes are used.
@@
 $\hat{\mathbf{v}}=\mathrm{E}_{\mathcal{R}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}}\{\mathbf{u}_{d}\},$ (8)
-$\mathrm{E}_{\mathcal{R}}\{\mathbf{u}_{g}\}=\frac{1}{|\mathcal{R}|}\sum_{(d,g)\in\mathcal{R}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{\mathcal{R}}}\{\mathbf{u}_{d}\}=\frac{1}{|\mathcal{R}|}\sum_{(d,g)\in\mathcal{R}}\mathbf{u}_{d}.$ (9)
+$\mathrm{E}_{\mathcal{\mathcal{R}}}\{\mathbf{u}_{d}\}=\frac{1}{|\mathcal{R}|}\sum_{(d,g)\in\mathcal{R}}\mathbf{u}_{d},\quad\mathrm{E}_{\mathcal{R}}\{\mathbf{u}_{g}\}=\frac{1}{|\mathcal{R}|}\sum_{(d,g)\in\mathcal{R}}\mathbf{u}_{g}.$ (9)
 To measure the performance of the estimator $\hat{\mathbf{v}}$ in Eq.
@@
 where $\mathrm{E}_{\mathcal{R}_{p}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}_{p}$.
-For easy comparison with Eq.
-(5), Eq.
-(14) is rewritten as the difference of mean vectors:
-$\hat{\mathbf{v}}_{p}=\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{d}\},$ (15)
-$\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{g}\}=\frac{1}{|\mathcal{R}_{p}|}\sum_{(d,g)\in\mathcal{R}_{p}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{d}\}=\frac{1}{|\mathcal{R}_{p}|}\sum_{(d,g)\in\mathcal{R}_{p}}\mathbf{u}_{d}.$ (16)
 To measure the performance of the estimator $\hat{\mathbf{v}}_{p}$, we prepare the evaluation of the analogy tasks.
@@
 (4).
-Therefore, similar to the estimator $\hat{\mathbf{v}}$ in Eq.
-(7), we define an estimator $\hat{\mathbf{v}}^{y}$ for the relation vector $\mathbf{v}^{y}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}^{y}$:
-$\hat{\mathbf{v}}^{y}:=\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}^{y}\right|}\sum_{(d,g)\in\mathcal{R}^{y}}(\mathbf{u}_{g}-\mathbf{u}_{d}),$ (20)
-where $\mathrm{E}_{\mathcal{R}^{y}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}^{y}$.
-For easy comparison with Eq.
-(5), Eq.
-(20) is rewritten as the difference of mean vectors:
-$\hat{\mathbf{v}}^{y}=\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{d}\},$ (21)
-$\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y}|}\sum_{(d,g)\in\mathcal{R}^{y}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y}|}\sum_{(d,g)\in\mathcal{R}^{y}}\mathbf{u}_{d}.$ (22)
+Therefore, we define the estimator $\hat{\mathbf{v}}^{y}$ for the relation vector $\mathbf{v}^{y}$ by using $\mathcal{R}^{y}$ instead of $\mathcal{R}$ in the estimator $\hat{\mathbf{v}}$ in Eq.
+(7).
 Given the above, similar to setting G, we perform the analogy tasks in the following setting.
@@
 In analogy tasks, we use “known” $\mathcal{R}^{y\mid L_{y}}$ and then predict the target genes $g$ from a drug $d$ for $(d,g)$ in “unknown” $\mathcal{R}^{y\mid U_{y}}$.
-Using the relation vector $\mathbf{v}^{y\mid L_{y}}$, which represents the drug-gene relations in $\mathcal{R}^{y\mid L_{y}}$, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}^{y}$ to $\mathbf{u}_{d}$:
+Using the relation vector $\mathbf{v}^{y\mid L_{y}}$, which represents the drug-gene relations in $\mathcal{R}^{y\mid L_{y}}$, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}^{y\mid L_{y}}$ to $\mathbf{u}_{d}$:
 $\mathbf{u}_{d}+\mathbf{v}^{y\mid L_{y}}\approx\mathbf{u}_{g}.$ (19)
@@
 (17).
-We define the estimator $\hat{\mathbf{v}}^{y\mid L_{y}}$ for the relation vector $\mathbf{v}^{y\mid L_{y}}$ by using $\mathcal{R}^{y\mid L_{y}}$ instead of $\mathcal{R}^{y}$ in the estimator $\hat{\mathbf{v}}^{y}$ in Eq.
+Therefore, we define the estimator $\hat{\mathbf{v}}^{y\mid L_{y}}$ for the relation vector $\mathbf{v}^{y\mid L_{y}}$ by using $\mathcal{R}^{y\mid L_{y}}$ instead of $\mathcal{R}$ in the estimator $\hat{\mathbf{v}}$ in Eq.
 (7).
@@
 To complement the pretrained BioConceptVec embeddings, we further trained 300-dimensional skip-gram embeddings on the publicly available PubMed abstracts.
-To train our embeddings, we used PubTator [14] to normalize six major biological concepts (genes, mutations, diseases, chemicals, cell lines, and species) in PubMed abstracts, and then tokenized them using NLTK [35].
+As with the original BioConceptVec, to train our embeddings, we used PubTator [14] to convert six major biological concepts (genes, mutations, diseases, chemicals, cell lines, and species) in PubMed abstracts into their respective IDs, followed by tokenization using NLTK [35].
 Note that since widely used embeddings such as word2vec and GloVe typically have 300 dimensions, we set the dimensions of our embeddings to 300 instead of the original 100.
@@
 For BioConceptVec and our skip-gram embeddings, Table 2 shows some basic statistics for setting G.
-Due to the difference in training data size and minimal word occurrence, the sizes of certain sets such as $|\mathcal{D}|$ and $|\mathcal{G}|$ differ significantly between BioConceptVec and our skip-gram embeddings, but the sizes of other sets show somewhat similar trends.
+Due to the difference in training data size and minimal word occurrence, the sizes of certain sets such as $|\mathcal{D}|$ and $|\mathcal{G}|$ differ significantly between BioConceptVec and our skip-gram embeddings, but the sizes of other sets, such as $|\mathcal{R}|$, which represents the size of overall relations, show somewhat similar trends.
+Figure S2 in Supplementary Information 1.4 shows the distribution of the sizes of the answer sets for each drug $d$ under settings G, P1, and P2 for both BioConceptVec and our skip-gram embeddings.
 Table S1 in Supplementary Information 1.1.1 shows the statistics for settings P1 and P2.
@@
 To evaluate the performance of predicting target genes by adding the relation vector to a drug embedding, we compare it to baseline methods.
-As a simple baseline, we randomly select one gene from the search space for each query.
+As a simple yet reasonable baseline, we randomly sampled genes from the set $G$ containing the genes that have drug-gene relations with at least one drug.
+To increase the probability of sampling genes with more drug-gene relations, a gene $g\in G$ was sampled with a probability proportional to $|[g]|$, the size of the set of drugs that have correct drug-gene relations with $g$.
+In other words, sampling was performed with probabilities proportional to the number of queries that have correct relations.
+For simplicity, this sampling method was consistently used across settings G, P1, and P2.
 We refer to this baseline as the random baseline.
-In the random baseline, we repeated the experiments 10 times and used the average score as the final result.
+In this baseline, we repeated the experiments 10 times and used the average score as the final result.
+Similar random baselines can also be applied to settings by year and to settings for predicting drugs from genes.
+See Supplementary Information 1.6 for details.
+The analogy task we consider can be regarded as a task of predicting relationships between entities.
+Therefore, we adopted Knowledge Graph Embedding (KGE) [37] as a general baseline method for predicting drug-gene relations.
+KGE learns embeddings for head entities, tail entities, and relations from triplets by leveraging the explicit structure of the knowledge graph.
+Using the learned embeddings, it predicts a specific tail corresponding to a given head and relation.
+Among various KGE methods, we used TransE [38], one of the most representative approaches, with the embedding dimensions set to the commonly used size of 500.
+Similar to the analogy computation, TransE predicts the tail embedding by adding the relation vector to the head embedding.
+However, while the analogy computation relies solely on document data and pathway information, TransE leverages explicit relations in the knowledge graph to directly learn embeddings and relation vectors.
+Word embeddings such as skip-gram are learned from large corpora (e.g., PubMed abstracts) without being tied to any particular research field or specific target task.
+Consequently, they capture broad linguistic and semantic information (not restricted to drug-gene relations), enabling them to be applied to various tasks (e.g., analogy tasks [39], sentence similarity tasks [5]).
+In contrast, KGE is learned based on knowledge and relational information tailored to a specific task, resulting in embeddings specialized for that task.
+For details on the dataset splits and hyperparameters used for training TransE, see Supplementary Information 1.8.
 In addition to this, as strong generative model baselines, we used the Generative Pre-trained Transformer (GPT) [40] series to predict the top 10 target genes from a query drug by zero-shot.
 We used the GPT-3.5 [41], GPT-4 [42], and GPT-4o (https://platform.openai.com/docs/models/gpt-4o) models with the temperature hyperparameter set to 00.
-See Supplementary Information 1.6 for more details on the models and prompt template for the predictions.
+See Supplementary Information 1.7 for more details on the models and prompt template for the predictions.
 Evaluation metrics
@@
 MRR is a statistical measure to evaluate search performance, expressed as the sum of the reciprocals of the ranks at which the predicted results first appear in the answer set.
-Note that since MRR cannot be calculated for the GPT models due to the limited number of prediction candidates output by GPT, only top-1 and top-10 accuracies are calculated.
+Note that, in the random baseline, all genes $g\in G$ are ranked without duplication through sampling with probabilities proportional to $|[g]|$.
+By treating these rankings as prediction results, we can calculate not only top-1 and top-10 accuracies but also MRR.
+Since MRR cannot be calculated for the GPT models due to the limited number of prediction candidates output by GPT, only top-1 and top-10 accuracies are calculated.
 When evaluating performance using top-1 accuracy, top-10 accuracy, and MRR, the embeddings of all genes $\mathcal{G}$ (or $\mathcal{G}^{y}$ in settings by year) are centered using the mean embedding $\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\}$ in Eq.
@@
 Model Setting Method Top1 Top10 MRR
-BioConceptVec G Random 0.000 0.000 0.000
-$\hat{\mathbf{v}}$ 0.304 0.645 0.419
-P1 Random 0.000 0.000 0.000
-$\hat{\mathbf{v}}_{p}$ 0.499 0.790 0.602
-P2 Random 0.000 0.000 0.000
-$\hat{\mathbf{v}}_{p}$ 0.522 0.807 0.624
-Our embeddings G Random 0.000 0.000 0.001
-$\hat{\mathbf{v}}$ 0.300 0.686 0.426
-P1 Random 0.000 0.001 0.001
-$\hat{\mathbf{v}}_{p}$ 0.589 0.862 0.685
-P2 Random 0.000 0.001 0.001
-$\hat{\mathbf{v}}_{p}$ 0.600 0.880 0.700
+BioConceptVec G Random 0.019 0.137 0.063
+TransE 0.576 0.682 0.615
+analogy 0.304 0.645 0.419
+P1 Random 0.018 0.140 0.063
+TransE 0.745 0.871 0.792
+analogy 0.499 0.790 0.602
+P2 Random 0.022 0.162 0.072
+TransE 0.813 0.896 0.842
+analogy 0.522 0.807 0.624
+Our embeddings G Random 0.020 0.140 0.065
+TransE 0.571 0.689 0.614
+analogy 0.300 0.686 0.426
+P1 Random 0.020 0.142 0.066
+TransE 0.752 0.873 0.798
+analogy 0.589 0.862 0.685
+P2 Random 0.023 0.164 0.074
+TransE 0.819 0.907 0.854
+analogy 0.600 0.880 0.700
 GPT-3.5 Best of all zero-shot 0.653 0.821
@@
 Table 3: Gene prediction performance in settings G, P1, and P2.
+In analogy tasks, the relation vector for prediction is $\hat{\mathbf{v}}$ in setting G and $\hat{\mathbf{v}}_{p}$ in settings P1 and P2.
+In the random baseline, a gene $g\in G$ is sampled with probabilities proportional to the size of the set of drugs related to $g$.
+TransE is evaluated using a different data split, while for the GPT series, only the best scores across settings are shown in the table.
+Since TransE’s performance depends on the size of the training data, Fig.
+4 compares analogy computation with TransE under varying proportions of the training data.
 (a) Y1
@@
 Since the analogy tasks are considered for each pathway, these settings are also likely to make the tasks easier than those in setting G.
-The random baseline in settings G, P1, and P2 showed nearly zero scores for all evaluation metrics, suggesting that random prediction is nearly impossible in these analogy tasks.
+The performance of the random baseline in settings G, P1, and P2 was consistently low across all evaluation metrics.
+Although random sampling was restricted to genes in the set $G$, these results demonstrate that random prediction is challenging in these analogy tasks.
 This also confirms the superior performance of prediction by adding relation vectors.
+TransE, as a KGE method, naturally outperforms our proposed approach across all three evaluation metrics.
+Interestingly, however, the top-10 accuracy of the prediction by vector addition is comparable to that of TransE.
+This suggests that our method has the latent potential to achieve performance close to TransE without directly exploiting explicit relational information.
+We then hypothesize that the strong performance of TransE observed in Table 3 depends on the size of the training data.
+In the Discussion section, we compare analogy computation and TransE under varying proportions of the training data (Fig.
+4).
 Among the GPT series, GPT-4o performed the best, followed by GPT-4, which outperformed GPT-3.5.
@@
 In setting Y2, the datasets for later years have fewer queries and larger search spaces, which tends to produce lower evaluation metric scores.
-In both settings, the evaluation metric scores for the random baseline are close to zero.
+In both settings, the performance of the random baseline is extremely high for older years such as $y=1975$ and $y=1980$, where the number of genes related to drugs is small.
+However, after $y=1985$, our method consistently outperforms the random baseline.
 For more detailed results for settings Y1 and Y2, see Table S12 in Supplementary Information 2.2.1.
@@
 In addition, the results for the older years show high evaluation metric scores in settings P1Y1 and P1Y2, similar to those in settings Y1 and Y2.
-The evaluation metric scores of the random baseline are close to zero.
+The performance of the random baseline also shows similar trends to those observed in settings Y1 and Y2.
 For the results in settings P2Y1 and P2Y2, see Fig.
-S2 in Supplementary Information 2.2.2.
+S4 in Supplementary Information 2.2.2.
 They are very similar to those in settings P1Y1 and P1Y2.
@@
 Discussion
+Connections to trends in existing studies
 Mikolov et al.
@@
 [18] reported an overall accuracy of 60.1% when using analogy computation of word embedding to predict 29,046 relationships across various concepts in materials science, including element names, crystal symmetries, and magnetic properties.
-Our results for drug-gene relationship prediction show a similar level of accuracy, suggesting that the approach of using analogy computation of word embedding can be effective in the biomedical domain as well.
+Our results for drug-gene relation prediction show a similar level of accuracy, suggesting that the approach of using analogy computation of word embedding can be effective in the biomedical domain as well.
 This demonstrates the potential of this method across different scientific disciplines.
@@
 This variability suggests that the effectiveness of word embedding analogies may depend on the specific type of relationship being predicted.
-In Table 3, predictions using $\mathbf{v}_{p}$ with our skip-gram embeddings in settings P1 and P2 performed on par with GPT-4o and outperformed both GPT-3.5 and GPT-4.
+Strengths and insights from prediction by adding the relation vector
+Figure 4: Changes in evaluation metrics on the test data for analogy computation and TransE under setting P1 as the proportion of training data varies.
+Analogy computation maintains consistent performance regardless of the proportion of training data, and outperforms TransE when the proportion of training data is small.
+See Supplementary Information 2.4 for experimental details.
+To test the hypothesis that the superior performance of TransE depends on the size of the training data, we compared the performance of our method with that of TransE in setting P1 by varying the proportion of training data used to compute relation vectors (see Supplementary Information 2.4 for experimental details).
+As Fig.
+4 shows, TransE’s performance degrades as the proportion of training data decreases, whereas our method maintains nearly constant performance.
+In particular, our method outperforms TransE when the available training data is small.
+This is because our approach derives relation vectors from already-trained word embeddings, while TransE learns all embeddings from scratch using the training data.
+These findings suggest that our method can function efficiently even with limited datasets.
+In addition, in Table 3, predictions using $\mathbf{v}_{p}$ with our skip-gram embeddings in settings P1 and P2 performed on par with GPT-4o and outperformed both GPT-3.5 and GPT-4.
 It is rather surprising that simple vector addition can achieve performance comparable to large language models.
 Given that the GPT series is trained on vast amounts of text data, using relation vector addition is an efficient method.
-Furthermore, the results showing that target genes can be predicted with performance comparable to or better than large language models imply that high-quality information about drug-gene relationships is encoded in the embeddings.
+Furthermore, the results showing that target genes can be predicted with performance comparable to or better than large language models imply that high-quality information about drug-gene relations is encoded in the embeddings.
 This also suggests that the computation of analogies may be a fundamental inference principle within large language models.
-In addition, we divided the datasets by year and redefined the vectors representing known relations to predict unknown relations.
+Furthermore, we divided the datasets by year and redefined the vectors representing known relations to predict unknown relations.
 The experimental results presented in Fig.
@@
 The experiments conducted with year-split datasets are a novel attempt not found in existing research, and confirmed the effectiveness of predicting unknown future events.
+Biological information intrinsic to the embedding space
 In Table 4, predicted genes in the ErbB signaling pathway are closely related to the genes or symptoms that have direct molecular interactions with the target genes, or alternative therapeutic targets for the target disease.
 This may suggest that the high-dimensional representations of genes and drugs calculated from biomedical texts not only capture the simple drug-gene relations we aimed to predict in this study but also integrate broader, higher-order information.
+Limitations
+This study has several limitations.
+First, we focus exclusively on drug-gene relations. Consequently, the applicability of this method to other biological relationships remains untested.
+Additionally, as shown in Table 3, our approach for predicting genes from drugs by simply adding relation vectors performs worse than KGE models such as TransE in terms of accuracy.
+While the goal of this study is not to improve performance but to explore the information intrinsic to the embedding space, efforts to improve performance remain an important future direction.
+Furthermore, the embeddings used in this study are static embeddings learned with skip-gram, which cannot handle out-of-vocabulary concepts.
+Assigning a single embedding to a concept that serves different roles depending on the context (e.g., concepts used differently in the past and present) may also be suboptimal.
+Future Directions
+Building on the limitations of this study, several future directions can be considered.
+First, since BioConceptVec includes concepts such as diseases and mutations, it would be useful to investigate whether analogy computations using relation vectors can also be applied to relationships beyond drugs and genes.
+This could serve as a valuable follow-up study.
+Additionally, using dynamic embeddings instead of static embeddings represents another promising direction.
+Since models like BERT [4] compute embeddings based on context, taking advantage of this feature may allow more accurate modeling of drug-gene relations.
+For example, this approach could help capture relations that vary with context, as well as drugs whose applications evolve over time.
+Furthermore, it would be interesting to investigate how the properties of drugs or genes affect performance in analogy tasks.
+For example, if the drug-gene relations of a drug have been extensively studied, can related genes for that drug be more easily predicted?
+To address this question, we conducted a simple experiment to examine the correlation between the size of the answer set for each drug and the rank of the search results predicted by adding relation vectors.
+Scatter plots of these values for settings G, P1, and P2, using BioConceptVec and our embeddings, are shown in Fig.
+S5 in Supplementary Information 2.3.
+Mathematically, a larger answer set is expected to make prediction easier, resulting in search result ranks closer to 1.
+The results showed a slight negative correlation in setting G, as expected (BioConceptVec: $-0.201$, our embeddings: $-0.196$).
+However, in settings P1 and P2, little to no correlation was observed (BioConceptVec: $0.054$ and $-0.102$, our embeddings: $0.079$ and $-0.041$).
+This suggests that in settings P1 and P2, incorporating pathway information into relation vectors helps maintain good search results even when the answer sets are small.
+By performing such detailed analyses, we can further deepen our understanding of the biological knowledge intrinsic to the embedding space.
 Conclusions
@@
 Acknowledgements
-This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3, JST SPRING JPMJSP2110.
+This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3.
 Author contributions statement
@@
 and H.S.
-conceived the experiments, H.Y., R.H., K.A., and K.M.
-conducted the experiments, H.Y., R.H., K.A., K.M., S.S., Ma.O., and H.S.
+conceived the experiments, H.Y., R.H., K.A., Y.Z.
+and K.M.
+conducted the experiments, H.Y., R.H., K.A., K.M., S.S., Y.Z., Ma.O., and H.S.
 analysed the results, R.H.
@@
 1.1.1 Pathway-wise setting
+For easy comparison with Eq. (5), Eq. (14) is rewritten as the difference of mean vectors:
+$\hat{\mathbf{v}}_{p}=\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{d}\},$ (S1)
+$\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{d}\}=\frac{1}{|\mathcal{R}_{p}|}\sum_{(d,g)\in\mathcal{R}_{p}}\mathbf{u}_{d},\quad\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{g}\}=\frac{1}{|\mathcal{R}_{p}|}\sum_{(d,g)\in\mathcal{R}_{p}}\mathbf{u}_{g}.$ (S2)
 In Eq.
@@
 1.2.1 Global setting by year
+Similar to the estimator $\hat{\mathbf{v}}$ in Eq.
+(7), we define an estimator $\hat{\mathbf{v}}^{y}$ for the relation vector $\mathbf{v}^{y}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}^{y}$:
+$\hat{\mathbf{v}}^{y}:=\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}^{y}\right|}\sum_{(d,g)\in\mathcal{R}^{y}}(\mathbf{u}_{g}-\mathbf{u}_{d}),$ (S5)
+where $\mathrm{E}_{\mathcal{R}^{y}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}^{y}$.
+For easy comparison with Eq.
+(5), Eq.
+(S5) is rewritten as the difference of mean vectors:
+$\hat{\mathbf{v}}^{y}=\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{d}\},$ (S6)
+$\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y}|}\sum_{(d,g)\in\mathcal{R}^{y}}\mathbf{u}_{d},\quad\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y}|}\sum_{(d,g)\in\mathcal{R}^{y}}\mathbf{u}_{g}.$ (S7)
 Similar to $\hat{\mathbf{v}}$ in Eq.
 (7), to measure the performance of the estimator $\hat{\mathbf{v}}^{y}$ in Eq.
-(20), we prepare the evaluation of the analogy tasks.
+(S5), we prepare the evaluation of the analogy tasks.
 Using the projection operations in Eq.
@@
 (S9) by using $\mathcal{R}^{y\mid L_{y}}$ instead of $\mathcal{R}^{y}$ in the estimator $\hat{\mathbf{v}}^{y}$ in Eq.
-(20).
+(S5).
 For easy comparison with Eq.
@@
 $\hat{\mathbf{v}}^{y\mid L_{y}}=\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\},$ (S10)
-$\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{d}.$ (S6)
+$\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{d},\quad\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{g}.$ (S11)
 Similar to $\hat{\mathbf{v}}$ in Eq.
@@
 $\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in G_{p}}\{|[g]_{p}|\}\}$ 5.050 4.847
-$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in\mathcal{D}_{p}\cap D}\{|[d]|\}\}$ 16.346 16.302
-$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in\mathcal{G}_{p}\cap G}\{|[g]|\}\}$ 101.765 91.488
+$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in\mathcal{D}_{p}\cap D}\{|[d]|\}\}$ 2.738 2.776
+$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in\mathcal{G}_{p}\cap G}\{|[g]|\}\}$ 7.131 6.714
 Table S1: Statistics for settings P1 and P2.
@@
 $\mathbf{u}_{d}+\mathbf{v}_{p}^{y}\approx\mathbf{u}_{g}.$ (S14)
-Similar to equations (13) and (19), Eq. (S9) corresponds to Eq. (4). Therefore, an estimator $\hat{\mathbf{v}}_{p}^{y}$ for the relation vector $\mathbf{v}_{p}^{y}$ in Eq. (S9) is defined as follows:
+Similar to equations (13) and (17), Eq.
+(S14) corresponds to Eq.
+(4).
+Therefore, an estimator $\hat{\mathbf{v}}_{p}^{y}$ for the relation vector $\mathbf{v}_{p}^{y}$ in Eq.
+(S14) is defined as follows:
 $\hat{\mathbf{v}}_{p}^{y}:=\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}_{p}^{y}\right|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}(\mathbf{u}_{g}-\mathbf{u}_{d}).$ (S15)
@@
 $\hat{\mathbf{v}}_{p}^{y}=\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{d}\},$ (S16)
-$\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}\mathbf{u}_{d}.$ (S12)
+$\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}\mathbf{u}_{d},\quad\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}\mathbf{u}_{g}.$ (S17)
 Similar to $\hat{\mathbf{v}}$ in Eq.
@@
 $\hat{\mathbf{v}}_{p}^{y\mid L_{y}}=\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{d}\},$ (S22)
-$\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}\mathbf{u}_{d}.$ (S18)
+$\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}\mathbf{u}_{d},\quad\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}\mathbf{u}_{g}.$ (S23)
 Similar to $\hat{\mathbf{v}}$ in Eq.
@@
 Table S4: Hyperparameter for our skip-gram.
-Model Version
-GPT-3.5 gpt-3.5-turbo-0125
-GPT-4 gpt-4-turbo-2024-04-09
-GPT-4o gpt-4o-2024-05-13
-Table S5: Version of each GPT model.
-Role Prompt
-system You are a biologist specializing in drug-target interactions. Your responses should be formatted as JSON, with data listed under the key ‘target_genes’.
-user List the top $10$ potential target genes for the drug drug name.
-Table S6: Prompt template used for the GPT series to predict the top 10 target genes for each query drug. Here, drug name is a specific name such as Bosutinib.
 Table S2 shows the settings for predicting genes from drugs, as used in the main text.
@@
 1.4 Embeddings
+Figure S2 shows the distribution of the sizes of the answer sets for each drug $d$ in settings G, P1, and P2 for BioConceptVec and our skip-gram embeddings.
+Similarly, Fig.
+S3 shows the distribution of the sizes of the answer sets for each gene $g$ in settings ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, and ${}^{{}^{\prime}}$ for BioConceptVec and our skip-gram embeddings.
 The hyperparameters used to train our skip-gram are shown in Table S4.
@@
 We also applied this conversion to drugs in the vocabulary of our trained skip-gram model.
+(a) BioConceptVec
+(b) Our embeddings
+Figure S2: Distribution of the sizes of the answer sets for each drug $d$ in the settings for predicting genes from drugs.
+The mean values are also shown: for setting G, the value of $\mathrm{E}_{d\in D}\{|[d]|\}$ from Table 2; for setting P1, the value of $\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in D_{p}}\{|[d]_{p}|\}\}$ from Table S1; and for setting P2, the value of $\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in\mathcal{D}_{p}\cap D}\{|[d]|\}\}$ from Table S1.
+(a) BioConceptVec
+(b) Our embeddings
+Figure S3: Distribution of the sizes of the answer sets for each gene $g$ in the settings for predicting drugs from genes.
+The mean values are also shown: for setting ${}^{{}^{\prime}}$, the value of $\mathrm{E}_{g\in G}\{|[g]|\}$ from Table 2; for setting ${}^{{}^{\prime}}$, the value of $\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in G_{p}}\{|[g]_{p}|\}\}$ from Table S1; and for setting ${}^{{}^{\prime}}$, the value of $\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in\mathcal{G}_{p}\cap G}\{|[g]|\}\}$ from Table S1.
+Model Version
+GPT-3.5 gpt-3.5-turbo-0125
+GPT-4 gpt-4-turbo-2024-04-09
+GPT-4o gpt-4o-2024-05-13
+Table S5: Version of each GPT model.
+Role Prompt
+system You are a biologist specializing in drug-target interactions. Your responses should be formatted as JSON, with data listed under the key ‘target_genes’.
+user List the top $10$ potential target genes for the drug drug name.
+Table S6: Prompt template used for the GPT series to predict the top 10 target genes for each query drug. Here, drug name is a specific name such as Bosutinib.
 Model Setting Method Centering Top1 Top10 MRR
-Our embeddings G Random 0.000 0.000 0.001
+Our embeddings G Random 0.020 0.140 0.065
 $\hat{\mathbf{v}}_{\text{naive}}$ 0.067 0.197 0.111
@@
 $\checkmark$ 0.300 0.686 0.426
-P1 Random 0.000 0.001 0.001
+P1 Random 0.020 0.142 0.066
 $\hat{\mathbf{v}}_{p,\text{naive}}$ 0.364 0.661 0.470
@@
 $\checkmark$ 0.589 0.862 0.685
-P2 Random 0.000 0.001 0.001
+P2 Random 0.023 0.164 0.074
 $\hat{\mathbf{v}}_{p,\text{naive}}$ 0.376 0.682 0.484
@@
 Model Setting Method Top1 Top10 MRR
-BioConceptVec ${}^{{}^{\prime}}$ Random 0.000 0.001 0.001
+BioConceptVec ${}^{{}^{\prime}}$ Random 0.011 0.075 0.036
 $\hat{\mathbf{v}}^{\prime}$ 0.233 0.560 0.345
-${}^{{}^{\prime}}$ Random 0.000 0.001 0.001
+${}^{{}^{\prime}}$ Random 0.010 0.065 0.032
 $\hat{\mathbf{v}}_{p}^{\prime}$ 0.426 0.679 0.515
-${}^{{}^{\prime}}$ Random 0.000 0.001 0.001
+${}^{{}^{\prime}}$ Random 0.009 0.066 0.031
 $\hat{\mathbf{v}}_{p}^{\prime}$ 0.248 0.507 0.337
-Our embeddings ${}^{{}^{\prime}}$ Random 0.000 0.003 0.002
+Our embeddings ${}^{{}^{\prime}}$ Random 0.010 0.078 0.037
 $\hat{\mathbf{v}}^{\prime}$ 0.240 0.577 0.351
-${}^{{}^{\prime}}$ Random 0.000 0.003 0.002
+${}^{{}^{\prime}}$ Random 0.012 0.070 0.034
 $\hat{\mathbf{v}}_{p}^{\prime}$ 0.478 0.751 0.571
-${}^{{}^{\prime}}$ Random 0.000 0.003 0.002
+${}^{{}^{\prime}}$ Random 0.009 0.070 0.033
 $\hat{\mathbf{v}}_{p}^{\prime}$ 0.279 0.516 0.359
@@
 Metric Setting Method 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020 2023
-Top1 Y1 Random 0.001 0.000 0.000 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+Top1 Y1 Random 0.143 0.066 0.054 0.036 0.027 0.022 0.024 0.022 0.022 0.019 0.019
 $\hat{\mathbf{v}}^{y}$ 0.365 0.262 0.249 0.236 0.247 0.272 0.275 0.287 0.323 0.301 0.300
-Y2 Random 0.000 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+Y2 Random 0.104 0.081 0.043 0.028 0.016 0.012 0.013 0.006 0.009 0.004
 $\hat{\mathbf{v}}^{y\mid L_{y}}$ 0.333 0.208 0.165 0.132 0.076 0.092 0.062 0.090 0.107 0.021
-Top10 Y1 Random 0.016 0.007 0.003 0.002 0.002 0.001 0.001 0.001 0.001 0.001 0.000
+Top10 Y1 Random 0.830 0.554 0.426 0.317 0.236 0.196 0.185 0.165 0.153 0.137 0.140
 $\hat{\mathbf{v}}^{y}$ 0.837 0.577 0.563 0.562 0.578 0.652 0.691 0.651 0.714 0.693 0.686
-Y2 Random 0.015 0.003 0.002 0.003 0.001 0.001 0.000 0.001 0.000 0.000
+Y2 Random 0.835 0.538 0.382 0.253 0.157 0.123 0.098 0.064 0.062 0.081
 $\hat{\mathbf{v}}^{y\mid L_{y}}$ 0.625 0.487 0.438 0.410 0.347 0.397 0.352 0.344 0.328 0.271
-MRR Y1 Random 0.011 0.005 0.003 0.002 0.001 0.001 0.001 0.001 0.001 0.001 0.000
+MRR Y1 Random 0.340 0.204 0.167 0.127 0.100 0.084 0.081 0.073 0.069 0.064 0.064
 $\hat{\mathbf{v}}^{y}$ 0.534 0.370 0.353 0.346 0.363 0.399 0.410 0.410 0.455 0.430 0.426
-Y2 Random 0.008 0.006 0.003 0.002 0.001 0.001 0.000 0.000 0.000 0.000
+Y2 Random 0.305 0.211 0.150 0.104 0.070 0.055 0.049 0.034 0.034 0.032
 $\hat{\mathbf{v}}^{y\mid L_{y}}$ 0.438 0.304 0.253 0.214 0.163 0.182 0.147 0.173 0.188 0.103
@@
 Table S14: Statistics for settings P1Y2 and P2Y2.
+1.6 Details of predictions using random baseline
+This section describes the sets used for sampling in the random baseline across different settings.
+For settings G, P1, and P2, the set used was $G$.
+Similarly, $D$ was used for ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, and ${}^{{}^{\prime}}$, while $G^{y}$ was used for settings Y1, P1Y1, and P2Y1.
+For settings Y2, P1Y2, and P2Y2, $G^{y\mid L_{y}}\cup G^{y\mid U_{y}}$ was used instead of $G^{y\mid U_{y}}$; This was intended to prevent extremely high random baseline performance when the size of $G^{y\mid U_{y}}$ is very small.
 1.7 Details of predictions using the GPT series
 In this study, we used GPT-3.5, GPT-4, and GPT-4o as baselines for generative models. The details of the models are shown in Table S5. The prompts are shown in Table S6.
+1.8 Details of predictions by TransE
+This section describes the details of the experimental settings for TransE.
+Embeddings are learned for each of settings G, P1, and P2.
+In our experiments, we utilized TransE with the following hyperparameters: a batch size of 1024, embedding dimensions set to 500, a learning rate of 0.0001, and a total of 800 iteration steps.
+Additionally, we adopted the self-adversarial negative sampling method from RotatE [60], configuring the negative sampling size to 512 and the sampling temperature to 0.5.
+For each setting, the triplets of drug $d$, gene $g$, and drug-gene relation $(d,g)$ are used to evaluate performance in analogy tasks and are then split into 60% training, 20% validation, and 20% test data.
+In setting G, only a single type of relation, the drug-gene relation, is considered.
+In settings P1 and P2, the drug-gene relation $(d,g)$ is distinguished based on each pathway $p\in\mathcal{P}$, resulting in $|\mathcal{P}|$ different relations.
+For entities in the validation and test data that do not appear in the training data, random embeddings are simply used.
+Since KGE typically uses the entire set of head and tail entities as its search space, we similarly define the search space for TransE as the entire set of drugs and genes that have drug-gene relations for each setting.
+Additionally, note that in Table 3, while random baseline, GPT series, and our method evaluate performance on the entire dataset, TransE evaluates performance on the 20% test data.
 2 Details of experimental results
@@
 The prediction results from the GPT models are shown in Table S9.
-The set of query drugs and the set of correct target genes depend not only on the settings G, P1, and P2 but also on the vocabulary set.
+The set of query drugs and the set of correct target genes depend not only on settings G, P1, and P2 but also on the vocabulary set.
 Here, we conducted experiments using vocabularies obtained from BioConceptVec or our skip-gram embeddings.
@@
 Metric Setting Method 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020 2023
-Top1 P1Y1 Random 0.001 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+Top1 P1Y1 Random 0.151 0.086 0.062 0.044 0.028 0.025 0.022 0.021 0.021 0.019 0.020
 $\hat{\mathbf{v}}_{p}^{y}$ 0.706 0.621 0.613 0.592 0.589 0.588 0.588 0.587 0.614 0.582 0.589
-P2Y1 Random 0.002 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+P2Y1 Random 0.178 0.083 0.060 0.043 0.034 0.030 0.027 0.029 0.023 0.021 0.022
 $\hat{\mathbf{v}}_{p}^{y}$ 0.638 0.562 0.568 0.570 0.581 0.581 0.596 0.597 0.626 0.591 0.600
-P1Y2 Random 0.003 0.001 0.001 0.000 0.000 0.001 0.000 0.000 0.000 0.000
+P1Y2 Random 0.147 0.079 0.054 0.033 0.017 0.012 0.013 0.007 0.004 0.006
 $\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.350 0.482 0.442 0.366 0.290 0.266 0.254 0.244 0.272 0.258
-P2Y2 Random 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+P2Y2 Random 0.107 0.074 0.051 0.032 0.019 0.014 0.009 0.007 0.005 0.003
 $\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.333 0.444 0.399 0.333 0.244 0.221 0.213 0.202 0.226 0.177
-Top10 P1Y1 Random 0.014 0.009 0.004 0.002 0.001 0.001 0.001 0.001 0.000 0.001 0.000
+Top10 P1Y1 Random 0.906 0.602 0.486 0.351 0.245 0.214 0.186 0.165 0.146 0.140 0.142
 $\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.871 0.862 0.835 0.847 0.881 0.889 0.867 0.886 0.869 0.862
-P2Y1 Random 0.015 0.006 0.004 0.003 0.002 0.002 0.001 0.001 0.001 0.001 0.001
+P2Y1 Random 0.862 0.601 0.492 0.366 0.278 0.243 0.214 0.199 0.171 0.162 0.170
 $\hat{\mathbf{v}}_{p}^{y}$ 0.915 0.797 0.812 0.811 0.842 0.877 0.899 0.881 0.900 0.884 0.880
-P1Y2 Random 0.018 0.006 0.004 0.002 0.002 0.002 0.001 0.001 0.000 0.002
+P1Y2 Random 0.922 0.583 0.448 0.269 0.156 0.118 0.098 0.068 0.061 0.076
 $\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.625 0.648 0.704 0.621 0.586 0.564 0.620 0.558 0.662 0.606
-P2Y2 Random 0.012 0.007 0.003 0.002 0.001 0.001 0.001 0.001 0.001 0.001
+P2Y2 Random 0.898 0.558 0.450 0.274 0.173 0.129 0.104 0.076 0.064 0.057
 $\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.595 0.615 0.655 0.575 0.530 0.495 0.562 0.506 0.560 0.479
-MRR P1Y1 Random 0.012 0.006 0.003 0.002 0.002 0.001 0.001 0.001 0.001 0.000 0.000
+MRR P1Y1 Random 0.366 0.231 0.186 0.140 0.102 0.090 0.080 0.073 0.068 0.064 0.065
 $\hat{\mathbf{v}}_{p}^{y}$ 0.849 0.724 0.712 0.684 0.680 0.690 0.694 0.688 0.713 0.684 0.685
-P2Y1 Random 0.008 0.006 0.003 0.002 0.002 0.001 0.001 0.001 0.001 0.001 0.001
+P2Y1 Random 0.369 0.228 0.186 0.144 0.115 0.101 0.091 0.088 0.076 0.072 0.074
 $\hat{\mathbf{v}}_{p}^{y}$ 0.772 0.658 0.664 0.661 0.673 0.684 0.704 0.700 0.726 0.696 0.700
-P1Y2 Random 0.008 0.007 0.002 0.002 0.001 0.001 0.001 0.000 0.000 0.000
+P1Y2 Random 0.358 0.224 0.169 0.113 0.071 0.054 0.049 0.034 0.031 0.034
 $\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.488 0.555 0.547 0.459 0.392 0.370 0.384 0.362 0.400 0.382
-P2Y2 Random 0.012 0.005 0.003 0.003 0.001 0.001 0.001 0.000 0.000 0.000
+P2Y2 Random 0.308 0.215 0.170 0.113 0.076 0.058 0.049 0.038 0.033 0.029
 $\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.466 0.514 0.499 0.423 0.340 0.316 0.333 0.312 0.333 0.277
@@
 Fig.
-S2 shows the plots of the results for settings P2Y1 and P2Y2.
+S4 shows the plots of the results for settings P2Y1 and P2Y2.
 For settings P1Y1, P2Y1, P1Y2, and P2Y2, Table S15 shows detailed results used in Fig.
 3 and Fig.
-S2.
+S4.
 As shown in Table 3, settings P1 and P2 show roughly the same trends.
 Similarly, in Table S15, settings P1Y1 and P2Y1, as well as P1Y2 and P2Y2, each show similar trends.
+2.3 Correlation between answer set sizes and predicted ranks
+(a) BioConceptVec
+(b) Our embeddings
+Figure S5: Scatter plots of the answer set sizes for each drug $d$ and the ranks of the predicted genes.
+Fig.
+S5 shows scatter plots and correlation coefficients between the answer set sizes for each drug $d$ and the ranks of the predicted genes obtained by adding relation vectors in settings G, P1, and P2.
+Since smaller answer set sizes are more frequent (see Fig.
+S2 in Supplementary Information 1.4 for the distribution of answer set sizes), we calculated a weighted Pearson correlation coefficient, where the weight is the reciprocal of the frequency of each answer set size.
+2.4 Details of comparison of analogy computation and TransE
+This section provides details regarding the experiments presented in Fig.
+4, comparing our analogy computation with TransE.
+We conducted experiments under setting P1, which is a simple setting and involves multiple drug-gene relations.
+For drug-gene relations in Setting P1, we fixed 20% of the data as validation data and another 20% as test data, following the same setting for TransE in Table 3, and used them consistently across both methods.
+From the remaining 60% of the data, we created training subsets ranging from 10% to 60%, ensuring that each smaller subset is contained within the next larger subset.
+Formally, if we define the 10% to 60% training sets as $\mathcal{T}_{10},\ldots,\mathcal{T}_{60}$, then $\mathcal{T}_{10}\subset\mathcal{T}_{20}\subset\mathcal{T}_{30}\subset\mathcal{T}_{40}\subset\mathcal{T}_{50}\subset\mathcal{T}_{60}$.
+We explain the experimental settings of our method.
+We used our skip-gram embeddings.
+At test time, for pathways present in the training data, we used the relation vector $\hat{\mathbf{v}}_{p}$ defined in Eq.
+(14).
+For pathways not present in the training data, we used the relation vector $\hat{\mathbf{v}}$ computed from all drug-gene relations in the training data, as defined in Eq.
+(7).
+The search space consisted of all genes appearing in the training, validation, and test data.
+The experimental settings for TransE follow those described in Supplementary Information 1.8.